### v1

# Benchmarking Multi-Modal Retrieval for Long Documents

Abstract.
Multi-modal document retrieval is designed to identify and retrieve various forms of multi-modal content, such as figures, tables, charts, and layout information from extensive documents. Despite its significance, there is a notable lack of a robust benchmark to effectively evaluate the performance of systems in multi-modal document retrieval. To address this gap, this work introduces a new benchmark, named as MMDocIR, encompassing two distinct tasks: page-level and layout-level retrieval. The former focuses on localizing the most relevant pages within a long document, while the latter targets the detection of specific layouts, offering a more fine-grained granularity than whole-page analysis. A layout can refer to a variety of elements such as textual paragraphs, equations, figures, tables, or charts. The MMDocIR benchmark comprises a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 173,843 questions, making it a pivotal resource for advancing multi-modal document retrieval for both training and evaluation. Through rigorous experiments, we reveal that (i) visual retrievers significantly outperform their text counterparts, (ii) MMDocIR train set can effectively benefit the training process of multi-modal document retrieval and (iii) text retrievers leveraging on VLM-text perform much better than those using OCR-text. These findings underscores the potential advantages of integrating visual elements for multi-modal document retrieval. MMDocIR is available at https://huggingface.co/MMDocIR.
Multi-modal Document Retrieval, Page Retrieval, Layout Retrieval
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Multi-modal Document Retrieval
Figure 1. MMDocIR comprises 313 lengthy documents across 10 different domains, along with 1,685 questions. For each question, page-level annotations are provided via selected screenshots. Red boundary boxes represent layout-level annotations.
Figure 2. Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark (Ma et al., 2024b). Note that the white spaces, headers, and footers are removed from the area counting.
1. Introduction
Multi-modal document retrieval (Hassan et al., 2013; Ma et al., 2024a; Faysse et al., 2024; Lee et al., 2024) aims to encode and retrieve various forms of content from visually rich documents based on user queries. Unlike traditional document retrieval (Zhang et al., 2022; Chen et al., 2023; Sarthi et al., 2024; Dong et al., 2024) which primarily deals with textual data, multi-modal document retrieval must handle documents that seamlessly integrate text with multi-modal elements such as images, tables, charts, and layout designs. These multi-modal content often carry significant information that plain text cannot convey (Cui et al., 2021; Zhu et al., 2024; Sassioui et al., 2023), thereby enriching the depth and relevance of retrieved content. For example, our analysis of MMLongBench-Doc benchmark (Ma et al., 2024b) reveals that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively, as shown in Figure 2. This significant presence of multiple data modalities requires retriever to leverage all data modalities present in a document.
However, current benchmarks (shown in Table 1) for evaluating multi-modal document retrievers are insufficient, lacking in certain aspects that are critical for a comprehensive assessment. The major shortcomings include: 1. Question Quality: The design and curation of questions in most benchmarks do not align with the specific needs of multi-modal document retrieval. The questions, often reused from Visual Question Answering (VQA) tasks, lack the specificity required for this purpose and require expert filtering to be truly effective for retrieval tasks. 2. Document Quality: A significant limitation of existing benchmarks is the missing of complete document pages, thus hindering accurate evaluation of retrieval capabilities within actual document contexts. Additionally, the range of document domains covered is often too narrow, limiting the benchmarks’ applicability across different fields. 3. Retrieval Granularity: Most benchmarks, except for SciMMIR (Wu et al., 2024), only allow retrieval at a page level which is insufficient, as user queries may pertain to specific parts of a page like a particular figure or table rather than the entire page.
Recognizing these gaps, we introduce MMDocIR Multi-Modal Document Information Retrieval benchmark. MMDocIR is innovatively structured around two critical tasks: page-level and layout-level retrieval. (1) The page-level retrieval task is designed to identify the most relevant pages within a document in response to a user query. (2) The layout-level retrieval aims to retrieve most relevant layouts. The layouts are defined as the fine-grained elements such as paragraphs, equations, figures, tables, and charts. This task allows for a more nuanced content retrieval, honing in on specific information that directly answers user queries. To support these tasks, we first develop MMDocIR evaluation set comprising 313 documents averaging 65.1 pages each, along with VQA 1,658 questions derived from MMLongBench-Doc (Ma et al., 2024b) and DocBench (Zou et al., 2024). Initially, we examine all VQA questions, filter out questions that are irrelevant to IR tasks, and revise remaining questions to be more suitable for IR. Then, we manually annotate these questions with page and layout labels. The page-level labels are the specific pages identified that contain answer evidence.  [^1] It is worth noting that while MMLongBench-Doc provided initial page labels, our meticulous review lead to corrections in about 21.3% of these labels. The layout-level labels are the bounding boxes meticulously delineated around the key evidence within the identified pages. Meanwhile, we introduce MMDocIR training set, comprising 73,843 QA pairs converted from 7 DocQA datasets. Specifically, we manually collect 6,878 documents and use either manual or automatic methods to annotate the ground truth labels.  [^2] Since layout annotations are missing for most datasets, our data annotating pipeline can only create layout-level labels for 4 subsets.
Benchmarks Question Document Label
Type By Expert? For IR? #Num Evidence Type Domain #Pages Source Page Layout
DocCVQA (Tito et al., 2021) VQA question 20 TXT/L Finance 1.0
SciMMIR (Wu et al., 2024) Image caption 530K TAB/I Science 1.0
ViDoRe (Faysse et al., 2024) VQA question 3,810 TXT/C/TAB/I Multi-domain 1.0
PDF-MVQA (Ding et al., 2024) Search query 260k TXT/TAB/I Biomedical 9.6
MMLongBench-Doc (Ma et al., 2024b) VQA question 1,082 TXT/L/C/TAB/I Multi-domain 47.5
Wiki-SS (Ma et al., 2024a) Natural question 3,610 TXT Wikipedia 1.0
DocMatix-IR (Ma et al., 2024a) VQA question 5.61M TXT/L/C/TAB/I Multi-domain 4.2
MMDocIR VQA question 1,658 TXT/C/TAB/I Multi-domain 65.1
Table 1. Comparison between our benchmark and previous Document IR datasets. TXT/L/C/TAB/I: pure text/generalized layout/chart/table/image.
In our research, we conduct a thorough evaluation of the prevailing multi-modal document retrieval baselines which can broadly be categorized into visual-driven and text-driven retrievers. Visual-driven retrievers (Ma et al., 2024a; Faysse et al., 2024), leverage vision-language models (VLMs) (Abdin et al., 2024; Beyer et al., 2024; Bai et al., 2023; Chen et al., 2024b) to capture rich multi-modal cues and generate embeddings for both queries and documents. In contrast, text-driven retrievers (Karpukhin et al., 2020; Khattab and Zaharia, 2020; Izacard et al., 2022; Wang et al., 2022; Xiao et al., 2023; Li et al., 2023) primarily rely on OCR or VLM to first convert the multi-modal content into text, subsequently employing language models (LMs)  (Devlin et al., 2019; Liu et al., 2019; Wang et al., 2024b) to generate embeddings for both queries and documents. Via extensive experiments, we discover that visual-driven retrievers consistently outperform their text-driven counterparts, often by a significant margin. This highlights the critical role that visual elements play in multi-modal document retrieval. In summary, our contributions are three-fold:
  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework (§$\S$ 2) that encompasses page-level and fine-grained layout-level multi-modal document retrieval.
  • MMDocIR Benchmark: We introduce Multi-Modal Document Information Retrieval (MMDocIR) benchmark. The evaluation set (§$\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 QA pairs. The training set (§$\S$ 4) consists of 6,878 documents and labels for 73,843 QA pairs.
  • Comprehensive Evaluation: We conduct an extensive evaluation of existing retrieval systems (§$\S$ 6). Our results reveal that visual-driven retrievers consistently outperform their text-driven counterparts, highlighting the importance of utilizing visual information in multi-modal document retrieval.
2. Dual-Task Retrieval Definition
Let $\mathcal{D}$ be a document corpora consisting of a set of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection [^3] A document page usually comprises about 5 to 15 layouts, depending on its complexity.. The objective is to perform document page-level and layout-level retrieval. Specifically, the task involves retrieving a subset of these pages and layouts that are most relevant to a query $Q$, identified by the top $k$ entries where $k$ is significantly smaller than $n$ or $m$. The relevance of pages ($p$) and layouts ($l$) to $Q$ is quantified by similarity scores, $\textrm{Sim}(Q,p)$ and $\textrm{Sim}(Q,l)$ respectively. The retrieval system can be decomposed into (1) an offline indexing phase in which each page and layout from $\mathcal{P}$ and $\mathcal{L}$ is transformed into a vector representation, and (2) an online querying phase in which a query $Q$ is converted into a vector form, and the vectorized query is then compared against the indexed vectors using the similarity scores $\textrm{Sim}(Q,p)$ for pages and $\textrm{Sim}(Q,l)$ for layouts.
3. MMDocIR: Evaluation Set
{tabu}
l—c@ c@ c@ c@ c@ c@ c—c@   c@  c@   c@   —c@   c@  c@   c@
Dataset Domain Document Statistics   Questions (%)   Modality Distribution (%)
#Doc #QA #Page #Lay #Page #Lay %Lay Text Image Table Lay/ Text Image Table Title
Label Label /Doc /Page Meta
MMDocIR Eval 313 1,658 2,107 2,638 65.1 8.3 41.8 44.7 21.7 37.4 11.5 60.4 18.8 16.7 4.1
[…]
- Research Report 34 200 318 400 39.4 6.0 39.1 45.0 17.5 74.5 13.5 45.6 40.0 5.9 8.4
- Admin & Industry 10 59 82 113 16.8 9.1 45.1 78.0 20.3 13.5 13.5 70.1 11.7 14.9 3.2
- Tut & Workshop 17 102 165 225 57.5 4.1 43.8 37.2 61.7 24.5 9.8 28.0 57.3 6.3 8.3
- Academic Paper 75 386 473 571 19.5 10.1 48.4 28.8 25.7 50.0 10.4 74.6 12.8 11.1 1.5
- Brochure 15 76 121 178 30.3 9.7 41.1 60.5 52.6 18.4 36.8 33.3 50.8 8.5 7.0
- Financial Report 51 343 394 477 169.5 9.2 44.8 28.0 13.1 54.5 5.3 60.3 7.9 29.2 2.6
- Guidebook 22 112 168 223 78.4 10.0 33.6 51.8 54.4 26.8 17.8 63.7 20.0 12.1 4.1
- Government 44 111 116 132 68.9 6.9 45.4 69.37 2.7 0 7.6 88.2 3.7 5.7 2.4
- Laws 44 132 133 149 58.5 6.0 31.2 62.1 0 10.6 27.3 83.8 1.6 12.3 2.2
- News 1 137 137 170 50.0 73.6 72.3 70.1 1.5 0 28.5 48.5 39.8 0.0 11.6
Table 2. Detailed statistics for MMDocIR evaluation set. “#Lay/Page” is the averaging layouts per page, reflecting page’s layout complexity. “%Lay” refers to the area ratio of useful layouts (excluding white spaces, headers, and footers) over entire page.
3.1. Document Corpora Collection
To facilitate the development of MMDocIR, we leverage resources from existing Document Visual Question Answering (DocVQA) benchmarks (Mathew et al., 2021; Tito et al., 2023; Tanaka et al., 2023; Zhu et al., 2022; Li et al., 2024; Landeghem et al., 2023; Ma et al., 2024b; Zou et al., 2024). Although DocVQA benchmarks are not primarily designed for IR, they offer valuable document corpora and questions that can be adapted for our purposes. To select useful DocVQA datasets, we adhere to following criteria:
  • Document Source: The dataset must include accessible original documents or sources for these documents. We need to access and enrich them to support more complex retrieval tasks.
  • Diverse Domain/Modality: The document collections must (1) encompass diverse domains e.g., academia, finance, publishing, and research, and (2) embrace multiple modalities, such as text, figures, tables, charts, and layouts.
  • Long Document: We seek documents of sufficient length as longer texts pose more significant challenges. This criterion can evaluate models in handling complex and lengthy documents.
  • QA Diversity and Comprehensiveness: The questions included in the dataset should be diverse and challenging. For example, some questions should require reasoning across both text and visual tables/figures, while some are multi-hop questions that require synthesizing information from multiple pages.
After considering aforementioned criteria, we choose to leverage on the document corpora and questions from MMLongBench-Doc (Ma et al., 2024b) and DocBench (Zou et al., 2024). MMLongBench-Doc is a long-context, multi-modal benchmark comprising 1,091 questions constructed upon 135 documents averaging 47.5 pages. Its corpora include domains of research report, administration and industry, tutorial and workshop, academic paper, brochure, financial report, and guidebook. DocBench focuses on long document understanding, consisting of 1,102 questions constructed upon 229 documents averaging 77.5 pages. The corpora span across domains: academia, finance, government, laws, and news. Both datasets provide expert-annotated questions, requiring content evidence of various modalities (e.g., text, tables, charts and images, as well as locations layouts) to answer. Consequently, we obtain a total of 364 documents and 2,193 questions for our subsequent annotation.
3.2. Annotation Process
Question Filtering and Revision
To ensure that the questions in MMDocIR are optimally suited for document retrieval tasks. To this end, we identify four specific types of questions that do not align well with the objectives of IR. By filtering out or modifying these questions, we ensure the integrity and relevance of MMDocIR.
  • Summarization Questions: These questions, such as “What does this book mainly illustrate?” require a comprehensive understanding of large sections of a document or even the entire document. The broad scope makes it hard to pinpoint specific content effectively, which contradicts the precise nature of IR tasks.
  • Overwhelm Statistical Questions: Questions that demand extensive data computation or collation, such as “How many words are there in total in the paper?” also fall outside our scope. These require a level of statistical analysis that goes beyond retrieval.
  • Online Search Questions: For example, queries like “What is the Google Scholar citation count of the main author of this article?” rely on information from external online resources. Since our focus is on retrieving information that is internally available within the documents, these questions are excluded.
  • Unanswerable Questions: These are designed to test if models generate answers based on non-existent information (model hallucinations). Since they do not facilitate the retrieval of factual document-based information, these questions are excluded.
Page-level annotation process
To address the challenge of accurately locating relevant information within extensive documents, we annotate page-level labels that pinpoint the exact pages containing evidence needed to answer specific questions. Considering that the average document in our dataset contains 65.1 pages, pinpointing the correct pages is similar to finding a needle in a haystack and requires meticulous effort. Our annotation process is:
  • For MMLongBench-Doc: we review and correct the answers and corresponding page labels for 794 questions. Specifically, we make corrections to 10 answers and revise 169 page labels [^4] Common errors in page labeling: annotators starting page indexing at 1, missing labels for questions spanning multiple pages, and incorrect or absent page labels. to ensure they accurately reflect the evidence locations.
  • For DocBench: we annotate page labels for all 864 questions from scratch. This involves a detailed examination of each document to accurately assign page labels that match the evidential requirements of each question.
Through these efforts and following a rigorous annotation and cross-validation process, we successfully compile page-level labels for a total of 1,658 questions in MMDocIR.
Layout-level annotation process
To achieve more fine-grained retrieval granularity, we go beyond page-level annotations by including layout-level labels. This advanced annotation process captures the specific layout elements within pages that contain necessary evidence to answer each question. Unlike page-level annotation, labeling layout-level information is more complex and labor-intensive due to its detailed nature. Our annotations are via:
  • To streamline annotation process, we utilize MinerU (Wang et al., 2024a) to parse all documents and extract the layout information (e.g., type and bounding boxes) for every layout element on a page. These elements are categorized into five main types: text, image, table, title, and equation. The bounding boxes are delineated by the coordinates of the top-right and bottom-left corners.
  • Based on the layouts drawn by MinerU, we identify the layouts that contain necessary eivdence to answer the question. In instances where MinerU fails to detect certain evidentiary elements, we manually draws the bounding boxes to ensure the precision of layout labels. This manual intervention is necessary for approximately 7% of our layout-level labels.
Ultimately, this meticulous process leads to the annotation of 2,638 layout labels for the 1,658 questions in MMDocIR.
3.3. Quality Control
Consistency Page Labels Layout Labels
Prec. Recall F1 Prec. Recall F1
[…]
Average 95.0 95.4 95.2 87.0 87.2 87.1
[…]
{tabu}
l—c—ccc@ c@ c—c@   c@  c@   c@   —c@   c@
Dataset Domain Document Statistics   Modality Distribution (%)   Labels
#Doc #QA #Page #Lay %Lay Text Image Table Title Page Lay
/Doc /Page
MMDocIR Train assorted docs 6,878 73,843 32.6 6.32 42.6 49.3 34.3 10.8 4.9
[…]
- MP-DocVQA (Tito et al., 2023) health/ind. docs 875 15,266 46.8 6.9 38.8 57.3 18.0 22.7 1.9
- SlideVQA (Tanaka et al., 2023) diverse slides 2,011 11,066 49.3 4.4 42.3 30.1 56.2 4.7 8.8
- TAT-DQA (Zhu et al., 2022) annual reports 163 15,814 147.3 9.2 42.2 66.4 4.4 26.5 2.7
- ArXivQA (Li et al., 2024) arXiv papers 1,579 12,314 18.4 7.9 50.0 70.4 22.3 2.8 1.0
- SciQAG (Wan et al., 2024) science papers 1,197 4,976 9.0 9.1 53.7 61.8 28.0 6.7 1.5
- DUDE (Landeghem et al., 2023) assorted docs 779 3,173 15.6 7.4 42.5 57.1 24.7 15.2 2.9
- CUAD (Hendrycks et al., 2021) legal contract 274 11,234 29.6 7.4 24.7 89.3 2.5 6.4 1.1
Table 4. Document statistics for Training Datasets collected.
To ensure annotation quality and reliability in MMDocIR, we have adopted a rigorous quality control process using a cross-validation method. This method involves splitting the document corpora into two main groups, labeled as Group A and Group B. Each group is responsible for annotating page-level and layout-level labels for approximately 1,000 QA pairs. Importantly, there is an overlap of 400 QA pairs between two groups, which serves as a validation checkpoint. The process for quality control works as follows:
  • Overlap Scoring: For 400 overlapping questions, each group’s annotations are scored against the other’s as a form of mutual validation. Specifically, Group A’s annotations are used as a ground truth to score Group B’s work, and conversely, Group B’s annotations serve as the ground truth for scoring Group A.
  • Cross-Evaluation: By cross-evaluating the outcomes from both groups, we are able to assess the consistency of the annotations and identify any discrepancies for improvement. Specifically, we achieved an F1 score of 95.2 for page-level annotations and 87.1 for layout-level annotations, as shown in Table 3.
  • Random Cross-Validation: Apart from the overlapping 400 questions, we randomly cross-validate around 50% of the remaining annotations. In the cases where we have different opinions, we discuss to achieve mutually-agreed annotations. This step ensures the high quality and reliability of MMDocIR.
3.4. Statistics and Analysis
Document Analysis. As shown in Table 3, MMDocIR evaluation set includes 313 long documents averaging 65.1 pages, categorized into ten main domains. Different domains feature distinct distributions of multi-modal information. For instance, research reports, tutorials, workshops, and brochures predominantly contain images, whereas financial and industry documents are table-rich. In contrast, government and legal documents primarily comprise text. Overall, the modality distribution is: Text (60.4%), Image (18.8%), Table (16.7%), and other modalities (4.1%).
Question and Annotation Analysis. MMDocIR encompasses 1,658 questions, 2,107 page labels, and 2,638 layout labels. The modalities required to answer these questions distribute across four categories: Text (44.7%), Image (21.7%), Table (37.4%), and Layout/Meta (11.5%). The “Layout/Meta” category encompasses questions related to layout information and meta-data statistics. Notably, the dataset poses several challenges: 254 questions necessitate cross-modal understanding, 313 questions demand evidence across multiple pages, and 637 questions require reasoning based on multiple layouts. These complexities highlight the need for advanced multi-modal reasoning and contextual understanding.
4. MMDocIR: Training Set
4.1. Document Corpus Collection
The training set corpora are mainly collected from DocVQA benchmarks, adhering to similar selection criteria (see Section 3.1).
  • MP-DocVQA (Tito et al., 2023) contains 47,952 images collected from Industry Documents Library (IDL)  [^5] https://www.industrydocuments.ucsf.edu/. We group the 47,952 document images into separate document files, and obtain 875 long documents (46.8 pages on average) with 15,266 QA pairs.
  • SlideVQA (Tanaka et al., 2023) contains 2,619 slide documents collected from slideshare  [^6] https://www.slideshare.net/ and covering 39 topics. Note that SlideVQA contains only the first 20 pages of each slide decks. In our research, we manually collect the complete slide decks, and obtain 2,011 long documents (averaging 49.3 pages) with 11,066 QA pairs.
  • TAT-DQA (Zhu et al., 2022) consists of 3,067 document pages from financial reports  [^7] https://www.annualreports.com/, dated between 2018 and 2020. Note that neither original documents nor links is provided. We use OCR to extract text in the pages, and use search engine to find relevant documents. After careful tracing and recognition, we identify 163 original documents (averaging 147.3 pages) with 15,814 QA pairs.
  • ArXivQA (Li et al., 2024) comprises 32k figures cropped from academic pages  [^8] https://arxiv.org/. We use the arXiv DOIs provided to collect the academic papers. After careful tracing, recognition, and document length filtering, we identify 1,579 documents averaging 18.4 pages.
  • SciQAG (Wan et al., 2024) consists of 22,728 papers in 24 scientific disciplines, collected from Web of Science (WoS) Core Collection database. We sample 50 documents from each discipline, and manually collect 1,197 papers using the DOIs provided.
  • DUDE (Landeghem et al., 2023) provides 5,019 documents from aggregate websites9991: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/. It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires. We filter out short documents and obtain 779 relatively long documents (averaging 15.6 pages) with 3,173 QA pairs.
  • CUAD (Hendrycks et al., 2021) provides 510 commercial legal contracts, collected from Electronic Data Gathering, Analysis, and Retrieval (EDGAR) [^10] https://www.sec.gov/search-filings. We filter out short documents in CUAD and obtain 274 long documents (29.6 pages on average) with 11,234 QA pairs.
4.2. Label Construction
The page labels can be converted for MP-DocVQA, SlideVQA, and DUDE datasets. Among them, only DUDE provides layout labels.
SciQAG provides only question and answer in texts. We utilize them to reversely obtain the page-level and layout-level labels. Specifically, we first use MinerU to obtain layout-level passage chunks. For each QA pair, we deploy E5 and BGE retrievers to obtain question-passage and answer-passage similarity scores against all extracted passage chunks. If both scores rank within top 3 for a certain passage chunk, we confirm this layout to be the layout-level labels for the given QA pair.
Similarly, ArXivQA provides only cropped images, without document page/layout labels. We first use MinerU to obtain layout-level image. For each cropped image, we calculate the image similarity between all extracted images, and keep only the highest scoring one. Subsequently, we manually examine if the selected image matches the cropped image. In this way, we filter around 20% unmatched images and their QA pairs, resulting 1,579 QA pairs with page-level and layout-level labels.
For TAT-DQA, the layout-level labels are provided for sampled single page. We need to localize the page index of the sampled page. We utilize PDF mapping technique to retrieve the best matched page for every sampled page in this document. Then, we manual examine if the retrieved page is identical to the given page, and correct the labels if there were any errors.
The overall statistics (e.g., document information, modality distribution, domain, etc) of MMDocIR training set are in Table 3.3.
{tabu}
lll—cccccccccc—cc
Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
Report &Indu. Worksh. Paper ure Report book ment Macro Micro
Recall@$k=1$
OCR-text
DPR 21.2 22.1 27.7 23.3 24.4 16.7 21.1 20.7 31.0 15.1 22.3 21.7
ColBERT 43.8 39.8 42.4 39.3 39.2 38.7 46.3 50.6 46.1 17.1 40.3 40.0
BGE 45.5 29.0 41.5 33.6 40.8 32.7 40.0 42.8 36.4 15.1 35.7 35.2
E5 44.2 30.8 39.9 33.2 33.0 32.3 40.4 41.7 38.9 15.8 35.0 34.7
Contriever 39.1 33.3 44.0 34.2 43.9 26.4 40.6 39.4 37.0 15.1 35.3 33.6
GTE 44.6 32.6 45.0 33.2 37.2 31.8 40.0 39.9 35.2 14.5 35.4 34.6
[…]
VLM-text
DPR 32.3 25.5 27.0 31.0 28.4 18.8 23.5 31.2 38.3 16.1 27.2 26.9
ColBERT 48.6 42.8 51.1 46.2 36.0 36.8 49.6 60.9 59.5 26.3 45.8 44.9
BGE 48.8 30.9 47.1 40.8 37.6 28.4 43.4 51.9 48.9 28.5 40.6 39.6
E5 48.1 30.0 50.4 39.4 41.1 29.7 40.9 52.8 51.1 24.1 40.8 39.5
Contriever 45.5 31.2 49.8 41.5 39.4 29.4 45.2 55.3 51.1 20.4 40.9 39.7
GTE 46.5 26.3 48.7 38.9 35.9 27.0 46.2 50.1 45.8 24.1 38.9 37.9
[…]
Image
DSEwiki-ss 53.0 50.0 54.0 48.7 45.1 43.0 51.5 46.9 54.2 33.6 48.0 47.5
DSEdocmatix 52.3 40.4 56.1 51.7 45.8 43.5 53.8 53.7 58.3 46.7 50.2 50.1
ColPali 56.0 51.8 58.6 55.9 52.0 47.2 57.9 53.9 64.0 32.8 53.0 52.7
DPR-Phi3ours 58.9 50.4 57.4 59.0 57.3 44.6 63.8 50.5 64.4 35.0 54.1 53.7
Col-Phi3ours 56.7 50.4 56.9 61.3 54.8 50.7 60.8 61.3 63.6 54.0 57.0 57.1
Recall@$k=3$
OCR-text
DPR 46.1 40.6 38.9 46.7 43.9 32.4 38.4 37.0 50.0 20.4 39.4 39.8
ColBERT 72.6 59.7 57.8 66.7 60.0 53.7 63.8 68.5 61.4 23.7 58.8 59.5
BGE 69.8 57.7 56.3 58.6 60.7 48.5 57.9 60.9 62.7 20.4 55.4 55.0
E5 66.6 48.7 59.0 58.0 60.9 48.8 63.7 61.4 60.8 20.4 54.8 54.6
Contriever 70.2 55.8 60.4 56.6 62.1 43.0 60.0 56.8 61.4 22.4 54.9 53.6
GTE 69.2 47.0 58.7 59.5 61.8 46.6 65.5 59.1 61.4 19.7 54.9 54.7
[…]
VLM-text
DPR 52.2 44.2 43.5 54.6 52.0 35.1 44.4 53.9 57.2 25.5 46.3 46.2
ColBERT 70.1 64.4 70.3 72.3 59.1 55.3 71.1 81.3 70.8 34.3 64.9 64.8
BGE 71.5 48.2 68.8 65.7 56.2 46.5 66.1 69.9 72.0 32.1 59.7 59.6
E5 68.4 45.7 68.1 63.7 60.1 44.0 69.3 72.3 78.8 32.8 60.3 59.3
Contriever 69.4 55.3 68.3 64.9 56.9 46.2 69.9 71.1 72.0 32.1 60.6 59.7
GTE 71.1 44.5 67.2 64.4 54.3 43.0 70.6 71.9 68.2 31.4 58.7 58.3
[…]
Image
DSEwiki-ss 75.4 65.0 73.9 79.8 69.5 63.5 75.4 71.5 81.4 50.4 70.6 71.4
DSEdocmatix 75.4 67.5 73.3 80.0 66.3 61.6 72.8 76.4 82.6 57.7 71.4 71.8
ColPali 77.6 71.8 79.4 83.4 72.6 66.1 80.0 80.4 86.4 49.6 74.7 75.0
DPR-Phi3ours 80.3 66.5 77.6 83.9 71.9 63.8 79.8 71.4 84.5 55.5 73.5 74.3
Col-Phi3ours 80.2 74.1 77.4 84.8 69.1 67.7 78.7 79.5 81.8 69.3 76.3 76.8
Recall@$k=5$
OCR-text
DPR 59.5 55.8 43.4 59.1 56.2 41.2 50.7 45.5 56.0 23.0 49.0 49.4
ColBERT 78.4 71.1 63.3 75.2 68.8 60.5 72.0 72.7 67.5 30.3 66.0 66.5
BGE 79.3 65.9 62.1 69.7 69.8 56.5 68.0 62.8 66.3 26.3 62.7 62.9
E5 79.3 62.4 67.0 70.3 71.8 57.6 72.5 67.1 67.5 25.7 64.1 64.2
Contriever 79.9 62.7 64.7 71.7 71.1 48.8 72.4 65.8 67.5 26.3 63.1 62.5
GTE 78.3 61.9 67.3 72.3 68.7 55.2 72.6 64.0 67.5 24.3 63.2 63.5
[…]
VLM-text
DPR 66.5 60.1 56.0 68.9 58.8 43.8 57.1 68.6 64.8 33.6 57.8 57.8
ColBERT 78.8 74.0 78.7 82.3 66.1 60.8 77.0 88.5 78.0 38.7 72.3 72.3
BGE 79.5 65.8 71.3 76.8 62.4 56.0 77.2 77.4 79.5 38.0 68.4 68.5
E5 76.9 64.2 75.3 74.4 67.4 52.0 78.5 78.6 82.6 40.9 69.1 67.9
Contriever 77.2 67.1 76.7 75.2 65.1 53.7 75.4 79.2 83.3 39.4 69.2 68.3
GTE 77.4 62.6 74.7 75.8 62.0 51.8 77.8 80.0 75.0 39.4 67.6 67.2
[…]
Image
DSEwiki-ss 84.0 80.2 78.7 87.0 75.7 73.0 82.0 77.3 88.3 58.4 78.5 79.2
DSEdocmatix 82.1 77.2 79.6 87.8 73.9 72.4 81.7 83.1 89.4 67.9 79.5 80.1
ColPali 84.6 79.3 82.3 89.0 79.8 72.1 86.7 84.9 92.4 56.9 80.8 81.0
DPR-Phi3ours 86.9 76.2 85.3 91.9 80.0 71.2 87.1 79.5 92.0 61.3 81.1 81.8
Col-Phi3ours 86.3 78.8 81.2 92.4 79.0 73.8 85.3 85.1 87.1 73.0 82.2 83.0
Table 5. Main results for page-level retrieval. “OCR-text” and “VLM-text” refer to converting multi-modal content in the document page using OCR and VLM respectively. “Image” refers to processing document page as screenshot image.
5. Model Training: DPR-Phi3&Col-Phi3
To evaluate the effectiveness of the MMDocIR training set, we train two visual retrievers based on Phi3-Vision (Abdin et al., 2024). Phi3-Vision ($\mathbf{M_{phi3v}}$) reuses the image tokenizer from clip-vit-large [^11] ViT-Large: https://huggingface.co/openai/clip-vit-large-patch14-336 ($\mathbf{M_{vit}}$). It can deal with high-resolution images by cropping them into sub-images, where each sub-image has $336\times 336$ pixels.
Document/Query Encoding
DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR (Karpukhin et al., 2020)) and a list of token-level embeddings (following ColBERT (Khattab and Zaharia, 2020)), respectively. Specifically, we follow Ma et al. (2024a) to concatenate document image with a text prompt: “¡s¿¡d¿ What is shown in this image?¡/s¿”. Here, the ¡d¿ token is a special placeholder token and is replaced by the sequence of patch latent embeddings from the vision encoder. We consider only text queries and use text prompt: “¡s¿ query: ¡q¿ ¡/s¿”. Similarly, the placeholder ¡q¿ token is replaced by input query.
(1) $\begin{split}\mathrm{E_{d}^{dpr}}&=\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\\ &\mathrm{E_{q}^{dpr}}=\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\end{split}$
where the end-of-sequence token ¡/s¿ from the last hidden state ($D_{1}=3072$) of $\mathbf{M_{phi3v}}$ is used to represent $\mathrm{E_{d}^{dpr}}$ and $\mathrm{E_{q}^{dpr}}$.
(2) $\begin{split}\mathrm{E_{d}^{col}}&=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)},\in\mathbb{R}^{N_{d}\times D_{2}}\\ &\mathrm{E_{q}^{col}}=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)},\in\mathbb{R}^{N_{q}\times D_{2}}\end{split}$
where $\mathbf{M_{proj}}$ is projection layer to map the last hidden states of $\mathbf{M_{phi3v}}$ into reduced dimension $D_{2}=128$. $N_{d}\approx 2500$ for a typical high-resolution page and $N_{q}$ is the number of query tokens.
Query-Doc Similarity
The similarity between the query and the document is computed as follows:
(3) $\text{Sim}(q,d)_{dpr}=\frac{\langle\mathrm{E_{q}^{dpr}}|\mathrm{E_{d}^{dpr}}\rangle}{\left\|\mathrm{E_{q}^{dpr}}\right\|\cdot\left\|\mathrm{E_{d}^{dpr}}\right\|}$
[…]
(4) $\text{Sim}(q,d)_{col}=\sum_{i\in[1,N_{q}]}\max_{j\in[1,N_{d}]}\langle\mathrm{E_{q}^{col}}^{(i)}|\mathrm{E_{d}^{col}}^{(j)}\rangle$
[…]
Contrastive Loss
Given the query $q$, we have the positive document $d^{+}$ and a set of negative documents $d^{-}$ including hard negatives and in-batch negatives. The hard negatives are negative pages within the document with highest $Sim(q,d^{-})$ scored by ColPali (Faysse et al., 2024) retriever. We calculate the loss as:
(5) ($\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$
where DPR-Phi3 is trained on the InfoNCE loss, and the temperature parameter $\tau=0.02$ in our experiments.
(6) ($\mathcal{L}^{col}_{(q,d^{+},d^{-})}=\log\Big{(}1+\exp\big{(}\max_{d_{i}\in d^{-}}(\text{Sim}^{col}_{(q,d_{i})})-\text{Sim}^{col}_{(q,d^{+})}\big{)}\Big{)}$
where Col-Phi3 is trained via the softplus loss based on the positive scores w.r.t. to the maximal negative scores.
6. Experiment
{tabu}
llll—cccccccccc—cc
Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
Report &Indu. Worksh. Paper ure Report book ment Macro Micro
Recall@$k=1$
Textual Retrieval
OCR-text
DPR 3.4 7.2 1.2 11.3 3.0 9.8 8.2 24.7 30.9 26.3 12.6 12.4
ColBERT 5.0 8.8 4.7 16.4 2.0 13.2 4.6 50.8 47.7 45.3 19.8 19.1
BGE 7.0 10.9 3.7 14.3 2.3 16.3 8.4 46.1 45.5 35.8 19.0 18.5
E5 6.3 6.0 2.6 14.0 3.5 14.4 6.1 44.7 45.4 40.5 18.4 17.9
Contriever 6.7 7.0 3.8 14.3 4.1 13.3 8.3 43.6 43.9 42.9 18.8 18.1
[…]
VLM-text
DPR 11.6 9.5 19.2 19.2 14.9 15.9 15.8 25.6 34.7 27.0 19.3 19.2
ColBERT 22.0 14.9 28.0 28.3 17.9 29.7 21.1 52.6 54.5 44.5 31.3 31.4
BGE 19.2 15.2 24.6 28.7 12.8 27.6 19.7 47.0 52.3 35.8 28.3 29.0
E5 15.9 8.8 27.7 24.3 14.6 21.8 14.7 45.6 53.0 40.5 26.7 26.4
Contriever 23.4 7.5 28.2 26.8 17.1 25.7 16.1 43.6 51.5 42.9 28.3 28.9
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
DSEdocmatix 19.9 11.4 31.5 30.1 17.8 30.0 20.8 46.5 39.4 31.4 27.9 29.1
ColPali 22.5 21.3 36.6 30.9 26.8 32.1 19.3 52.5 51.8 33.6 32.7 32.5
DPR-Phi3ours 21.1 22.1 36.8 35.2 25.6 28.7 24.1 38.3 35.4 27.4 29.5 30.2
Col-Phi3ours 22.6 22.0 37.5 34.9 28.9 30.3 22.7 50.2 45.1 26.3 31.1 31.6
[…]
Hybrid
DSEwiki-ss 14.0 10.4 29.8 18.0 13.7 20.4 13.5 46.0 45.1 34.7 24.6 23.4
DSEdocmatix 18.2 11.6 32.7 24.0 17.7 27.2 16.7 48.1 45.5 33.0 27.5 27.4
ColPali 17.7 12.3 30.0 18.4 19.0 25.5 20.6 49.7 51.2 40.9 28.5 27.1
DPR-Phi3ours 28.3 11.1 35.5 39.2 29.3 25.2 27.4 37.0 32.9 22.2 28.8 30.5
Col-Phi3ours 26.4 12.6 33.7 37.5 30.1 27.9 24.6 46.2 47.4 21.9 30.8 32.0
Recall@$k=5$
Textual Retrieval
OCR-text
DPR 7.3 12.5 5.6 24.0 8.9 16.9 13.6 47.2 50.2 51.1 23.7 23.5
ColBERT 10.9 23.8 10.2 32.2 6.8 25.5 17.0 78.7 63.4 63.2 33.2 32.6
BGE 11.9 20.3 13.6 30.0 11.7 27.7 18.8 68.5 65.4 59.1 32.7 32.2
E5 12.8 16.2 8.9 31.9 10.9 23.6 19.9 76.1 68.8 63.9 33.3 32.7
Contriever 11.9 17.9 11.9 28.8 9.3 24.6 18.1 64.3 68.0 62.7 31.7 31.2
[…]
VLM-text
DPR 31.0 25.7 36.7 44.9 33.0 34.1 34.9 49.9 56.3 51.1 39.8 40.4
ColBERT 41.8 37.7 53.7 61.8 35.1 52.4 46.1 83.2 70.1 62.5 54.4 56.0
BGE 41.0 28.1 52.7 59.2 36.7 46.0 50.7 72.0 71.5 59.9 51.8 53.2
E5 35.4 28.1 51.7 58.5 33.2 41.2 40.2 79.7 77.9 64.6 51.1 51.8
Contriever 40.2 29.2 54.1 57.9 36.8 47.1 44.6 68.8 76.2 61.9 51.7 53.0
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
DSEdocmatix 39.6 36.2 53.9 57.5 33.7 52.5 42.8 69.4 63.1 48.9 49.8 51.9
ColPali 40.7 45.9 54.9 58.5 42.6 51.2 45.7 76.8 74.5 48.9 54.0 54.3
DPR-Phi3ours 45.5 37.7 57.0 62.9 41.4 51.1 45.5 65.1 60.8 49.3 51.6 53.9
Col-Phi3ours 46.4 38.2 53.1 61.8 45.0 54.6 45.7 68.8 65.7 43.8 52.3 54.5
[…]
Hybrid
DSEwiki-ss 31.8 29.5 51.1 43.0 34.5 39.3 38.3 71.3 71.3 57.1 46.7 45.6
DSEdocmatix 37.3 26.7 48.2 49.7 34.5 48.6 41.0 72.2 69.4 54.2 48.2 49.3
ColPali 40.1 38.3 55.2 49.2 42.7 47.6 40.8 78.6 68.7 61.0 52.2 51.4
DPR-Phi3ours 54.2 27.4 53.8 64.9 36.6 45.3 49.6 69.5 63.8 53.2 51.8 54.4
Col-Phi3ours 50.9 25.5 49.1 66.2 40.3 48.1 48.3 62.3 60.6 48.5 50.0 53.5
Recall@$k=10$
Textual Retrieval
OCR-text
DPR 10.5 21.1 8.8 32.0 14.9 19.6 17.0 59.7 56.4 58.9 29.9 29.3
ColBERT 14.1 31.1 13.1 38.4 9.1 31.0 22.4 83.9 67.9 65.4 37.6 37.4
BGE 15.7 24.3 15.9 35.9 17.9 31.6 25.3 76.3 73.4 62.1 37.8 37.3
E5 16.9 24.5 13.8 40.1 15.8 26.7 24.7 83.5 77.9 66.1 39.0 38.3
Contriever 15.1 25.7 14.2 36.8 15.9 27.9 24.2 76.8 71.8 64.9 37.3 36.6
[…]
VLM-text
DPR 42.2 33.1 52.1 56.2 39.9 43.5 44.0 62.8 61.7 59.7 49.5 50.5
ColBERT 51.0 48.7 60.6 69.8 43.9 61.6 53.7 88.4 74.8 66.4 61.9 63.7
BGE 51.1 38.7 62.1 71.5 41.9 55.6 58.7 80.8 78.7 63.5 60.3 62.4
E5 45.3 38.6 62.0 70.5 45.6 50.0 55.3 87.1 82.4 66.8 60.4 61.2
Contriever 49.9 41.3 62.0 70.5 44.8 56.5 54.5 81.3 78.0 64.9 60.4 62.2
[…]
Visual Retrieval
Pure-Image
DSEwiki-ss 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
DSEdocmatix 53.7 43.3 59.6 66.5 44.7 59.1 50.3 75.4 69.2 53.7 57.5 59.9
ColPali 53.6 54.1 64.4 69.5 48.8 60.7 54.0 81.9 82.5 50.4 62.0 63.2
DPR-Phi3ours 58.1 49.1 67.0 74.7 48.4 57.9 57.8 68.7 66.2 54.4 60.2 62.8
Col-Phi3ours 57.7 50.5 66.6 72.3 50.7 59.3 53.6 68.5 74.8 57.5 61.1 63.3
[…]
Hybrid
DSEwiki-ss 44.1 34.3 57.6 56.3 42.6 50.7 48.6 81.1 79.1 63.7 55.8 56.0
DSEdocmatix 49.9 37.3 57.3 61.3 45.9 57.9 50.1 77.9 74.9 61.5 57.4 58.9
ColPali 52.1 46.4 65.0 64.4 50.7 53.8 51.0 82.7 71.4 62.5 60.0 60.2
DPR-Phi3ours 65.2 33.7 60.3 72.9 42.4 54.1 52.9 79.1 72.5 65.6 59.9 62.9
Col-Phi3ours 59.1 38.7 57.0 77.7 43.9 57.7 51.7 72.4 68.0 60.7 58.7 62.8
Table 6. Main results for layout-level retrieval. “OCR-text” and “VLM-text” refer to converting multi-modal layouts using OCR and VLM respectively. “Pure-Image” and “Hybrid” refer to reading textual layouts in image and text format respectively.
6.1. Evaluation Metric
The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores. Recall@$k$ is defined as the proportion of the ground truth page/layout evidence that is successfully retrieved. Note for layout matching, we calculate recall based on the overlaps of bounding boxes between retrieved layouts and gold layouts.
6.2. Baseline Models and Setting
We evaluate 6 state-of-the-art text retrievers: namely DPR (Karpukhin et al., 2020), ColBERT (Khattab and Zaharia, 2020), BGE (Xiao et al., 2023), E5 (Wang et al., 2022), Contriever (Izacard et al., 2022), and GTE (Li et al., 2023) (refer to Section 7.1). Meanwhile, we evaluate 5 VLM-based retrievers: 3 off-the-shelf models named DSEwiki-ss (Ma et al., 2024a), DSEdocmatix (Ma et al., 2024a), and ColPali (Faysse et al., 2024) (refer to Section 7.2), and 2 our trained models (see section 5). All retrievers are adapted into dual-task setting:
  • Page Retrieval: For textual retrievers, we use MinerU (Wang et al., 2024a) to extract the layout of all document pages. Non-textual layouts are converted into ‘OCR-text” and “VLM-text” by using Tesseract OCR  [^12] https://github.com/tesseract-ocr/tesseract and GPT-4o respectively. The document page is represent by concatenating the text of all layouts within it. For visual retrievers, we directly utilize document page screenshots.
  • Layout Retrieval: Textual retrievers process non-textual layouts using “OCR-text” or “VLM-text”. Visual retrievers process textual layouts in either “Pure Image” (using cropped image of textual area as input) or “Hybrid” (using its original text, as VLM can directly encode text as well).
6.3. Main Results for Page-level Retrieval
For page-level retrieval (Table 4.2), our key findings are as follows:
  • Superiority of Visual Retrievers: Visual retrievers consistently outperform text retrievers across various domains and retrieval metrics (e.g., Top $k=1,3,5$). This highlights the significance of leveraging document screenshots to capture multi-modal information, which is often lost when relying solely on OCR-text.
  • Effectiveness of MMDocIR: Among the visual retrievers, DPR-Phi3ours and Col-Phi3ours trained on theMMDocIR train set demonstrate superior performance. It suggests dataset quality to effectively enhances the retrieval ability.
  • Effectiveness of VLM-Text: Although VLM-text approaches underperform visual retrievers, they achieve much better performance than the OCR-text methods. This indicates benefits of using GPT-4o to preserve visual cues in text.
  • Necessity of Token-level Embeddings: Token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) compared with their document/page-level counterparts (e.g., BGE, DSE, DPR-Phi3ours), achieve more advantageous results in Recall@1 and marginal performance increase in Recall@5/10. Yet, the storage overhead of token-level embeddings can be 10 times more than page-level embeddings.
  • Top 5 Coverage: Retrieving top 5 pages provides substantial coverage, ensuring that relevant information is captured.
6.4. Main Results for Layout-level Retrieval
For layout-level retrieval (Table 6), our key findings are as follows:
  • Superiority of Visual Retrievers: Visual retrievers exhibit substantial performance advantages over text retrievers utilizing OCR-text. This highlights the limitations of OCR in capturing the multi-modal nature of documents, where visual context can significantly enhance retrieval accuracy.
  • Effectiveness of VLM-Text: Interestingly, VLM-text approaches achieve comparable performance as visual retrievers. This indicates strong image description capabilities of state-of-the-art VLM, which can offer significant benefits to textual retrievers in understanding multi-modal information.
  • Comparison of Hybrid vs. Pure Image Sequences: Visual retrievers relying on hybrid image-text sequences generally perform less effectively than those utilizing pure image sequences. This suggests that current VLMs may have stronger capabilities in modeling images than text within the multi-modal framework.
  • Necessity of Token-level Embeddings: Token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) compared with their document/page-level counterparts (e.g., BGE, DSE, DPR-Phi3ours), do not achieve advantageous results in layout retrieval.
  • Top 10 Coverage: Retrieving top 10 layouts does not provide comprehensive coverage of the ground truth layouts for document retrieval tasks. It suggests the challenge of layout retrieval.
6.5. Analysis of OCR and VLM Text
Interestingly, text retrievers leveraging VLM-text can significantly outperform OCR-text, in both page and layout retrieval. OCR-text which is the the raw text extracted via OCR tools, and as experimental results suggest, it is not suitable for multi-modal retrieval. In comparison, VLM-text can facilitate multi-modal retrieval, suggesting that VLMs can largely preserve rich multi-modal information. The averaged word length and distribution of OCR and VLM text obtained from the tables and figures in MMDocIR in shown in Figure 3. Observe that the VLM-text is much longer than OCR-text. Specifically, the length of VLM-text is 1.5 and 3.8 times more than that of OCR-text in table and figure. Tables tend to contain numerical or structured layouts, which can be well recorded by text. Figures contain mostly graphical elaboration and visual cues. Hence, making OCR tools struggle in getting textual information from figures. Although VLM-text can offer much comprehensive text information than OCR-text, it comes at the cost of higher computational overhead and longer latency.
(a) Avg word length
(b) Distribution density of word length
Figure 3. Avg length and distribution of OCR and VLM text.
7. Related Work
7.1. Text-Centric Document Retrieval
Document indexing.
The process of indexing a multi-modal document involves multiple steps, including Document Parsing (Chao and Fan, 2004; Wang et al., 2024a), Optical Character Recognition (OCR) (Chaudhuri et al., 2017; Borovikov, 2014; Mori et al., 1999), Layout Detection (Sassioui et al., 2023; Xu et al., 2020, 2021), Chunking (Chen et al., 2024a; Raina and Gales, 2024; Dong et al., 2024), and Image Captioning (You et al., 2016; Aneja et al., 2018). These steps are time-consuming and can introduce errors that impact the overall retrieval performance.
Text retrieval.
Current text retrieval are primarily classified into sparse and dense retrieval. For two widely-used sparse retrievers: TF-IDF (Salton et al., 1983) calculates the relevance via word frequency with the inverse document frequency, and BM25 (Robertson et al., 1994) introduces nonlinear word frequency saturation and length normalization. Dense retrievers encode content into vector representations. DPR (Karpukhin et al., 2020) is the pioneering work of dense vector representations for QA tasks. Similarly, ColBERT (Khattab and Zaharia, 2020) introduces an efficient question-document interaction model with late fine-grained term matching. Contriever (Izacard et al., 2022) leverages contrastive learning to improve content dense encoding. E5 (Wang et al., 2022) and BGE (Xiao et al., 2023) propose novel training and data preparation techniques to enhance retrieval performance. Moreover, GTE (Li et al., 2023) integrates graph-based techniques to enhance dense embedding. Most systems focus on text-based representations, neglecting the valuable visual information present in documents.
7.2. Vision-Driven Document Retrieval
Vision Language Models (VLMs) (Abdin et al., 2024; Beyer et al., 2024; Bai et al., 2023; Chen et al., 2024b) have gained popularity for their ability to understand and generate text based on combined text and visual inputs. This advancement has led to the development of cutting-edge visual-driven retrievers, such as ColPali (Faysse et al., 2024) and DSE (Ma et al., 2024a). These models specifically leverage PaliGemma (Beyer et al., 2024) and Phi3-V (Abdin et al., 2024) to directly encode document page screenshots for multi-modal document retrieval. ColPali adopts similar question-document interaction as ColBERT, and represent each document page in token-level embeddings. By contrast, DSE is similar to DPR to encode each page with a single dense embedding. Visual Retrievers can directly model useful visual information, allowing for the direct utilization of multi-modal content without first converting it into text. Despite these advancements, visual retrievers face challenges, particularly in dealing with text details when document page resolutions are high. The high resolution of document pages substantially increases the computational cost and complexity of the embedding process, which may hinder the model’s performance.
7.3. Multi-modal Document Retrieval Datasets
As described in Section 1 and Table 1, there is a notable lack of a robust benchmark for multi-modal document retrieval. DocCVQA (Tito et al., 2021) is the first multi-modal document retrieval-answering task, which extracts information from a document image collection and then provides the answer. However, DocCVQA provides only 20 questions, and the corpora is limited to finance domain specifically in political disclosure. PDF-MVQA (Ding et al., 2024) is tailored for multi-modal retrieval in research articles. Note that PDF-MVQA is annotated by Chat-GPT (GPT-3.5-turbo) but not experts, and its corpora contains only biomedical articles. SciMMIR (Wu et al., 2024) proposes multi-modal retrieval in scientific research papers. However, it provides only the image-caption pairs, not the natural user query with the targeting document pages. Ma et al. (2024a) introduce two relevant datasets, namely Wiki-SS and DocMatix-IR. Wiki-SS is constructed based on Natural Questions (Kwiatkowski et al., 2019) where the evidence passage is the screenshot wikipedia webpage. Natural Questions are mainly designed for text rather than multi-modal retrieval. We also notice that the screenshot may not include the ground-truth evidence as only the front-page is screenshotted. DocMatix-IR is constructed from the largest document understanding dataset DocMatix (Laurençon et al., 2024) via filtering and hard negative mining. However, the VQA questions in DocMatix is constructed by Phi-3-small (Abdin et al., 2024) rather than experts, and these VQA questions are not de-contextualized for retrieval.
ViDoRe (Faysse et al., 2024) is the most relevant benchmark to MMDocIR. It integrates multiple DocVQA datasets (Mathew et al., 2021, 2022; Zhu et al., 2022; Li et al., 2024), and provides new documents in scientific, medical, administrative, and environment domains. Upon thorough examination of the 1,180 questions within the ViDoRe dataset, we find that more than 80% of these questions present drawbacks in terms of their complexity and context-specific nature. Typical examples include “What’s the title of the picture” and “what’s the author’s name”. They tend to be either overly simplistic or excessively dependent on a specific context which does not necessarily reflect the broader, more complex needs of effective IR. Additionally, a significant limitation of the ViDoRe dataset is its reliance on sampled document pages or cropped images rather than providing the complete document corpora.
8. Conclusion
In conclusion, multi-modal document retrieval presents a complex challenge that necessitates the integration of diverse data modalities beyond plain text. Our contributions address this challenge by introducing the MMDocIR benchmark, which includes innovative dual-task retrieval functionalities targeting both page-level and layout-level specifics of documents. The MMDocIR benchmark comprises a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 173,843 questions, making it a pivotal resource for advancing multi-modal document retrieval for both training and evaluation purposes. Our comprehensive evaluation reveals that visual-driven retrievers significantly outperform text-driven ones, underscoring the critical importance of incorporating visual content in enhancing the effectiveness of document retrieval systems. Future work can expand on optimizing these retrieval algorithms to improve the accuracy and efficiency of multi-modal document retrieval systems.

### v2

@@
-Benchmarking Multi-Modal Retrieval for Long Documents
-Abstract.
-Multi-modal document retrieval is designed to identify and retrieve various forms of multi-modal content, such as figures, tables, charts, and layout information from extensive documents.
-Despite its significance, there is a notable lack of a robust benchmark to effectively evaluate the performance of systems in multi-modal document retrieval.
-To address this gap, this work introduces a new benchmark, named as MMDocIR, encompassing two distinct tasks: page-level and layout-level retrieval.
-The former focuses on localizing the most relevant pages within a long document, while the latter targets the detection of specific layouts, offering a more fine-grained granularity than whole-page analysis.
-A layout can refer to a variety of elements such as textual paragraphs, equations, figures, tables, or charts.
-The MMDocIR benchmark comprises a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 173,843 questions, making it a pivotal resource for advancing multi-modal document retrieval for both training and evaluation.
-Through rigorous experiments, we reveal that (i) visual retrievers significantly outperform their text counterparts, (ii) MMDocIR train set can effectively benefit the training process of multi-modal document retrieval and (iii) text retrievers leveraging on VLM-text perform much better than those using OCR-text.
-These findings underscores the potential advantages of integrating visual elements for multi-modal document retrieval.
-MMDocIR is available at https://huggingface.co/MMDocIR.
-Multi-modal Document Retrieval, Page Retrieval, Layout Retrieval
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Multi-modal Document Retrieval
-Figure 1.
-MMDocIR comprises 313 lengthy documents across 10 different domains, along with 1,685 questions.
-For each question, page-level annotations are provided via selected screenshots.
-Red boundary boxes represent layout-level annotations.
-Figure 2.
-Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark (Ma et al., 2024b).
-Note that the white spaces, headers, and footers are removed from the area counting.
-1. Introduction
-Multi-modal document retrieval (Hassan et al., 2013; Ma et al., 2024a; Faysse et al., 2024; Lee et al., 2024) aims to encode and retrieve various forms of content from visually rich documents based on user queries.
-Unlike traditional document retrieval (Zhang et al., 2022; Chen et al., 2023; Sarthi et al., 2024; Dong et al., 2024) which primarily deals with textual data, multi-modal document retrieval must handle documents that seamlessly integrate text with multi-modal elements such as images, tables, charts, and layout designs.
-These multi-modal content often carry significant information that plain text cannot convey (Cui et al., 2021; Zhu et al., 2024; Sassioui et al., 2023), thereby enriching the depth and relevance of retrieved content.
-For example, our analysis of MMLongBench-Doc benchmark (Ma et al., 2024b) reveals that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively, as shown in Figure 2.
-This significant presence of multiple data modalities requires retriever to leverage all data modalities present in a document.
-However, current benchmarks (shown in Table 1) for evaluating multi-modal document retrievers are insufficient, lacking in certain aspects that are critical for a comprehensive assessment.
-The major shortcomings include: 1.
-Question Quality: The design and curation of questions in most benchmarks do not align with the specific needs of multi-modal document retrieval.
-The questions, often reused from Visual Question Answering (VQA) tasks, lack the specificity required for this purpose and require expert filtering to be truly effective for retrieval tasks.
+MMDocIR: Benchmarking Multimodal Retrieval for Long Documents
+Abstract
+Multimodal document retrieval aims to identify and retrieve various forms of multimodal content, such as figures, tables, charts, and layout information from extensive documents.
+Despite its increasing popularity, there is a notable lack of a comprehensive and robust benchmark to effectively evaluate the performance of systems in such tasks.
+To address this gap, this work introduces a new benchmark, named MMDocIR, that encompasses two distinct tasks: page-level and layout-level retrieval.
+The former evaluates the performance of identifying the most relevant pages within a long document, while the later assesses the ability of detecting specific layouts, providing a more fine-grained measure than whole-page analysis.
+A layout refers to a variety of elements, including textual paragraphs, equations, figures, tables, or charts.
+The MMDocIR benchmark comprises a rich dataset featuring 1,685 questions annotated by experts and 173,843 questions with bootstrapped labels, making it a valuable resource in multimodal document retrieval for both training and evaluation.
+Through rigorous experiments, we demonstrate that (i) visual retrievers significantly outperform their text counterparts, (ii) MMDocIR training set effectively enhances the performance of multimodal document retrieval and (iii) text retrievers leveraging VLM-text significantly outperforms retrievers relying on OCR-text.
+Our dataset is available at https://mmdocrag.github.io/MMDocIR/.
+MMDocIR: Benchmarking Multimodal Retrieval for Long Documents
+[…]
+Figure 1: MMDocIR evaluation set comprises 313 long documents and 1,658 queries across 10 domains.
+For each query, page-level labels are provided via selected screenshots.
+Red boundary boxes represent layout-level labels.
+$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally
+Figure 2: Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark.
+1 Introduction
+Benchmarks Question Document Label
+Type Expert? IR? #Num Evidence Type Domain #Pages Source Page Layout
+DocCVQA VQA question 20 TXT/L Finance 1.0
+SciMMIR Image caption 530k TAB/I Science 1.0
+ViDoRe VQA question 3,810 TXT/C/TAB/I Multiple 1.0
+PDF-MVQA Search query 260k TXT/TAB/I Biomedical 9.6
+MMLongBench-Doc VQA question 1,082 TXT/C/TAB/I Multiple 47.5
+Wiki-SS Natural question 3,610 TXT Wikipedia 1.0
+DocMatix-IR VQA question 5.61m TXT/C/TAB/I Multiple 4.2
+MMDocIR (eval) VQA question 1,658 TXT/C/TAB/I Multiple 65.1
+MMDocIR (train) VQA question 73.8k TXT/C/TAB/I Multiple 49.3
+Table 1: MMDocIR versus existing document IR datasets. TXT/C/TAB/I refers to text/chart/table/image.
+Multimodal document retrieval Hassan et al.
+(2013); Lee et al.
+(2024) aims to retrieve information from visually rich documents based on user queries.
+Unlike traditional document retrieval Zhang et al.
+(2022); Chen et al.
+(2023) which primarily deals with textual data, multimodal document retrieval imposes substantially greater demands on understanding multimodal elements such as images, tables, charts, and layout designs.
+Such elements often carry significant information that plain text fails to convey Cui et al.
+(2021); Sassioui et al.
+(2023): tables reveal structured data patterns, charts visualize trends or correlations, images offer contextual and semantic cues, etc.
+Combining these visual elements enriches the quality of retrieved content.
+Our analysis of MMLongBench-Doc benchmark Ma et al.
+(2024b) in Figure 2 shows that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively.
+This highlights the need for retrieval systems that effectively handle multimodal information.
+However, as shown on (see Table 1), existing benchmarks exhibit several critical limitations that undermine comprehensive evaluation of multimodal retrieval systems.
+The key limitations include: 1.
+Question Quality: Many questions used in existing benchmarks are directly sourced from datasets for Visual Question Answering (VQA) tasks.
+Some questions often assume the input is already relevant, making it not suited for meaningful evaluation of retrieval capabilities.
 2.
-Document Quality: A significant limitation of existing benchmarks is the missing of complete document pages, thus hindering accurate evaluation of retrieval capabilities within actual document contexts.
-Additionally, the range of document domains covered is often too narrow, limiting the benchmarks’ applicability across different fields.
+Document Completeness and Diversity: Existing benchmarks often provide only partial documents, limiting the ability to evaluate within full document context.
+Additionally, the narrow range of document domains further restricts their applicability across diverse use-cases in real-world.
 3.
-Retrieval Granularity: Most benchmarks, except for SciMMIR (Wu et al., 2024), only allow retrieval at a page level which is insufficient, as user queries may pertain to specific parts of a page like a particular figure or table rather than the entire page.
-Recognizing these gaps, we introduce MMDocIR Multi-Modal Document Information Retrieval benchmark.
-MMDocIR is innovatively structured around two critical tasks: page-level and layout-level retrieval.
-(1) The page-level retrieval task is designed to identify the most relevant pages within a document in response to a user query.
-(2) The layout-level retrieval aims to retrieve most relevant layouts.
-The layouts are defined as the fine-grained elements such as paragraphs, equations, figures, tables, and charts.
-This task allows for a more nuanced content retrieval, honing in on specific information that directly answers user queries.
-To support these tasks, we first develop MMDocIR evaluation set comprising 313 documents averaging 65.1 pages each, along with VQA 1,658 questions derived from MMLongBench-Doc (Ma et al., 2024b) and DocBench (Zou et al., 2024).
-Initially, we examine all VQA questions, filter out questions that are irrelevant to IR tasks, and revise remaining questions to be more suitable for IR.
-Then, we manually annotate these questions with page and layout labels.
-The page-level labels are the specific pages identified that contain answer evidence.
-[^1] It is worth noting that while MMLongBench-Doc provided initial page labels, our meticulous review lead to corrections in about 21.3% of these labels.
-The layout-level labels are the bounding boxes meticulously delineated around the key evidence within the identified pages.
-Meanwhile, we introduce MMDocIR training set, comprising 73,843 QA pairs converted from 7 DocQA datasets.
-Specifically, we manually collect 6,878 documents and use either manual or automatic methods to annotate the ground truth labels.
-[^2] Since layout annotations are missing for most datasets, our data annotating pipeline can only create layout-level labels for 4 subsets.
-Benchmarks Question Document Label
-Type By Expert? For IR? #Num Evidence Type Domain #Pages Source Page Layout
-DocCVQA (Tito et al., 2021) VQA question 20 TXT/L Finance 1.0
-SciMMIR (Wu et al., 2024) Image caption 530K TAB/I Science 1.0
-ViDoRe (Faysse et al., 2024) VQA question 3,810 TXT/C/TAB/I Multi-domain 1.0
-PDF-MVQA (Ding et al., 2024) Search query 260k TXT/TAB/I Biomedical 9.6
-MMLongBench-Doc (Ma et al., 2024b) VQA question 1,082 TXT/L/C/TAB/I Multi-domain 47.5
-Wiki-SS (Ma et al., 2024a) Natural question 3,610 TXT Wikipedia 1.0
-DocMatix-IR (Ma et al., 2024a) VQA question 5.61M TXT/L/C/TAB/I Multi-domain 4.2
-MMDocIR VQA question 1,658 TXT/C/TAB/I Multi-domain 65.1
-Table 1. Comparison between our benchmark and previous Document IR datasets. TXT/L/C/TAB/I: pure text/generalized layout/chart/table/image.
-In our research, we conduct a thorough evaluation of the prevailing multi-modal document retrieval baselines which can broadly be categorized into visual-driven and text-driven retrievers.
-Visual-driven retrievers (Ma et al., 2024a; Faysse et al., 2024), leverage vision-language models (VLMs) (Abdin et al., 2024; Beyer et al., 2024; Bai et al., 2023; Chen et al., 2024b) to capture rich multi-modal cues and generate embeddings for both queries and documents.
-In contrast, text-driven retrievers (Karpukhin et al., 2020; Khattab and Zaharia, 2020; Izacard et al., 2022; Wang et al., 2022; Xiao et al., 2023; Li et al., 2023) primarily rely on OCR or VLM to first convert the multi-modal content into text, subsequently employing language models (LMs)  (Devlin et al., 2019; Liu et al., 2019; Wang et al., 2024b) to generate embeddings for both queries and documents.
-Via extensive experiments, we discover that visual-driven retrievers consistently outperform their text-driven counterparts, often by a significant margin.
-This highlights the critical role that visual elements play in multi-modal document retrieval.
-In summary, our contributions are three-fold:
-  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework (§$\S$ 2) that encompasses page-level and fine-grained layout-level multi-modal document retrieval.
-  • MMDocIR Benchmark: We introduce Multi-Modal Document Information Retrieval (MMDocIR) benchmark.
-The evaluation set (§$\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 QA pairs.
-The training set (§$\S$ 4) consists of 6,878 documents and labels for 73,843 QA pairs.
-  • Comprehensive Evaluation: We conduct an extensive evaluation of existing retrieval systems (§$\S$ 6).
-Our results reveal that visual-driven retrievers consistently outperform their text-driven counterparts, highlighting the importance of utilizing visual information in multi-modal document retrieval.
-2. Dual-Task Retrieval Definition
-Let $\mathcal{D}$ be a document corpora consisting of a set of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection [^3] A document page usually comprises about 5 to 15 layouts, depending on its complexity..
-The objective is to perform document page-level and layout-level retrieval.
-Specifically, the task involves retrieving a subset of these pages and layouts that are most relevant to a query $Q$, identified by the top $k$ entries where $k$ is significantly smaller than $n$ or $m$.
-The relevance of pages ($p$) and layouts ($l$) to $Q$ is quantified by similarity scores, $\textrm{Sim}(Q,p)$ and $\textrm{Sim}(Q,l)$ respectively.
-The retrieval system can be decomposed into (1) an offline indexing phase in which each page and layout from $\mathcal{P}$ and $\mathcal{L}$ is transformed into a vector representation, and (2) an online querying phase in which a query $Q$ is converted into a vector form, and the vectorized query is then compared against the indexed vectors using the similarity scores $\textrm{Sim}(Q,p)$ for pages and $\textrm{Sim}(Q,l)$ for layouts.
-3. MMDocIR: Evaluation Set
-{tabu}
-l—c@ c@ c@ c@ c@ c@ c—c@   c@  c@   c@   —c@   c@  c@   c@
-Dataset Domain Document Statistics   Questions (%)   Modality Distribution (%)
+Retrieval Granularity: Most benchmarks support only page-level retrieval.
+Such granularity is often insufficient, as user queries frequently target specific elements, such as figures or tables, rather than entire pages.
+To address these gaps, we introduce MMDocIR, a multimodal document information retrieval benchmark.
+MMDocIR is designed for two key tasks: page-level and layout-level retrieval.
+(1) The page-level retrieval identifies the most relevant pages within a document to answer user query.
+(2) The layout-level retrieval targets the most relevant fine-grained elements such as paragraphs, equations, figures, tables, and charts (see Appendix E.2 for more examples).
+Such task supports more precise and context-aware retrieval that pinpoint specific elements to address user queries.
+To support both tasks, we develop MMDocIR evaluation set that comprises 313 documents, each averaging 65.1 pages, along with 1,658 modified queries derived from MMLongBench-Doc and DocBench Zou et al.
+(2024).
+The queries are annotated with 2,107 page-level and 2,638 layout-level labels.
+The page labels are specific pages that contain the evidence needed to answer the query.
+[^1] While MMLongBench-Doc provided initial page labels, our meticulous review lead to corrections in 21.3% of them.
+The layout labels consist of precisely drawn bounding boxes around the key evidence within the identified pages.
+In addition, we introduce the MMDocIR training set, designed to support retriever training.
+It contains 73,843 questions sourced from 7 DocQA datasets.
+To construct this set, we manually collect 6,878 documents and apply a semi-automatic pipeline to annotate the ground truth labels.
+By leveraging MMDocIR, we conduct a comprehensive evaluation on multimodal document retrieval across two retriever types: visual-driven and text-driven.
+Visual-driven retrievers Ma et al.
+(2024a); Faysse et al.
+(2024), leverage vision language models (VLMs) to capture rich multimodal cues and generate embeddings for both queries and documents.
+In contrast, text-driven retrievers Karpukhin et al.
+(2020); Khattab and Zaharia (2020); Xiao et al.
+(2023) rely on OCR or VLM to first convert the multimodal content into text, subsequently employing language models (LMs) to generate embeddings for both queries and documents.
+Our extensive experiments reveal that visual-driven retrievers consistently outperform their text-driven counterparts, often by a significant margin.
+In summary, our contributions are threefold:
+  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework (§$\S$ 2) that supports page-level and fine-grained layout-level multimodal document retrieval.
+  • MMDocIR Benchmark: We introduce Multimodal Document Information Retrieval benchmark.
+The evaluation set (§$\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 questions.
+The training set (§$\S$ 4) consists of 6,878 documents and labels for 73,843 questions.
+  • We conduct extensive experiments and comparisons of both text and visual retrievers (§$\S$ 5), demonstrating clear advantage of incorporating visual content in multimodal retrieval tasks.
+2 Dual-Task Retrieval Definition
+Let $\mathcal{D}$ be a document corpora consisting of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection.
+The objective is to perform document retrieval at both page-level and layout-level.
+Specifically, given query $Q$, the task is to retrieve the top $k$ pages and layouts most relevant to $Q$, where $k<<n$ and $k<<m$.
+The relevance of pages ($p$) and layouts ($l$) to $Q$ is measured by similarity scores, $\textrm{Sim}(Q,p)$ and $\textrm{Sim}(Q,l)$ respectively.
+The retrieval system consists of two phases: (1) an offline indexing phase, where pages and layouts from $\mathcal{P}$ and $\mathcal{L}$ are encoded into vectors, and (2) an online querying phase, in which a query $Q$ is encoded into a vector, which is then compared against the offline-indexed vectors using similarity scores $\textrm{Sim}(Q,p)$ for pages and $\textrm{Sim}(Q,l)$ for layouts.
+3 MMDocIR: Evaluation Set
+3.1 Document Corpora Collection
+After a comprehensive review of existing DocVQA datasets, we select MMLongBench-Doc Ma et al.
+(2024b) and DocBench Zou et al.
+(2024) to facilitate our benchmark construction (see Appendix B.2 for our selection criteria).
+MMLongBench-Doc is a long-context, multimodal benchmark comprising 1,091 questions across 135 documents with 47.5 pages on average.
+DocBench emphasizes long document understanding, consisting of 1,102 questions across 229 documents, each with an average length of 77.5 pages.
+Both datasets offer corpora from diverse domains with expert-annotated questions that require evidence from various modalities.
+Consequently, we curate a set of 364 documents and 2,193 questions for our subsequent annotation.
+3.2 Annotation Process
+Question Filtering and Revision.
+To ensure that the questions in MMDocIR are optimally suited for document retrieval tasks, we identify four specific types of questions (see Appendix B.3) that do not align well with the objectives of IR.
+By filtering and refining these questions, we ensure the integrity and relevance of MMDocIR, resulting in 1,658 questions for subsequent annotation.
+Page-level Annotation.
+We annotate page labels that precisely identify the exact pages containing ground truth evidence.
+Given that documents in MMDocIR contain 65.1 pages on average, pinpointing relevant pages is highly non-trivial, akin to finding a needle in haystack, which demands careful inspection and document understanding.
+  • For DocBench: we manually annotate page labels for all 864 questions from scratch, by carefully reviewing each document and locating the pages containing answer evidence.
+  • For MMLongBench-Doc: we rigorously review and validate the answers and page labels of 794 questions.
+This effort results in corrections to 10 answers and 169 page labels [^2] Common errors in page labeling: annotators starting page indexing at 1 rather than 0, missing labels for questions spanning multiple pages, and incorrect or absent page labels..
+Layout-level Annotation.
+To enhance the granularity of our benchmark, we extend our annotations to include layout-level labels, identifying specific layout elements as evidence.
+Compared to page annotation, layout-level labeling is significantly more complex and labor-intensive.
+  • Layout Detection. We begin by utilizing MinerU Wang et al. (2024) to automatically parse all documents and detect all layouts (e.g., layout type and bounding boxes).
+  • Evidence Identification.
+We identify the layouts that contain necessary answer evidence.
+In case where MinerU fails to detect evidentiary element, we manually annotate the the bounding boxes, accounting for 7% of the total layout-level labels.
+3.3 Quality Control
+Consistency Page Labels Layout Labels
+Prec. Recall F1 Prec. Recall F1
+[…]
+[…]
+Average 95.0 95.4 95.2 87.0 87.2 87.1
+Table 2: Annotation consistency between group A & B.
+To ensure annotation quality and reliability in MMDocIR, we have adopted a rigorous 3-stage quality control process.
+We split questions into two parts.
+Each group is responsible for annotating approximately 1,000 questions, with an overlap of 400 questions serving the need for cross-validation.
+[…]
+  • Cross-Evaluation: We cross-evaluate and achieve F1 score of 95.2 and 87.1 for page and layout labels, as shown in Table 2. We then identify and fix the discrepancies.
+  • Random Cross-Validation: We randomly cross-validate 50% of the remaining annotations. In the cases where we have different opinions, we discuss to achieve mutually-agreed annotations.
+(a) #words/layout
+(b) Distribution density: #words/layout
+(c) Distribution density: #words/page
+Figure 3: Distribution of OCR/VLM-text by length.
+3.4 Multimodal content as OCR/VLM-text
+To apply multimodal retrieval to text retrievers, we convert multimodal layouts (e.g., tables or figures) into text.
+Specifically, we extract text using OCR Smith (2007) (“OCR-text”) and generate detailed descriptions using VLMs OpenAI (2024); Qwen-Team (2024) (“VLM-text”).
+As a result, each image layout is represented in three formats: original image, OCR-text, and VLM-text.
+For layouts, the average word length and distribution of OCR-text and VLM-text of MMDocIR are shown in Figure 3(a) and 3(b).
+Notably, the length of VLM-text is 1.5 and 3.8 times of OCR-text for table (with more structured numbers) and figure (mostly with visual elaboration), respectively.
+For pages, we construct two variants by combining the natural text with either OCR-text and VLM-text for each page, resulting in OCR-page and VLM-page representations.
+The average word length is 477 and 505 for OCR-page and VLM-page respectively, with their distribution shown in Figure 3(c).
+MMDocIR Document Statistics Questions (%) Modality Distribution %
 #Doc #QA #Page #Lay #Page #Lay %Lay Text Image Table Lay/ Text Image Table Title
 Label Label /Doc /Page Meta
-MMDocIR Eval 313 1,658 2,107 2,638 65.1 8.3 41.8 44.7 21.7 37.4 11.5 60.4 18.8 16.7 4.1
-[…]
+Eval Domains 313 1,658 2,107 2,638 65.1 8.3 41.8 44.7 21.7 37.4 11.5 60.4 18.8 16.7 4.1
 - Research Report 34 200 318 400 39.4 6.0 39.1 45.0 17.5 74.5 13.5 45.6 40.0 5.9 8.4
@@
 - News 1 137 137 170 50.0 73.6 72.3 70.1 1.5 0 28.5 48.5 39.8 0.0 11.6
-Table 2.
-Detailed statistics for MMDocIR evaluation set.
-“#Lay/Page” is the averaging layouts per page, reflecting page’s layout complexity.
-“%Lay” refers to the area ratio of useful layouts (excluding white spaces, headers, and footers) over entire page.
-3.1. Document Corpora Collection
-To facilitate the development of MMDocIR, we leverage resources from existing Document Visual Question Answering (DocVQA) benchmarks (Mathew et al., 2021; Tito et al., 2023; Tanaka et al., 2023; Zhu et al., 2022; Li et al., 2024; Landeghem et al., 2023; Ma et al., 2024b; Zou et al., 2024).
-Although DocVQA benchmarks are not primarily designed for IR, they offer valuable document corpora and questions that can be adapted for our purposes.
-To select useful DocVQA datasets, we adhere to following criteria:
-  • Document Source: The dataset must include accessible original documents or sources for these documents. We need to access and enrich them to support more complex retrieval tasks.
-  • Diverse Domain/Modality: The document collections must (1) encompass diverse domains e.g., academia, finance, publishing, and research, and (2) embrace multiple modalities, such as text, figures, tables, charts, and layouts.
-  • Long Document: We seek documents of sufficient length as longer texts pose more significant challenges. This criterion can evaluate models in handling complex and lengthy documents.
-  • QA Diversity and Comprehensiveness: The questions included in the dataset should be diverse and challenging.
-For example, some questions should require reasoning across both text and visual tables/figures, while some are multi-hop questions that require synthesizing information from multiple pages.
-After considering aforementioned criteria, we choose to leverage on the document corpora and questions from MMLongBench-Doc (Ma et al., 2024b) and DocBench (Zou et al., 2024).
-MMLongBench-Doc is a long-context, multi-modal benchmark comprising 1,091 questions constructed upon 135 documents averaging 47.5 pages.
-Its corpora include domains of research report, administration and industry, tutorial and workshop, academic paper, brochure, financial report, and guidebook.
-DocBench focuses on long document understanding, consisting of 1,102 questions constructed upon 229 documents averaging 77.5 pages.
-The corpora span across domains: academia, finance, government, laws, and news.
-Both datasets provide expert-annotated questions, requiring content evidence of various modalities (e.g., text, tables, charts and images, as well as locations layouts) to answer.
-Consequently, we obtain a total of 364 documents and 2,193 questions for our subsequent annotation.
-3.2. Annotation Process
-Question Filtering and Revision
-To ensure that the questions in MMDocIR are optimally suited for document retrieval tasks.
-To this end, we identify four specific types of questions that do not align well with the objectives of IR.
-By filtering out or modifying these questions, we ensure the integrity and relevance of MMDocIR.
-  • Summarization Questions: These questions, such as “What does this book mainly illustrate?” require a comprehensive understanding of large sections of a document or even the entire document.
-The broad scope makes it hard to pinpoint specific content effectively, which contradicts the precise nature of IR tasks.
-  • Overwhelm Statistical Questions: Questions that demand extensive data computation or collation, such as “How many words are there in total in the paper?” also fall outside our scope.
-These require a level of statistical analysis that goes beyond retrieval.
-  • Online Search Questions: For example, queries like “What is the Google Scholar citation count of the main author of this article?” rely on information from external online resources.
-Since our focus is on retrieving information that is internally available within the documents, these questions are excluded.
-  • Unanswerable Questions: These are designed to test if models generate answers based on non-existent information (model hallucinations).
-Since they do not facilitate the retrieval of factual document-based information, these questions are excluded.
-Page-level annotation process
-To address the challenge of accurately locating relevant information within extensive documents, we annotate page-level labels that pinpoint the exact pages containing evidence needed to answer specific questions.
-Considering that the average document in our dataset contains 65.1 pages, pinpointing the correct pages is similar to finding a needle in a haystack and requires meticulous effort.
-Our annotation process is:
-  • For MMLongBench-Doc: we review and correct the answers and corresponding page labels for 794 questions.
-Specifically, we make corrections to 10 answers and revise 169 page labels [^4] Common errors in page labeling: annotators starting page indexing at 1, missing labels for questions spanning multiple pages, and incorrect or absent page labels.
-to ensure they accurately reflect the evidence locations.
-  • For DocBench: we annotate page labels for all 864 questions from scratch.
-This involves a detailed examination of each document to accurately assign page labels that match the evidential requirements of each question.
-Through these efforts and following a rigorous annotation and cross-validation process, we successfully compile page-level labels for a total of 1,658 questions in MMDocIR.
-Layout-level annotation process
-To achieve more fine-grained retrieval granularity, we go beyond page-level annotations by including layout-level labels.
-This advanced annotation process captures the specific layout elements within pages that contain necessary evidence to answer each question.
-Unlike page-level annotation, labeling layout-level information is more complex and labor-intensive due to its detailed nature.
-Our annotations are via:
-  • To streamline annotation process, we utilize MinerU (Wang et al., 2024a) to parse all documents and extract the layout information (e.g., type and bounding boxes) for every layout element on a page.
-These elements are categorized into five main types: text, image, table, title, and equation.
-The bounding boxes are delineated by the coordinates of the top-right and bottom-left corners.
-  • Based on the layouts drawn by MinerU, we identify the layouts that contain necessary eivdence to answer the question.
-In instances where MinerU fails to detect certain evidentiary elements, we manually draws the bounding boxes to ensure the precision of layout labels.
-This manual intervention is necessary for approximately 7% of our layout-level labels.
-Ultimately, this meticulous process leads to the annotation of 2,638 layout labels for the 1,658 questions in MMDocIR.
-3.3. Quality Control
-Consistency Page Labels Layout Labels
-Prec. Recall F1 Prec. Recall F1
-[…]
-[…]
-Average 95.0 95.4 95.2 87.0 87.2 87.1
-[…]
-{tabu}
-l—c—ccc@ c@ c—c@   c@  c@   c@   —c@   c@
-Dataset Domain Document Statistics   Modality Distribution (%)   Labels
+Table 3: Detailed statistics of MMDocIR evaluation set.
+“#Lay/Page” is averaging layouts per page, reflecting layout complexity.
+“%Lay” is the area ratio of useful layouts (excluding white spaces, headers, and footers).
+MMDocIR Domain Document Statistics Evidence Modality (%) Labels
 #Doc #QA #Page #Lay %Lay Text Image Table Title Page Lay
 /Doc /Page
-MMDocIR Train assorted docs 6,878 73,843 32.6 6.32 42.6 49.3 34.3 10.8 4.9
+Train Subsets assorted docs 6,878 73,843 32.6 6.32 42.6 49.3 34.3 10.8 4.9
 […]
-- MP-DocVQA (Tito et al., 2023) health/ind. docs 875 15,266 46.8 6.9 38.8 57.3 18.0 22.7 1.9
-- SlideVQA (Tanaka et al., 2023) diverse slides 2,011 11,066 49.3 4.4 42.3 30.1 56.2 4.7 8.8
-- TAT-DQA (Zhu et al., 2022) annual reports 163 15,814 147.3 9.2 42.2 66.4 4.4 26.5 2.7
-- ArXivQA (Li et al., 2024) arXiv papers 1,579 12,314 18.4 7.9 50.0 70.4 22.3 2.8 1.0
-- SciQAG (Wan et al., 2024) science papers 1,197 4,976 9.0 9.1 53.7 61.8 28.0 6.7 1.5
-- DUDE (Landeghem et al., 2023) assorted docs 779 3,173 15.6 7.4 42.5 57.1 24.7 15.2 2.9
-- CUAD (Hendrycks et al., 2021) legal contract 274 11,234 29.6 7.4 24.7 89.3 2.5 6.4 1.1
-Table 4. Document statistics for Training Datasets collected.
-To ensure annotation quality and reliability in MMDocIR, we have adopted a rigorous quality control process using a cross-validation method.
-This method involves splitting the document corpora into two main groups, labeled as Group A and Group B.
-Each group is responsible for annotating page-level and layout-level labels for approximately 1,000 QA pairs.
-Importantly, there is an overlap of 400 QA pairs between two groups, which serves as a validation checkpoint.
-The process for quality control works as follows:
-  • Overlap Scoring: For 400 overlapping questions, each group’s annotations are scored against the other’s as a form of mutual validation.
-Specifically, Group A’s annotations are used as a ground truth to score Group B’s work, and conversely, Group B’s annotations serve as the ground truth for scoring Group A.
-  • Cross-Evaluation: By cross-evaluating the outcomes from both groups, we are able to assess the consistency of the annotations and identify any discrepancies for improvement.
-Specifically, we achieved an F1 score of 95.2 for page-level annotations and 87.1 for layout-level annotations, as shown in Table 3.
-  • Random Cross-Validation: Apart from the overlapping 400 questions, we randomly cross-validate around 50% of the remaining annotations.
-In the cases where we have different opinions, we discuss to achieve mutually-agreed annotations.
-This step ensures the high quality and reliability of MMDocIR.
-3.4. Statistics and Analysis
+- MP-DocVQA health/ind. docs 875 15,266 46.8 6.9 38.8 57.3 18.0 22.7 1.9
+- SlideVQA diverse slides 2,011 11,066 49.3 4.4 42.3 30.1 56.2 4.7 8.8
+- TAT-DQA annual reports 163 15,814 147.3 9.2 42.2 66.4 4.4 26.5 2.7
+- arXivQA arXiv papers 1,579 12,314 18.4 7.9 50.0 70.4 22.3 2.8 1.0
+- SciQAG science papers 1,197 4,976 9.0 9.1 53.7 61.8 28.0 6.7 1.5
+- DUDE assorted docs 779 3,173 15.6 7.4 42.5 57.1 24.7 15.2 2.9
+- CUAD legal contracts 274 11,234 29.6 7.4 24.7 89.3 2.5 6.4 1.1
+Table 4: MMDocIR training set statistics about our collected documents, questions, and constructed labels.
+3.5 Statistics and Analysis
 Document Analysis.
-As shown in Table 3, MMDocIR evaluation set includes 313 long documents averaging 65.1 pages, categorized into ten main domains.
-Different domains feature distinct distributions of multi-modal information.
-For instance, research reports, tutorials, workshops, and brochures predominantly contain images, whereas financial and industry documents are table-rich.
-In contrast, government and legal documents primarily comprise text.
-Overall, the modality distribution is: Text (60.4%), Image (18.8%), Table (16.7%), and other modalities (4.1%).
+As shown in Table 3, MMDocIR evaluation set includes 313 long documents, averaging 65.1 pages, categorized into 10 domains.
+Different domains feature distinct multimodal distribution.
+The overall modality distribution is as follows: text (60.4%), image (18.8%), table (16.7%), and others (4.1%), with fine-grained distribution shown in Figure 4(a).
 Question and Annotation Analysis.
-MMDocIR encompasses 1,658 questions, 2,107 page labels, and 2,638 layout labels.
-The modalities required to answer these questions distribute across four categories: Text (44.7%), Image (21.7%), Table (37.4%), and Layout/Meta (11.5%).
-The “Layout/Meta” category encompasses questions related to layout information and meta-data statistics.
-Notably, the dataset poses several challenges: 254 questions necessitate cross-modal understanding, 313 questions demand evidence across multiple pages, and 637 questions require reasoning based on multiple layouts.
-These complexities highlight the need for advanced multi-modal reasoning and contextual understanding.
-4. MMDocIR: Training Set
-4.1. Document Corpus Collection
-The training set corpora are mainly collected from DocVQA benchmarks, adhering to similar selection criteria (see Section 3.1).
-  • MP-DocVQA (Tito et al., 2023) contains 47,952 images collected from Industry Documents Library (IDL)  [^5] https://www.industrydocuments.ucsf.edu/.
-We group the 47,952 document images into separate document files, and obtain 875 long documents (46.8 pages on average) with 15,266 QA pairs.
-  • SlideVQA (Tanaka et al., 2023) contains 2,619 slide documents collected from slideshare  [^6] https://www.slideshare.net/ and covering 39 topics.
-Note that SlideVQA contains only the first 20 pages of each slide decks.
-In our research, we manually collect the complete slide decks, and obtain 2,011 long documents (averaging 49.3 pages) with 11,066 QA pairs.
-  • TAT-DQA (Zhu et al., 2022) consists of 3,067 document pages from financial reports  [^7] https://www.annualreports.com/, dated between 2018 and 2020.
-Note that neither original documents nor links is provided.
-We use OCR to extract text in the pages, and use search engine to find relevant documents.
-After careful tracing and recognition, we identify 163 original documents (averaging 147.3 pages) with 15,814 QA pairs.
-  • ArXivQA (Li et al., 2024) comprises 32k figures cropped from academic pages  [^8] https://arxiv.org/.
-We use the arXiv DOIs provided to collect the academic papers.
-After careful tracing, recognition, and document length filtering, we identify 1,579 documents averaging 18.4 pages.
-  • SciQAG (Wan et al., 2024) consists of 22,728 papers in 24 scientific disciplines, collected from Web of Science (WoS) Core Collection database.
-We sample 50 documents from each discipline, and manually collect 1,197 papers using the DOIs provided.
-  • DUDE (Landeghem et al., 2023) provides 5,019 documents from aggregate websites9991: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/.
-It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires.
-We filter out short documents and obtain 779 relatively long documents (averaging 15.6 pages) with 3,173 QA pairs.
-  • CUAD (Hendrycks et al., 2021) provides 510 commercial legal contracts, collected from Electronic Data Gathering, Analysis, and Retrieval (EDGAR) [^10] https://www.sec.gov/search-filings.
-We filter out short documents in CUAD and obtain 274 long documents (29.6 pages on average) with 11,234 QA pairs.
-4.2. Label Construction
-The page labels can be converted for MP-DocVQA, SlideVQA, and DUDE datasets. Among them, only DUDE provides layout labels.
-SciQAG provides only question and answer in texts.
-We utilize them to reversely obtain the page-level and layout-level labels.
-Specifically, we first use MinerU to obtain layout-level passage chunks.
-For each QA pair, we deploy E5 and BGE retrievers to obtain question-passage and answer-passage similarity scores against all extracted passage chunks.
-If both scores rank within top 3 for a certain passage chunk, we confirm this layout to be the layout-level labels for the given QA pair.
-Similarly, ArXivQA provides only cropped images, without document page/layout labels.
-We first use MinerU to obtain layout-level image.
-For each cropped image, we calculate the image similarity between all extracted images, and keep only the highest scoring one.
-Subsequently, we manually examine if the selected image matches the cropped image.
-In this way, we filter around 20% unmatched images and their QA pairs, resulting 1,579 QA pairs with page-level and layout-level labels.
-For TAT-DQA, the layout-level labels are provided for sampled single page.
-We need to localize the page index of the sampled page.
-We utilize PDF mapping technique to retrieve the best matched page for every sampled page in this document.
-Then, we manual examine if the retrieved page is identical to the given page, and correct the labels if there were any errors.
-The overall statistics (e.g., document information, modality distribution, domain, etc) of MMDocIR training set are in Table 3.3.
+MMDocIR includes 1,658 questions, and 2,107 page and 2,638 layout labels.
+The evidence spans 4 modalities: text (44.7%), image (21.7%), table (37.4%), and layout/meta (11.5%).
+Notably, MMDocIR presents several challenges: 254 questions require cross-modal understanding, 313 questions require evidence across multiple pages, and 637 questions require reasoning over multiple layouts.
+4 MMDocIR: Training Set
+4.1 Document Corpus Collection
+After screening related DocVQA datasets, we collect our training set corpora from 7 datasets, namely MP-DocVQA Tito et al.
+(2023), SlideVQA Tanaka et al.
+(2023), TAT-DQA Zhu et al.
+(2022), SciQAG Wan et al.
+(2024), DUDE Landeghem et al.
+(2023), and CUAD Hendrycks et al.
+(2021).
+Since most of these datasets do not provide original document, we invest significant efforts in tracing and recovering the original document, as detailed in Appendix B.4.
+4.2 Label Construction and statistics
+We use semi-automated construction pipeline to generate page-level and layout-level labels for datasets that lack them, referring to Appendix B.5 and B.6 for more details of construction process.
+Notably, layout annotations are missing from most existing datasets, as we manage to obtain or construct layout-level labels for only 4 datasets.
+The overall statistics (e.g., document information, modality distribution, domain, etc) of MMDocIR training set are summarized in Table 4.
 {tabu}
-lll—cccccccccc—cc
+lll|cccccccccc|cc
+Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
+Report &Indu. Worksh. Paper ure Report book ment Macro Micro
+Recall@$k=1$
+VLM-text
+DPR 32.3 25.5 27.0 31.0 28.4 18.8 23.5 31.2 38.3 16.1 27.2 26.9
+ColBERT 48.6 42.8 51.1 46.2 36.0 36.8 49.6 60.9 59.5 26.3 45.8 44.9
+BGE 48.8 30.9 47.1 40.8 37.6 28.4 43.4 51.9 48.9 28.5 40.6 39.6
+E5 48.1 30.0 50.4 39.4 41.1 29.7 40.9 52.8 51.1 24.1 40.8 39.5
+Contriever 45.5 31.2 49.8 41.5 39.4 29.4 45.2 55.3 51.1 20.4 40.9 39.7
+GTE 46.5 26.3 48.7 38.9 35.9 27.0 46.2 50.1 45.8 24.1 38.9 37.9
+[…]
+Image
+DSEwiki-ss 53.0 50.0 54.0 48.7 45.1 43.0 51.5 46.9 54.2 33.6 48.0 47.5
+DSEdocmatix 52.3 40.4 56.1 51.7 45.8 43.5 53.8 53.7 58.3 46.7 50.2 50.1
+ColPali 56.0 51.8 58.6 55.9 52.0 47.2 57.9 53.9 64.0 32.8 53.0 52.7
+DPR-Phi3ours 58.9 50.4 57.4 59.0 57.3 44.6 63.8 50.5 64.4 35.0 54.1 53.7
+Col-Phi3ours 56.7 50.4 56.9 61.3 54.8 50.7 60.8 61.3 63.6 54.0 57.0 57.1
+Recall@$k=3$
+VLM-text
+DPR 52.2 44.2 43.5 54.6 52.0 35.1 44.4 53.9 57.2 25.5 46.3 46.2
+ColBERT 70.1 64.4 70.3 72.3 59.1 55.3 71.1 81.3 70.8 34.3 64.9 64.8
+BGE 71.5 48.2 68.8 65.7 56.2 46.5 66.1 69.9 72.0 32.1 59.7 59.6
+E5 68.4 45.7 68.1 63.7 60.1 44.0 69.3 72.3 78.8 32.8 60.3 59.3
+Contriever 69.4 55.3 68.3 64.9 56.9 46.2 69.9 71.1 72.0 32.1 60.6 59.7
+GTE 71.1 44.5 67.2 64.4 54.3 43.0 70.6 71.9 68.2 31.4 58.7 58.3
+[…]
+Image
+DSEwiki-ss 75.4 65.0 73.9 79.8 69.5 63.5 75.4 71.5 81.4 50.4 70.6 71.4
+DSEdocmatix 75.4 67.5 73.3 80.0 66.3 61.6 72.8 76.4 82.6 57.7 71.4 71.8
+ColPali 77.6 71.8 79.4 83.4 72.6 66.1 80.0 80.4 86.4 49.6 74.7 75.0
+DPR-Phi3ours 80.3 66.5 77.6 83.9 71.9 63.8 79.8 71.4 84.5 55.5 73.5 74.3
+Col-Phi3ours 80.2 74.1 77.4 84.8 69.1 67.7 78.7 79.5 81.8 69.3 76.3 76.8
+Recall@$k=5$
+VLM-text
+DPR 66.5 60.1 56.0 68.9 58.8 43.8 57.1 68.6 64.8 33.6 57.8 57.8
+ColBERT 78.8 74.0 78.7 82.3 66.1 60.8 77.0 88.5 78.0 38.7 72.3 72.3
+BGE 79.5 65.8 71.3 76.8 62.4 56.0 77.2 77.4 79.5 38.0 68.4 68.5
+E5 76.9 64.2 75.3 74.4 67.4 52.0 78.5 78.6 82.6 40.9 69.1 67.9
+Contriever 77.2 67.1 76.7 75.2 65.1 53.7 75.4 79.2 83.3 39.4 69.2 68.3
+GTE 77.4 62.6 74.7 75.8 62.0 51.8 77.8 80.0 75.0 39.4 67.6 67.2
+[…]
+Image
+DSEwiki-ss 84.0 80.2 78.7 87.0 75.7 73.0 82.0 77.3 88.3 58.4 78.5 79.2
+DSEdocmatix 82.1 77.2 79.6 87.8 73.9 72.4 81.7 83.1 89.4 67.9 79.5 80.1
+ColPali 84.6 79.3 82.3 89.0 79.8 72.1 86.7 84.9 92.4 56.9 80.8 81.0
+DPR-Phi3ours 86.9 76.2 85.3 91.9 80.0 71.2 87.1 79.5 92.0 61.3 81.1 81.8
+Col-Phi3ours 86.3 78.8 81.2 92.4 79.0 73.8 85.3 85.1 87.1 73.0 82.2 83.0
+Table 5: Main results for page-level retrieval, with the best results in boldface and second best results underlined. For clarity, we omit results using VLM-text (Refer to Table A for full results).
+5 Experiment
+{tabu}
+lll|cccccccccc|cc
+Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
+Report &Indu. Worksh. Paper ure Report book ment Macro Micro
+Recall@$k=1$
+VLM-text
+DPR 11.6 9.5 19.2 19.2 14.9 15.9 15.8 25.6 34.7 27.0 19.3 19.2
+ColBERT 22.0 14.9 28.0 28.3 17.9 29.7 21.1 52.6 54.5 44.5 31.3 31.4
+BGE 19.2 15.2 24.6 28.7 12.8 27.6 19.7 47.0 52.3 35.8 28.3 29.0
+E5 15.9 8.8 27.7 24.3 14.6 21.8 14.7 45.6 53.0 40.5 26.7 26.4
+Contriever 23.4 7.5 28.2 26.8 17.1 25.7 16.1 43.6 51.5 42.9 28.3 28.9
+GTE 17.5 10.5 23.0 27.2 14.5 26.3 14.4 39.8 49.2 38.3 26.1 27.1
+[…]
+Image
+DSEwiki-ss 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
+DSEdocmatix 19.9 11.4 31.5 30.1 17.8 30.0 20.8 46.5 39.4 31.4 27.9 29.1
+ColPali 22.5 21.3 36.6 30.9 26.8 32.1 19.3 52.5 51.8 33.6 32.7 32.5
+DPR-Phi3ours 21.1 22.1 36.8 35.2 25.6 28.7 24.1 38.3 35.4 27.4 29.5 30.2
+Col-Phi3ours 22.6 22.0 37.5 34.9 28.9 30.3 22.7 50.2 45.1 26.3 31.1 31.6
+Recall@$k=5$
+VLM-text
+DPR 31.0 25.7 36.7 44.9 33.0 34.1 34.9 49.9 56.3 51.1 39.8 40.4
+ColBERT 41.8 37.7 53.7 61.8 35.1 52.4 46.1 83.2 70.1 62.5 54.4 56.0
+BGE 41.0 28.1 52.7 59.2 36.7 46.0 50.7 72.0 71.5 59.9 51.8 53.2
+E5 35.4 28.1 51.7 58.5 33.2 41.2 40.2 79.7 77.9 64.6 51.1 51.8
+Contriever 40.2 29.2 54.1 57.9 36.8 47.1 44.6 68.8 76.2 61.9 51.7 53.0
+GTE 36.7 25.6 51.2 56.6 39.7 46.6 46.4 72.2 74.3 63.2 51.3 52.3
+[…]
+Image
+DSEwiki-ss 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
+DSEdocmatix 39.6 36.2 53.9 57.5 33.7 52.5 42.8 69.4 63.1 48.9 49.8 51.9
+ColPali 40.7 45.9 54.9 58.5 42.6 51.2 45.7 76.8 74.5 48.9 54.0 54.3
+DPR-Phi3ours 45.5 37.7 57.0 62.9 41.4 51.1 45.5 65.1 60.8 49.3 51.6 53.9
+Col-Phi3ours 46.4 38.2 53.1 61.8 45.0 54.6 45.7 68.8 65.7 43.8 52.3 54.5
+Recall@$k=10$
+VLM-text
+DPR 42.2 33.1 52.1 56.2 39.9 43.5 44.0 62.8 61.7 59.7 49.5 50.5
+ColBERT 51.0 48.7 60.6 69.8 43.9 61.6 53.7 88.4 74.8 66.4 61.9 63.7
+BGE 51.1 38.7 62.1 71.5 41.9 55.6 58.7 80.8 78.7 63.5 60.3 62.4
+E5 45.3 38.6 62.0 70.5 45.6 50.0 55.3 87.1 82.4 66.8 60.4 61.2
+Contriever 49.9 41.3 62.0 70.5 44.8 56.5 54.5 81.3 78.0 64.9 60.4 62.2
+GTE 48.6 41.1 61.5 68.8 44.3 56.9 58.0 83.0 77.5 66.9 60.7 62.2
+[…]
+Image
+DSEwiki-ss 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
+DSEdocmatix 53.7 43.3 59.6 66.5 44.7 59.1 50.3 75.4 69.2 53.7 57.5 59.9
+ColPali 53.6 54.1 64.4 69.5 48.8 60.7 54.0 81.9 82.5 50.4 62.0 63.2
+DPR-Phi3ours 58.1 49.1 67.0 74.7 48.4 57.9 57.8 68.7 66.2 54.4 60.2 62.8
+Col-Phi3ours 57.7 50.5 66.6 72.3 50.7 59.3 53.6 68.5 74.8 57.5 61.1 63.3
+Table 6: Main results for layout-level retrieval (Refer to Table A for full results with VLM-text and Hybrid inputs).
+5.1 Evaluation Metric
+The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores.
+Recall@$k$ is defined as the proportion of ground truth page/layout evidence successfully retrieved.
+For layout matching, we calculate recall based on the overlaps between the bounding boxes of retrieved layouts and gold-standard layouts.
+5.2 Baseline Models and Setting
+We evaluate 6 state-of-the-art text retrievers: namely DPR, ColBERT, BGE, E5, Contriever, and GTE (see Appendix D.1).
+Additionally, we evaluate 5 VLM-based retrievers: 3 off-the-shelf models, namely DSEwiki-ss, DSEdocmatix, and ColPali (see Appendix D.2), and 2 models trained using MMDocIR training set (see Appendix C).
+Among all retrievers, ColBERT, ColPali, and Col-Phi3ours represent query/document as a list of token-level embeddings, while the other retrievers represent query/document as a single dense embedding.
+All retrievers are adapted to a dual-task setting:
+  • Page Retrieval: For text retrievers, we use the text from OCR-page or VLM page as described in Section 3.4. For visual retrievers, we directly utilize document page screenshots.
+  • Layout Retrieval: Text retrievers process multimodal layouts using OCR or VLM text (see Section 3.4).
+Visual retrievers process textual layouts using either Image input (cropped image of textual area) or Hybrid input (original text, as VLM can directly encode text).
+5.3 Main Results for Page-level Retrieval
+Table 4.2 presents the main results for page-level retrieval. Our key findings are as follows:
+  • Superiority of Visual Retrievers: Visual retrievers consistently outperform text retrievers across various domains and retrieval metrics, highlighting the advantage of using screenshots to retain multimodal cues often lost in text conversion.
+  • Effectiveness of MMDocIR: The visual retrievers trained on the MMDocIR training set demonstrate superior performance, demonstrating the value of high-quality training data.
+  • Necessity of Token-level Embeddings: Compared to dense-level retrievers (e.g., BGE, DSE, DPR-Phi3ours), token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) achieve more advantageous results in Recall@1 and have marginal performance improvement in Recall@3/5.
+However, token-level embedding can incur storage costs of 10 times more than a single embedding (DSE requires 0.24GB for indexing MMDocIR while ColPali requires 10.0GB).
+  • Top 5 Coverage: Retrieving top 5 pages provides substantial coverage of relevant information.
+5.4 Main Results for Layout-level Retrieval
+Table 5 shows the main results for layout-level retrieval. Our key findings are as follows:
+  • Effectiveness of VLM-Text: Interestingly, VLM-text approaches perform comparably to visual retrievers, demonstrating the promising image description capabilities of state-of-the-art VLM.
+This greatly benefits textual retrievers in multimodal understanding.
+  • Necessity of Token-level Embeddings: For layout retrieval tasks, token-level retrievers marginally outperform dense-level retrievers, demonstrating its importance of in such task.
+  • Top 10 Coverage: For layout retrieval tasks, retrieving top 10 layouts does not guarantee comprehensive coverage of the ground truth layout labels, emphasizing the complexity of the tasks.
+{tabu}
+ll|ccc|ccc
+Method   Page recall   Layout recall
+OCR VLM $\Delta$ OCR VLM $\Delta$
+$k=1$
+DPR 22.3 27.2 +4.9 12.6 19.3 +6.7
+ColBERT 40.3 45.8 +5.5 19.8 31.3 +11.5
+BGE 35.7 40.6 +4.9 19.0 28.3 +9.3
+E5 35.0 40.8 +5.8 18.4 26.7 +8.3
+Contriever 35.3 40.9 +5.6 18.8 28.3 +9.5
+GTE 35.4 38.9 +3.5 18.2 26.1 +7.9
+$k=3\text{ or }5$
+DPR 39.4 46.3 +6.9 23.7 39.8 +16.1
+ColBERT 58.8 64.9 +6.1 33.2 54.4 +21.2
+BGE 55.4 59.7 +4.3 32.7 51.8 +19.1
+E5 54.8 60.3 +5.5 33.3 51.1 +17.8
+Contriever 54.9 60.6 +5.7 31.7 51.7 +20.0
+GTE 54.9 58.7 +3.8 33.5 51.3 +17.8
+$k=5\text{ or }10$
+DPR 49.0 57.8 +8.8 29.9 49.5 +19.6
+ColBERT 66.0 72.3 +6.3 37.6 61.9 +24.3
+BGE 62.7 68.4 +5.7 37.8 60.3 +22.5
+E5 64.1 69.1 +5.0 39.0 60.4 +21.4
+Contriever 63.1 69.2 +6.1 37.3 60.4 +23.1
+GTE 63.2 67.6 +4.4 40.9 60.7 +19.8
+Table 7: Results of text retrievers using OCR/VLM-text.
+5.5 Text Retrieval: OCR-text vs VLM-Text
+Text retrievers leveraging VLM-text significantly outperform those using OCR-text in both tasks.
+Based on results, OCR-text is insufficient for multimodal retrieval, while VLM-text retains richer multimodal information.
+Although VLM-text offers much more comprehensive text information than OCR-text, it also introduces higher computational overhead and longer inference time.
+Most text retrievers based on on BERT Devlin et al.
+(2019), truncate input that exceed 512 tokens (approximately 380 english words).
+As shown in Figure 3(c), there are many pages containing more than 380 words (62.9% for OCR-page and 61.1% for VLM-page).
+Those pages suffer from critical information loss during page retrieval if the ground truth evidence is in the truncated part.
+In contrast, only a small fraction of layouts contain more than 380 tokens (3.9% for OCR-text, 4.8% for VLM-text, 0.5% for natural-text).
+Hence, as reflected in Table 4.2 and 5, text retriever demonstrates stronger performance on layout-level retrieval than on page-level retrieval.
+{tabu}
+ll|ccc
+Method   Layout recall
+Hybrid Image $\Delta$
+$k=1$
+DSEwiki-ss 24.6 28.2 +3.6
+DSEdocmatix 27.5 27.9 +0.4
+ColPali 28.5 32.7 +4.2
+DPR-Phi3ours 28.9 29.5 +0.6
+Col-Phi3ours 29.8 31.1 +1.3
+$k=5$
+DSEwiki-ss 46.7 50.2 +3.5
+DSEdocmatix 48.2 49.8 +1.6
+ColPali 52.2 54.0 +1.8
+DPR-Phi3ours 50.1 51.6 +1.5
+Col-Phi3ours 50.0 52.3 +2.3
+$k=10$
+DSEwiki-ss 55.8 58.5 +2.7
+DSEdocmatix 57.4 57.5 +0.1
+ColPali 60.0 62.0 +2.0
+DPR-Phi3ours 55.5 60.2 +4.7
+Col-Phi3ours 58.7 61.1 +2.4
+Table 8: Results of visual retrievers: image vs hybrid.
+5.6 Visual Retrieval: Image vs Hybrid input
+Visual retrievers tend to perform better when encoding text as images via visual encoders, rather than processing native textual input with LLM backbones.
+This advantage largely stems from their training setup: visual retrievers are typically optimized using text queries paired with image-based passages or documents, but are not fine-tuned directly on purely textual passages.
+However, encoding text as images incurs substantial computational overhead.
+Representing text as image tokens requires significantly more resources than native text encoding.
+To address this inefficiency and promote balanced retrieval capabilities, we advocate for future visual retrievers to be jointly trained or fine-tuned on both text and visual retrieval tasks.
+Such hybrid training would enable models to efficiently process text when appropriate, without compromising performance on visual inputs.
+{tabu}
+l|ll|ccc Model Store Index Search
+(GB) (MM:SS) (MM:SS)
+Page Processing
+Text
+DPR 0.06 6:53 00:02
+ColBERT 3.45 14:12 00:04
+BGE 0.08 7:31 00:03
+E5 0.08 8:26 00:03
+[…]
+[.2pt on 1pt off 1.5pt]2-6
+Image
+DPR-Phi3 0.24 101:20 00:04
+ColPali 10.00 47:14 00:05
+Col-Phi3 24.56 106:23 00:07
+Layout Processing
+Text
+DPR 0.51 41:33 00:15
+ColBERT 26.72 94:56 02:25
+BGE 0.66 55:05 00:18
+E5 0.66 60:21 00:18
+[…]
+[.2pt on 1pt off 1.5pt]2-6
+Image
+DPR-Phi3 1.99 735:51 00:44
+ColPali 83.50 262:29 09:06
+Col-Phi3 204.32 784:07 10:56
+[…]
+Hybrid
+DPR-Phi3 1.84 130:38 01:09
+ColPali 12.06 73:50 04:24
+Col-Phi3 22.72 140:44 02:41
+Table 9: Efficiency analysis different retrievers.
+5.7 Efficiency Analysis
+We evaluate the inference efficiency by measuring three key metrics : storage consumption, indexing time and, retrieval latency, as shown in Table  5.6.
+Experiments are conducted on a NVIDIA A100 40GB GPU with batch size of 4 for image and 256 for text.
+DPR-styled retrievers which generate single vector embeddings, demonstrates higher efficiency and lower computation across all metrics, compared to ColBERT-styled retrievers that produce token-level embeddings.
+Although DPR-styled retrievers slightly underperform in retrieval accuracy, their smaller embeddings size provide a significant advantage in the inference stage when storage space and inference time are concerned.
+Another key finding is that textual inputs are significantly more efficient than the visual inputs across all metrics.
+Meanwhile, hybrid retrieval system, which processes text in the image through LLM rather than visual encoders, further reduces memory and time consumption.
+Hence, future works on training hybrid retrieval system are encouraged as it offers a strong balance between computational efficiency and retrieval performance.
+6 Related Work
+DocCVQA Tito et al.
+(2021) proposes extracting information from a document image collection.
+However, it is limited by its small question set (20 questions).
+While PDF-MVQA Ding et al.
+(2024) is tailored for multimodal retrieval in biomedical articles, it is annotated by GPT-3.5-turbo rather than experts.
+SciMMIR Wu et al.
+(2024) also investigates multimodal retrieval but only provides image-caption pairs, lacking user queries paired with the corresponding document pages.
+Ma et al.
+(2024a) introduces two relevant datasets, namely Wiki-SS and DocMatix-IR.
+Wiki-SS is derived from natural questions Kwiatkowski et al.
+(2019) , wherein evidence passages are screenshots of Wikipedia pages.
+However, natural questions are primarily designed for text retrieval, and the provided screenshots may not consistently capture the ground-truth evidence, as only the front page is considered.
+DocMatix-IR is constructed from the large-scale DocMatix Laurençon et al.
+(2024) dataset using filtering and hard negative mining.
+However, the questions are generated by Phi-3-small Abdin et al.
+(2024) rather than human experts, and are not de-contextualized for retrieval task.
+ViDoRe Faysse et al.
+(2024) is the most relevant benchmark to MMDocIR.
+It integrates multiple DocVQA datasets and provides new industrial documents.
+Upon a thorough examination of the 2,400 questions, we find that over 80% questions exhibit notable limitations in terms of their complexity, contextual clarity, and the absence of complete document corpora.
+Refer to Appendix F for detailed quantitative and qualitative analysis of ViDoRe.
+7 Conclusion
+In conclusion, multimodal document retrieval presents a complex challenge that requires both understanding and integrating diverse data modalities beyond plain text.
+To more effectively evaluate these capabilities, we introduce the MMDocIR benchmark, which features the innovative dual-task retrieval capabilities targeting page-level and layout-level document granularity.
+The MMDocIR includes a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 73,843 questions, serving as a valuable resource for both training and evaluation of multimodal document retrieval.
+Our comprehensive empirical studies show that visual-driven retrievers significantly outperform text-driven ones, underscoring the importance of visual information in improving retrieval performance.
+Future work can expand upon these findings by optimizing retrieval algorithms to enhance both accuracy and efficiency of multimodal document retrieval systems.
+Limitations
+The limitations of MMDocIR are summarized as follows:
+  • Incomplete layout label annotations for training set: For 3 out of 7 training subsets, our semi-automated pipelines could not extract layout labels.
+These pipelines are optimized for datasets with single text or image layouts and cannot handle complex or cross-modal layouts.
+Future work should explore leveraging advanced vision-language models (VLMs) to facilitate annotation of layout labels for these subsets.
+  • Lack of joint text and visual training: As demonstrated in Section 5.6, all visual retrievers are suboptimal at modeling text passages, compared to modeling text as image screenshots.
+Our current visual retrievers do not explicitly utilize text query-document pairs to address this limitation.
+Future research should consider integrating both text and visual passages for joint training or finetuning to improve performance on both retrieval tasks.
+Appendix Table of Contents
+Due to page limits, we provide meaningful descriptions, results, implementation details, case studies, and other supplementary materials in Appendix. Our 21-page appendix is organized as follows:
+  • Appendix A: Supplementary experimental results for page and layout retrieval.
+  • Appendix B: process and implementation for MMDocIR curation.
+    • Related DocVQA benchmarks (§B.1)
+    • Multimodal corpora selection (§B.2)
+    • Question filtering for IR context (§B.3)
+    • Collection of training multimodal documents (§B.4)
+    • Curation of ground-truth training labels (§B.5)
+    • Selection of hard negatives (§B.6)
+    • Modality distribution (§B.7)
+    • Resource and Artifact URLs (§B.8)
+  • Appendix C: Training visual retrievers with MMDocIR training set.
+    • Encoding text queries and images (§C.1)
+    • Similarity calculation (§C.2)
+    • Training objectives (§C.3)
+    • Implementation details (§C.4)
+  • Appendix D: Text and visual retriever details.
+    • Text-centric retrieval (§D.1)
+    • Vision-driven retrieval (§D.2)
+    • Retriever implementation (§D.3)
+  • Appendix E: Examples from MMDocIR.
+    • Document pages (10 domains) (§E.1)
+    • Document layouts (§E.2)
+    • Page and layout annotations (§E.3)
+  • Appendix F: Analyses on ViDoRe benchmark.
+    • Search query and new annotation (§F.1)
+    • Corpora analysis (§F.2)
+  • Appendix G: License agreements.
+  • Appendix H: Ethical considerations.
+Appendix A Supplementary Experimental Results
+In this section, we provide the full results of page-level and layout-level retrievals, which supplement our main results discussion in Section 5.3 and Section 5.4 respectively.
+Specifically, Table A extends the main page-level results shown in Table 4.2 with the results of text retrievers using OCR-text.
+Table A extends the main layout-level results shown in Table 5 with the results of (i) text retrievers using OCR-text and (ii) visual retrievers using hybrid inputs.
+{tabu}
+lll|cccccccccc|cc
 Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
@@
 Col-Phi3ours 86.3 78.8 81.2 92.4 79.0 73.8 85.3 85.1 87.1 73.0 82.2 83.0
-Table 5.
-Main results for page-level retrieval.
-“OCR-text” and “VLM-text” refer to converting multi-modal content in the document page using OCR and VLM respectively.
+Table 10: Main results for page-level retrieval.
+“OCR-text” and “VLM-text” refer to converting multi-modal content using OCR and VLM respectively.
 “Image” refers to processing document page as screenshot image.
-5. Model Training: DPR-Phi3&Col-Phi3
-To evaluate the effectiveness of the MMDocIR training set, we train two visual retrievers based on Phi3-Vision (Abdin et al., 2024).
-Phi3-Vision ($\mathbf{M_{phi3v}}$) reuses the image tokenizer from clip-vit-large [^11] ViT-Large: https://huggingface.co/openai/clip-vit-large-patch14-336 ($\mathbf{M_{vit}}$).
-It can deal with high-resolution images by cropping them into sub-images, where each sub-image has $336\times 336$ pixels.
-Document/Query Encoding
-DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR (Karpukhin et al., 2020)) and a list of token-level embeddings (following ColBERT (Khattab and Zaharia, 2020)), respectively.
-Specifically, we follow Ma et al.
-(2024a) to concatenate document image with a text prompt: “¡s¿¡d¿ What is shown in this image?¡/s¿”.
-Here, the ¡d¿ token is a special placeholder token and is replaced by the sequence of patch latent embeddings from the vision encoder.
-We consider only text queries and use text prompt: “¡s¿ query: ¡q¿ ¡/s¿”.
-Similarly, the placeholder ¡q¿ token is replaced by input query.
-(1) $\begin{split}\mathrm{E_{d}^{dpr}}&=\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\\ &\mathrm{E_{q}^{dpr}}=\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\end{split}$
-where the end-of-sequence token ¡/s¿ from the last hidden state ($D_{1}=3072$) of $\mathbf{M_{phi3v}}$ is used to represent $\mathrm{E_{d}^{dpr}}$ and $\mathrm{E_{q}^{dpr}}$.
-(2) $\begin{split}\mathrm{E_{d}^{col}}&=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)},\in\mathbb{R}^{N_{d}\times D_{2}}\\ &\mathrm{E_{q}^{col}}=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)},\in\mathbb{R}^{N_{q}\times D_{2}}\end{split}$
-where $\mathbf{M_{proj}}$ is projection layer to map the last hidden states of $\mathbf{M_{phi3v}}$ into reduced dimension $D_{2}=128$.
-$N_{d}\approx 2500$ for a typical high-resolution page and $N_{q}$ is the number of query tokens.
-Query-Doc Similarity
-The similarity between the query and the document is computed as follows:
-(3) $\text{Sim}(q,d)_{dpr}=\frac{\langle\mathrm{E_{q}^{dpr}}|\mathrm{E_{d}^{dpr}}\rangle}{\left\|\mathrm{E_{q}^{dpr}}\right\|\cdot\left\|\mathrm{E_{d}^{dpr}}\right\|}$
-[…]
-(4) $\text{Sim}(q,d)_{col}=\sum_{i\in[1,N_{q}]}\max_{j\in[1,N_{d}]}\langle\mathrm{E_{q}^{col}}^{(i)}|\mathrm{E_{d}^{col}}^{(j)}\rangle$
-[…]
-Contrastive Loss
-Given the query $q$, we have the positive document $d^{+}$ and a set of negative documents $d^{-}$ including hard negatives and in-batch negatives.
-The hard negatives are negative pages within the document with highest $Sim(q,d^{-})$ scored by ColPali (Faysse et al., 2024) retriever.
-We calculate the loss as:
-(5) ($\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$
-where DPR-Phi3 is trained on the InfoNCE loss, and the temperature parameter $\tau=0.02$ in our experiments.
-(6) ($\mathcal{L}^{col}_{(q,d^{+},d^{-})}=\log\Big{(}1+\exp\big{(}\max_{d_{i}\in d^{-}}(\text{Sim}^{col}_{(q,d_{i})})-\text{Sim}^{col}_{(q,d^{+})}\big{)}\Big{)}$
-where Col-Phi3 is trained via the softplus loss based on the positive scores w.r.t. to the maximal negative scores.
-6. Experiment
 {tabu}
-llll—cccccccccc—cc
+llll|cccccccccc|cc
 Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
@@
 Contriever 6.7 7.0 3.8 14.3 4.1 13.3 8.3 43.6 43.9 42.9 18.8 18.1
+GTE 5.7 7.0 3.9 17.2 2.8 15.8 9.4 40.7 41.6 38.3 18.2 18.4
 […]
@@
 Contriever 23.4 7.5 28.2 26.8 17.1 25.7 16.1 43.6 51.5 42.9 28.3 28.9
+GTE 17.5 10.5 23.0 27.2 14.5 26.3 14.4 39.8 49.2 38.3 26.1 27.1
 […]
@@
 ColPali 17.7 12.3 30.0 18.4 19.0 25.5 20.6 49.7 51.2 40.9 28.5 27.1
-DPR-Phi3ours 28.3 11.1 35.5 39.2 29.3 25.2 27.4 37.0 32.9 22.2 28.8 30.5
-Col-Phi3ours 26.4 12.6 33.7 37.5 30.1 27.9 24.6 46.2 47.4 21.9 30.8 32.0
+DPR-Phi3ours 28.3 11.1 35.5 18.8 29.3 24.0 27.4 38.0 41.9 34.5 28.9 27.3
+Col-Phi3ours 26.4 12.6 33.7 27.3 30.1 27.9 24.6 46.2 47.4 21.9 29.8 29.6
 Recall@$k=5$
@@
 Contriever 11.9 17.9 11.9 28.8 9.3 24.6 18.1 64.3 68.0 62.7 31.7 31.2
+GTE 10.0 18.2 12.8 32.9 15.2 29.4 21.0 67.7 65.3 62.4 33.5 33.5
 […]
@@
 Contriever 40.2 29.2 54.1 57.9 36.8 47.1 44.6 68.8 76.2 61.9 51.7 53.0
+GTE 36.7 25.6 51.2 56.6 39.7 46.6 46.4 72.2 74.3 63.2 51.3 52.3
 […]
@@
 ColPali 40.1 38.3 55.2 49.2 42.7 47.6 40.8 78.6 68.7 61.0 52.2 51.4
-DPR-Phi3ours 54.2 27.4 53.8 64.9 36.6 45.3 49.6 69.5 63.8 53.2 51.8 54.4
-Col-Phi3ours 50.9 25.5 49.1 66.2 40.3 48.1 48.3 62.3 60.6 48.5 50.0 53.5
+DPR-Phi3ours 54.2 27.4 53.8 39.9 36.6 45.4 49.6 67.2 66.8 59.8 50.1 49.3
+Col-Phi3ours 50.9 25.5 49.1 58.3 41.9 48.1 49.2 62.3 60.6 48.5 50.0 51.8
 Recall@$k=10$
@@
 Contriever 15.1 25.7 14.2 36.8 15.9 27.9 24.2 76.8 71.8 64.9 37.3 36.6
+GTE 19.1 29.0 21.4 39.0 19.2 32.9 29.2 78.8 74.5 66.2 40.9 40.1
 […]
@@
 Contriever 49.9 41.3 62.0 70.5 44.8 56.5 54.5 81.3 78.0 64.9 60.4 62.2
+GTE 48.6 41.1 61.5 68.8 44.3 56.9 58.0 83.0 77.5 66.9 60.7 62.2
 […]
@@
 ColPali 52.1 46.4 65.0 64.4 50.7 53.8 51.0 82.7 71.4 62.5 60.0 60.2
-DPR-Phi3ours 65.2 33.7 60.3 72.9 42.4 54.1 52.9 79.1 72.5 65.6 59.9 62.9
+DPR-Phi3ours 65.2 33.7 60.3 51.3 42.4 52.4 52.9 79.1 72.5 65.6 55.5 53.4
 Col-Phi3ours 59.1 38.7 57.0 77.7 43.9 57.7 51.7 72.4 68.0 60.7 58.7 62.8
-Table 6.
-Main results for layout-level retrieval.
-“OCR-text” and “VLM-text” refer to converting multi-modal layouts using OCR and VLM respectively.
-“Pure-Image” and “Hybrid” refer to reading textual layouts in image and text format respectively.
-6.1. Evaluation Metric
-The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores.
-Recall@$k$ is defined as the proportion of the ground truth page/layout evidence that is successfully retrieved.
-Note for layout matching, we calculate recall based on the overlaps of bounding boxes between retrieved layouts and gold layouts.
-6.2. Baseline Models and Setting
-We evaluate 6 state-of-the-art text retrievers: namely DPR (Karpukhin et al., 2020), ColBERT (Khattab and Zaharia, 2020), BGE (Xiao et al., 2023), E5 (Wang et al., 2022), Contriever (Izacard et al., 2022), and GTE (Li et al., 2023) (refer to Section 7.1).
-Meanwhile, we evaluate 5 VLM-based retrievers: 3 off-the-shelf models named DSEwiki-ss (Ma et al., 2024a), DSEdocmatix (Ma et al., 2024a), and ColPali (Faysse et al., 2024) (refer to Section 7.2), and 2 our trained models (see section 5).
-All retrievers are adapted into dual-task setting:
-  • Page Retrieval: For textual retrievers, we use MinerU (Wang et al., 2024a) to extract the layout of all document pages.
-Non-textual layouts are converted into ‘OCR-text” and “VLM-text” by using Tesseract OCR  [^12] https://github.com/tesseract-ocr/tesseract and GPT-4o respectively.
-The document page is represent by concatenating the text of all layouts within it.
-For visual retrievers, we directly utilize document page screenshots.
-  • Layout Retrieval: Textual retrievers process non-textual layouts using “OCR-text” or “VLM-text”.
-Visual retrievers process textual layouts in either “Pure Image” (using cropped image of textual area as input) or “Hybrid” (using its original text, as VLM can directly encode text as well).
-6.3. Main Results for Page-level Retrieval
-For page-level retrieval (Table 4.2), our key findings are as follows:
-  • Superiority of Visual Retrievers: Visual retrievers consistently outperform text retrievers across various domains and retrieval metrics (e.g., Top $k=1,3,5$).
-This highlights the significance of leveraging document screenshots to capture multi-modal information, which is often lost when relying solely on OCR-text.
-  • Effectiveness of MMDocIR: Among the visual retrievers, DPR-Phi3ours and Col-Phi3ours trained on theMMDocIR train set demonstrate superior performance.
-It suggests dataset quality to effectively enhances the retrieval ability.
-  • Effectiveness of VLM-Text: Although VLM-text approaches underperform visual retrievers, they achieve much better performance than the OCR-text methods.
-This indicates benefits of using GPT-4o to preserve visual cues in text.
-  • Necessity of Token-level Embeddings: Token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) compared with their document/page-level counterparts (e.g., BGE, DSE, DPR-Phi3ours), achieve more advantageous results in Recall@1 and marginal performance increase in Recall@5/10.
-Yet, the storage overhead of token-level embeddings can be 10 times more than page-level embeddings.
-  • Top 5 Coverage: Retrieving top 5 pages provides substantial coverage, ensuring that relevant information is captured.
-6.4. Main Results for Layout-level Retrieval
-For layout-level retrieval (Table 6), our key findings are as follows:
-  • Superiority of Visual Retrievers: Visual retrievers exhibit substantial performance advantages over text retrievers utilizing OCR-text.
-This highlights the limitations of OCR in capturing the multi-modal nature of documents, where visual context can significantly enhance retrieval accuracy.
-  • Effectiveness of VLM-Text: Interestingly, VLM-text approaches achieve comparable performance as visual retrievers.
-This indicates strong image description capabilities of state-of-the-art VLM, which can offer significant benefits to textual retrievers in understanding multi-modal information.
-  • Comparison of Hybrid vs.
-Pure Image Sequences: Visual retrievers relying on hybrid image-text sequences generally perform less effectively than those utilizing pure image sequences.
-This suggests that current VLMs may have stronger capabilities in modeling images than text within the multi-modal framework.
-  • Necessity of Token-level Embeddings: Token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) compared with their document/page-level counterparts (e.g., BGE, DSE, DPR-Phi3ours), do not achieve advantageous results in layout retrieval.
-  • Top 10 Coverage: Retrieving top 10 layouts does not provide comprehensive coverage of the ground truth layouts for document retrieval tasks. It suggests the challenge of layout retrieval.
-6.5. Analysis of OCR and VLM Text
-Interestingly, text retrievers leveraging VLM-text can significantly outperform OCR-text, in both page and layout retrieval.
-OCR-text which is the the raw text extracted via OCR tools, and as experimental results suggest, it is not suitable for multi-modal retrieval.
-In comparison, VLM-text can facilitate multi-modal retrieval, suggesting that VLMs can largely preserve rich multi-modal information.
-The averaged word length and distribution of OCR and VLM text obtained from the tables and figures in MMDocIR in shown in Figure 3.
-Observe that the VLM-text is much longer than OCR-text.
-Specifically, the length of VLM-text is 1.5 and 3.8 times more than that of OCR-text in table and figure.
-Tables tend to contain numerical or structured layouts, which can be well recorded by text.
-Figures contain mostly graphical elaboration and visual cues.
-Hence, making OCR tools struggle in getting textual information from figures.
-Although VLM-text can offer much comprehensive text information than OCR-text, it comes at the cost of higher computational overhead and longer latency.
-(a) Avg word length
-(b) Distribution density of word length
-Figure 3. Avg length and distribution of OCR and VLM text.
-7. Related Work
-7.1. Text-Centric Document Retrieval
-Document indexing.
-The process of indexing a multi-modal document involves multiple steps, including Document Parsing (Chao and Fan, 2004; Wang et al., 2024a), Optical Character Recognition (OCR) (Chaudhuri et al., 2017; Borovikov, 2014; Mori et al., 1999), Layout Detection (Sassioui et al., 2023; Xu et al., 2020, 2021), Chunking (Chen et al., 2024a; Raina and Gales, 2024; Dong et al., 2024), and Image Captioning (You et al., 2016; Aneja et al., 2018).
+Table 11: Main results for layout-level retrieval. “Pure-Image” and “Hybrid” refer to reading textual layouts in image and text format respectively.
+Appendix B Dataset Construction
+B.1 Related DocVQA Benchmarks
+Early DocVQA benchmarks primarily address single-page visual question answering (VQA), exemplified by datasets such as DocVQA Mathew et al.
+(2021), InfoVQA Mathew et al.
+(2022), and TAT-DQA Zhu et al.
+(2022).
+To overcome the limitations of single-page inputs, subsequent datasets like DUDE Landeghem et al.
+(2023), MP-DocVQA Tito et al.
+(2023), and SlideVQA Tanaka et al.
+(2023) have extended the context length to an average of 5.7, 8.3, and 20 pages, respectively.
+More recent benchmarks, including MMLongBench-Doc Ma et al.
+(2024b) and DocBench Zou et al.
+(2024), treat DocVQA as a long-context task, accommodating entire documents that average between 50–70 pages in length.
+As document lengths increase, retrieval becomes essential.
+Relevant pages must first be identified, followed by answer generation based on the retrieved multimodal evidence.
+B.2 Document Corpora Collection Criteria
+To facilitate the development of MMDocIR, we leverage resources from existing DocVQA benchmarks described in Appendix B.1.
+Despite not being curated for IR, they offer valuable document corpora and questions that can be adapted for IR tasks.
+We select relevant DocVQA datasets based on the following criteria:
+  • Document Source: The dataset must include accessible original documents or sources for these documents. We need to access and enrich them to support more complex retrieval tasks.
+  • Diverse Domain/Modality: The document collections must (1) encompass diverse domains suitable for generalized evaluation, and (2) contain multiple modalities, such as text, figures, tables, charts, and layouts.
+  • Long Document: We choose documents with extensive texts as longer texts pose more significant challenges. This criterion can evaluate models in handling complex and lengthy documents.
+  • Question Diversity and Comprehensiveness: The questions included in the dataset should be diverse and challenging.
+For example, cross-modal questions require reasoning across both text and visual tables/figures; multi-hop questions require reasoning over multiple steps; multi-page questions require combining information from multiple pages.
+Considering these criteria, we utilize document corpora and questions from datasets as follows:
+  • Evaluation: MMLongBench-Doc Ma et al. (2024b) and DocBench Zou et al. (2024).
+  • Training: MP-DocVQA Tito et al. (2023), SlideVQA Tanaka et al. (2023), TAT-DQA Zhu et al. (2022), SciQAG Wan et al. (2024), DUDE Landeghem et al. (2023), and CUAD Hendrycks et al. (2021).
+B.3 Question Filtering Guidelines
+We filter questions based on the following criteria:
+  • Summarization Questions: Questions such as “What does this book mainly illustrate?”“What does this story mainly tell?” require understanding of large sections or even the entire document.
+The broad scope makes it hard to pinpoint specific content and contradicts the IR nature of our task.
+  • Overwhelm Statistical Questions: Questions that demand extensive data computation or collation, such as “How many words are there in total in the paper?”“How many pictures are there in total in the document?” are also excluded from our scope.
+  • Online Search Questions: Questions like “What is the Google Scholar citation count of the author?” rely on information from external online resources.
+We focus only on retrieving information within the documents, and therefore exclude these questions.
+  • Unanswerable Questions: These are designed to test if models generate answers based on non-existent information (model hallucinations).
+Since they do not facilitate the retrieval of factual document-based information, these questions are excluded.
+B.4 Training Document Collection
+We collect the training datasets as follows:
+  • MP-DocVQA Tito et al.
+(2023) contains 47,952 images collected from Industry Documents Library (IDL)  [^3] https://www.industrydocuments.ucsf.edu/.
+IDL is a crucial resource for public health research, containing millions of documents produced by industries such as tobacco, drug, chemical, and food, which have had significant impacts on public health.
+We group the 47,952 document images into separate document files, and obtain 875 long documents (46.8 pages on average) with 15,266 QA pairs.
+  • SlideVQA Tanaka et al.
+(2023) contains 2,619 slide documents collected from slideshare  [^4] https://www.slideshare.net/ and covering 39 topics.
+SlideVQA hosts a wide variety of slide presentations across various categories such as business, mobile, social media, marketing, technology, arts, career, design, education, and government & nonprofit, among others, which can enrich the diversity of our corpus.
+Note that SlideVQA contains only the first 20 pages for each slide deck.
+In our research, we manually collect the remaining missing pages, and obtain 2,011 long documents (averaging 49.3 pages) with 11,066 QA pairs.
+SlideVQA requires complex reasoning, including single-hop, multi-hop, and numerical reasoning, and also provides annotated arithmetic expressions of numerical answers for enhancing the ability of numerical reasoning.
+  • TAT-DQA Zhu et al.
+(2022) consists of 3,067 document pages from financial reports  [^5] https://www.annualreports.com/, dated between 2018 and 2020.
+AnnualReports.com provides access to a comprehensive collection of corporate annual reports from over 10,320 companies worldwide.
+Note that neither original documents nor links are provided.
+We use OCR to extract text in the pages, and use search engine to find relevant documents.
+After careful tracing and recognition, we identify 163 original documents (averaging 147.3 pages) with 15,814 QA pairs.
+  • arXivQA Li et al.
+(2024) comprises 32k figures cropped from academic pages  [^6] https://arxiv.org/.
+The papers on arXiv cover a wide range of disciplines including physics, mathematics, computer science, quantitative biology, quantitative finance, statistics, engineering, and systems science, and economics, etc.
+We use the arXiv DOIs provided to collect the academic papers.
+Due to the missing of paper versions, extra efforts are made to identify paper versions.
+After careful tracing, recognition, and document length filtering, we identify 1,579 documents averaging 18.4 pages.
+  • SciQAG Wan et al.
+(2024) consists of 22,728 papers and 188,042 QA pairs in 24 scientific disciplines, collected from Web of Science (WoS) Core Collection database.
+WoS provides comprehensive scientific literature in natural sciences, social sciences, arts, and humanities.
+We sample 50 documents from each discipline, and manually collect 1,197 papers using the DOIs provided.
+  • DUDE Landeghem et al.
+(2023) provides 5,019 documents from aggregator websites7771: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/.
+It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires.
+We filter out short documents and obtain 779 relatively long documents (averaging 15.6 pages) with 3,173 QA pairs.
+  • CUAD Hendrycks et al.
+(2021) provides 510 commercial legal contracts, collected from Electronic Data Gathering, Analysis, and Retrieval (EDGAR) [^8] https://www.sec.gov/search-filings.
+EDGAR contracts are usually more complex and heavily negotiated than the general population of all legal contracts.
+We filter out short documents in CUAD and obtain 274 long documents (29.6 pages on average) with 11,234 QA pairs.
+(a) 313 documents in MMDocIR evaluation set in Section 3
+. Refer to caption
+(b) 313 documents in MMDocIR training set in Section 4
+. Refer to caption
+Figure 4: Area ratio of different modalities (1) in overall and (2) by domains/datasets in MMDocIR evaluation and training set.
+Note that white spaces, headers, and footers are excluded from the area calculations.
+B.5 Training Dataset Label Construction
+The page labels can be directly obtained in the MP-DocVQA, SlideVQA, and DUDE datasets. Among these, only DUDE provides layout labels.
+SciQAG provides only question and answer in texts.
+We use these information to infer the page-level and layout-level labels.
+Specifically, we first use MinerU to obtain layout-level passage chunks.
+For each QA pair, we deploy E5 and BGE retrievers to obtain question-passage and answer-passage similarity scores against all extracted passage chunks.
+If both scores rank within top 3 for a specific passage chunk, we assign this layout as the layout-level labels for the given QA pair.
+Similarly, arXivQA provides only cropped images, without document page/layout labels.
+We first use MinerU to obtain layout-level images.
+For each cropped image, we calculate its similarity against all extracted images using brute-force matcher [^9] https://opencv.org/, and select the most similar one.
+Subsequently, we manually examine if the selected image matches the cropped image.
+In this way, we filter around 20% unmatched images, resulting 1,579 questions with page and layout level labels.
+For TAT-DQA, layout-level labels are provided for each sampled page.
+To localize the page index of the sampled pages, we first utilize PDF mapping tool [^10] https://github.com/pymupdf/PyMuPDF to retrieve best matched page in the document.
+Then, we manually verify whether the retrieved page matches the given page, and correct the labels if there were any errors.
+B.6 Hard Negative Sampling
+In addition to annotating ground truth (positive) page labels, we enhance our training data with negative labels.
+In the context of retrieval, hard negatives are particularly informative non-relevant documents that closely resemble true positives according to the model’s current scoring function.
+Unlike randomly selected negatives, hard negatives are challenging for the model to distinguish from relevant documents, thus providing stronger supervision.
+In our framework, hard negatives are crucial for improving retrieval performance.
+By training the model on these challenging examples, we encourage it to learn more discriminative representations, ultimately enhancing its robustness and reducing false positives during retrieval.
+As described in Appendix C.3, training is conducted using a contrastive loss, where the model aims to separate relevant documents from irrelevant ones.
+Specifically, we obtain hard negatives using the ColPali retriever Faysse et al.
+(2024), which scores all document pages for a given query.
+The irrelevant pages with the highest top-$k$ scores (i.e., those most likely to be confused with positives) are selected as hard negatives for training.
+Artifacts Purpose Referred Section Resource URL
+MMLong’-Doc Eval-set curation Section 3 https://github.com/mayubo2333/MMLongBench-Doc
+DocBench Eval-set curation https://github.com/Anni-Zou/DocBench
+MP-DocVQA Train-set curation Section 4 Appendix B.4 https://rrc.cvc.uab.es/?ch=17&com=tasks
+SlideVQA Train-set curation https://github.com/nttmdlab-nlp/SlideVQA
+TAT-DQA Train-set curation https://github.com/NExTplusplus/TAT-DQA
+arXivQA Train-set curation https://huggingface.co/datasets/taesiri/arxiv_qa
+SciQAG Train-set curation https://github.com/MasterAI-EAM/SciQAG
+DUDE Train-set curation https://github.com/duchallenge-team/dude
+CUAD Train-set curation https://www.atticusprojectai.org/cuad
+MinerU Doc Parsing Section 3.2 https://github.com/opendatalab/MinerU
+Tesseract OCR OCR-text Section 3.4 https://github.com/tesseract-ocr/tesseract
+GPT-4o VLM-text https://openai.com/index/hello-gpt-4o/
+QwenVL2.5 VLM-text https://github.com/QwenLM/Qwen2.5-VL
+PyMuPDF Page matching Appendix B.5 https://github.com/pymupdf/PyMuPDF
+OpenCV Image matching https://opencv.org/
+Text Retriever Layout location BGE and E5 (see Table D.1)
+Visual Retriever Hard negatives Appendix B.6 Colpali (see Table D.1)
+Table 12: Artifacts used to facilitate construction of MMDocIR evaluation & train set.
+B.7 Fine-grained Modality Distribution
+MMDocIR evaluation set includes 313 long documents with an average length of 65.1 pages, categorized into ten main domains: research reports, administration&industry, tutorials&workshops, academic papers, brochures, financial reports, guidebooks, government documents, laws, and news articles.
+Overall, the modality distribution is: Text (60.4%), Image (18.8%), Table (16.7%), and other modalities (4.1%), as shown in Figure 4(a) Different domains exhibit different distributions of multimodal information.
+For instance, research reports, tutorials, workshops, and brochures predominantly contain images, whereas financial and industry documents are table-rich.
+In contrast, government and legal documents primarily comprise text.
+MMDocIR training set includes 6,878 long documents with an average length of 32.6 pages, categorized into seven Document VQA or QA datasets: MP-DocVQA Tito et al.
+(2023), SlideVQA Tanaka et al.
+(2023), TAT-DQA Zhu et al.
+(2022), arXivQA Li et al.
+(2024), SciQAG Wan et al.
+(2024), and DUDE Landeghem et al.
+(2023).
+Overall, the modality distribution is: Text (49.3%), Image (34.3%), Table (10.8%), and other modalities (4.9%), as shown in Figure 4(b).
+Each dataset features unique distributions of multimodal content.
+The legal documents and academic papers are text-intensive.
+The slides consist mostly of visual features.
+Industrial documents and financial reports are table-intensive.
+B.8 Resource URL of Artifacts
+In this section, we summarize the artifacts used to facilitate the construction of MMDocIR’s evaluation and train set, as shown in Table 12.
+These artifacts mainly includes: datasets used for curating MMDocIR evaluation and training sets, tools for parsing documents, packages for locating evidence, and etc.
+Appendix C Model Training: DPR-Phi3&Col-Phi3
+To evaluate the effectiveness of the MMDocIR training set, we train two visual retrievers based on Phi3-Vision Abdin et al.
+(2024).
+Phi3-Vision ($\mathbf{M_{phi3v}}$) reuses the image tokenizer from clip-vit-large [^11] ViT-Large: https://huggingface.co/openai/clip-vit-large-patch14-336 ($\mathbf{M_{vit}}$).
+It can deal with high-resolution images by cropping them into sub-images, where each sub-image has $336\times 336$ pixels.
+C.1 Document/Query Encoding
+DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR Karpukhin et al.
+(2020)) and a list of token-level embeddings (following ColBERT Khattab and Zaharia (2020)), respectively.
+Specifically, we follow Ma et al.
+(2024a) to concatenate document image with a text prompt: “<s><d> What is shown in this image?</s>”.
+Here, the <d> token is a special placeholder token and is replaced by the sequence of patch latent embeddings from the vision encoder.
+We consider only text queries and use text prompt: “<s> query: <q> </s>”.
+Similarly, the placeholder <q> token is replaced by input query.
+We encode query $q$ and document $d$ in two ways:
+$\begin{split}\mathrm{E_{d}^{dpr}}&=\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\\ \mathrm{E_{q}^{dpr}}&=\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)}[-1],\in\mathbb{R}^{D_{1}}\end{split}$ (1)
+where the end-of-sequence token </s> from the last hidden state ($D_{1}=3072$) of $\mathbf{M_{phi3v}}$ is used to represent $\mathrm{E_{d}^{dpr}}$ and $\mathrm{E_{q}^{dpr}}$.
+$\begin{split}\mathrm{E_{d}^{col}}&=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}\mathbf{M_{vit}}(d),\text{prompt}\big{)}\\ \mathrm{E_{q}^{col}}&=\mathbf{M_{proj}}\cdot\mathbf{M_{phi3v}}\big{(}q,\text{prompt}\big{)}\end{split}$ (2)
+where $\mathrm{E_{d}^{col}}\in\mathbb{R}^{N_{d}\times D_{2}}$ and $\mathrm{E_{q}^{col}}\in\mathbb{R}^{N_{q}\times D_{2}}$, and $\mathbf{M_{proj}}$ is projection layer to map the last hidden states of $\mathbf{M_{phi3v}}$ into reduced dimension $D_{2}=128$.
+$N_{d}\approx 2500$ for a typical high-resolution page and $N_{q}$ is the number of query tokens.
+C.2 Query-Doc Similarity
+The similarity between the query and the document is computed as follows:
+$\text{Sim}(q,d)_{dpr}=\frac{\langle\mathrm{E_{q}^{dpr}}|\mathrm{E_{d}^{dpr}}\rangle}{\left\|\mathrm{E_{q}^{dpr}}\right\|\cdot\left\|\mathrm{E_{d}^{dpr}}\right\|}$ (3)
+[…]
+$\text{Sim}(q,d)_{col}=\sum_{i\in[1,N_{q}]}\max_{j\in[1,N_{d}]}\langle\mathrm{E_{q}^{col}}^{(i)}|\mathrm{E_{d}^{col}}^{(j)}\rangle$ (4)
+[…]
+C.3 Contrastive Loss
+Given the query $q$, we have the positive document $d^{+}$ and a set of negative documents $d^{-}$ including hard negatives and in-batch negatives.
+The hard negatives are negative pages within the document with highest $Sim(q,d^{-})$ scored by ColPali Faysse et al.
+(2024) retriever, refer to Appendix B.6 for more details on hard negative selection.
+We calculate the loss as:
+($\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$ (5)
+where DPR-Phi3 is trained on the InfoNCE loss, and the temperature parameter $\tau=0.02$ in our experiments.
+($\mathcal{L}^{col}_{(q,d^{+},d^{-})}=$ $\log\Big{(}1+\exp\big{(}\max_{d_{i}\in d^{-}}(\text{Sim}^{col}_{(q,d_{i})})$
+$-\text{Sim}^{col}_{(q,d^{+})}\big{)}\Big{)}$ (6)
+where Col-Phi3 is trained via the softplus loss based on the positive scores w.r.t. to the maximal negative scores.
+C.4 Training Implementation Details
+In summary, we train two visual retrievers based on Phi3-Vision Abdin et al.
+(2024).
+DPR-Phi3 and Col-Phi3 represent document page or query using a single dense embedding (following DPR Karpukhin et al.
+(2020)) and a list of token-level embeddings (following ColBERT Khattab and Zaharia (2020)), respectively.
+To train the model, we employ memory-efficient techniques such as LoRA Hu et al.
+(2022), FlashAttention Dao (2024), and DeepSpeed Rasley et al.
+(2020).
+The model is trained with a batch size of 64 for one epoch on MMDocIR training set.
+The model weights are shared between the language models for document screenshot and query encoding.
+In both tasks, each training query is paired with one positive document and one hard negative document.
+The document screenshots are resized to $344\times 1,$ 344 × 1 , 344 pixels and cropped into $4\times 4$ sub-images.
+The training process is performed on four A100 80GB GPUs.
+Appendix D Retrievers: Introduction and Implementation Details
+D.1 Text-Centric Document Retrieval
+For text retrieval, the first step is to convert multimodal document into text using techniques, e.g., Document Parsing Chao and Fan (2004); Wang et al.
+(2024), Optical Character Recognition (OCR) Chaudhuri et al.
+(2017); Borovikov (2014); Mori et al.
+(1999), Layout Detection Sassioui et al.
+(2023); Xu et al.
+(2020, 2021), Chunking Chen et al.
+(2024a); Raina and Gales (2024); Dong et al.
+(2024), and Image Captioning You et al.
+(2016); Aneja et al.
+(2018).
 These steps are time-consuming and can introduce errors that impact the overall retrieval performance.
-Text retrieval.
-Current text retrieval are primarily classified into sparse and dense retrieval.
-For two widely-used sparse retrievers: TF-IDF (Salton et al., 1983) calculates the relevance via word frequency with the inverse document frequency, and BM25 (Robertson et al., 1994) introduces nonlinear word frequency saturation and length normalization.
+Current text retrieval are primarily categorized as sparse or dense retrieval.
+For two widely-used sparse retrievers: TF-IDF Salton et al.
+(1983) calculates the relevance via word frequency with the inverse document frequency, and BM25 Robertson et al.
+(1994) introduces nonlinear word frequency saturation and length normalization.
 Dense retrievers encode content into vector representations.
-DPR (Karpukhin et al., 2020) is the pioneering work of dense vector representations for QA tasks.
-Similarly, ColBERT (Khattab and Zaharia, 2020) introduces an efficient question-document interaction model with late fine-grained term matching.
-Contriever (Izacard et al., 2022) leverages contrastive learning to improve content dense encoding.
-E5 (Wang et al., 2022) and BGE (Xiao et al., 2023) propose novel training and data preparation techniques to enhance retrieval performance.
-Moreover, GTE (Li et al., 2023) integrates graph-based techniques to enhance dense embedding.
-Most systems focus on text-based representations, neglecting the valuable visual information present in documents.
-7.2. Vision-Driven Document Retrieval
-Vision Language Models (VLMs) (Abdin et al., 2024; Beyer et al., 2024; Bai et al., 2023; Chen et al., 2024b) have gained popularity for their ability to understand and generate text based on combined text and visual inputs.
-This advancement has led to the development of cutting-edge visual-driven retrievers, such as ColPali (Faysse et al., 2024) and DSE (Ma et al., 2024a).
-These models specifically leverage PaliGemma (Beyer et al., 2024) and Phi3-V (Abdin et al., 2024) to directly encode document page screenshots for multi-modal document retrieval.
-ColPali adopts similar question-document interaction as ColBERT, and represent each document page in token-level embeddings.
-By contrast, DSE is similar to DPR to encode each page with a single dense embedding.
-Visual Retrievers can directly model useful visual information, allowing for the direct utilization of multi-modal content without first converting it into text.
+DPR Karpukhin et al.
+(2020) is the pioneering work of dense vector representations for QA tasks.
+Similarly, ColBERT Khattab and Zaharia (2020) introduces an efficient question-document interaction model with late fine-grained term matching.
+Contriever Izacard et al.
+(2022) leverages contrastive learning to improve content dense encoding.
+E5 Wang et al.
+(2022) and BGE Xiao et al.
+(2023) propose novel training and data preparation techniques to enhance retrieval performance.
+Moreover, GTE Li et al.
+(2023) integrates graph-based techniques to enhance dense embedding.
+However, most text retrieval systems overlook valuable visual information present in documents.
+{tabu}
+l|l|c|c|l Model Dimension Base Model HuggingFace Checkpoint
+Text
+DPR 768 BERT-base facebook/dpr-ctx_encoder-multiset-base facebook/dpr-question_encoder-multiset-base
+[…]
+ColBERT $N_{\mathrm{tok}}\times$768 BERT-base colbert-ir/colbertv2.0
+[…]
+Contriever 768 BERT-base facebook/contriever-msmarco
+[…]
+E5 1,024 BERT-large intfloat/e5-large-v2
+[…]
+BGE 1,024 RetroMAE BAAI/bge-large-en-v1.5
+[…]
+GTE 1,024 BERT-large thenlper/gte-large
+Visual
+DSEwiki-ss 3,072 Phi-3-Vision Tevatron/dse-phi3-v1.0
+[…]
+DSEdocmatix 3,072 Phi-3-Vision Tevatron/dse-phi3-docmatix-v2
+[…]
+ColPali $N_{\mathrm{tok}}\times$1,024 PaliGemma vidore/colpali
+Table 13: Implementation details for Text and Vision Retrieval Models
+D.2 Vision-Driven Document Retrieval
+Vision Language Models (VLMs) Abdin et al.
+(2024); Beyer et al.
+(2024); Bai et al.
+(2023); Chen et al.
+(2024b) can understand and generate text based on combined text and visual inputs.
+This advancement has led to the development of cutting-edge visual-driven retrievers, such as ColPali Faysse et al.
+(2024) and DSE Ma et al.
+(2024a).
+These models specifically leverage PaliGemma Beyer et al.
+(2024) and Phi3-Vision Abdin et al.
+(2024) to directly encode document page screenshots for multimodal document retrieval.
+ColPali adopts a similar question-document interaction as ColBERT, and represents each document page in token-level embeddings.
+By contrast, DSE is similar to DPR in that it encodes each page with a single dense embedding.
+Visual retrievers are capable of modeling useful visual information, allowing direct utilization of multimodal content without first converting it into text first.
 Despite these advancements, visual retrievers face challenges, particularly in dealing with text details when document page resolutions are high.
 The high resolution of document pages substantially increases the computational cost and complexity of the embedding process, which may hinder the model’s performance.
-7.3. Multi-modal Document Retrieval Datasets
-As described in Section 1 and Table 1, there is a notable lack of a robust benchmark for multi-modal document retrieval.
-DocCVQA (Tito et al., 2021) is the first multi-modal document retrieval-answering task, which extracts information from a document image collection and then provides the answer.
-However, DocCVQA provides only 20 questions, and the corpora is limited to finance domain specifically in political disclosure.
-PDF-MVQA (Ding et al., 2024) is tailored for multi-modal retrieval in research articles.
-Note that PDF-MVQA is annotated by Chat-GPT (GPT-3.5-turbo) but not experts, and its corpora contains only biomedical articles.
-SciMMIR (Wu et al., 2024) proposes multi-modal retrieval in scientific research papers.
-However, it provides only the image-caption pairs, not the natural user query with the targeting document pages.
-Ma et al.
-(2024a) introduce two relevant datasets, namely Wiki-SS and DocMatix-IR.
-Wiki-SS is constructed based on Natural Questions (Kwiatkowski et al., 2019) where the evidence passage is the screenshot wikipedia webpage.
-Natural Questions are mainly designed for text rather than multi-modal retrieval.
-We also notice that the screenshot may not include the ground-truth evidence as only the front-page is screenshotted.
-DocMatix-IR is constructed from the largest document understanding dataset DocMatix (Laurençon et al., 2024) via filtering and hard negative mining.
-However, the VQA questions in DocMatix is constructed by Phi-3-small (Abdin et al., 2024) rather than experts, and these VQA questions are not de-contextualized for retrieval.
-ViDoRe (Faysse et al., 2024) is the most relevant benchmark to MMDocIR.
-It integrates multiple DocVQA datasets (Mathew et al., 2021, 2022; Zhu et al., 2022; Li et al., 2024), and provides new documents in scientific, medical, administrative, and environment domains.
-Upon thorough examination of the 1,180 questions within the ViDoRe dataset, we find that more than 80% of these questions present drawbacks in terms of their complexity and context-specific nature.
-Typical examples include “What’s the title of the picture” and “what’s the author’s name”.
-They tend to be either overly simplistic or excessively dependent on a specific context which does not necessarily reflect the broader, more complex needs of effective IR.
-Additionally, a significant limitation of the ViDoRe dataset is its reliance on sampled document pages or cropped images rather than providing the complete document corpora.
-8. Conclusion
-In conclusion, multi-modal document retrieval presents a complex challenge that necessitates the integration of diverse data modalities beyond plain text.
-Our contributions address this challenge by introducing the MMDocIR benchmark, which includes innovative dual-task retrieval functionalities targeting both page-level and layout-level specifics of documents.
-The MMDocIR benchmark comprises a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 173,843 questions, making it a pivotal resource for advancing multi-modal document retrieval for both training and evaluation purposes.
-Our comprehensive evaluation reveals that visual-driven retrievers significantly outperform text-driven ones, underscoring the critical importance of incorporating visual content in enhancing the effectiveness of document retrieval systems.
-Future work can expand on optimizing these retrieval algorithms to improve the accuracy and efficiency of multi-modal document retrieval systems.
+D.3 Implementation Details
+In our experiments (refer to Section 5.2), we implement 9 off-the-shelf retrievers including 6 text retrievers and 3 visual retrievers.
+The text retrieval models deployed are namely DPR, ColBERT, Contriever, E5, BGE and GTE.
+These models use the WordPiece tokenizer from BERT and also inherit the maximum input length of 512 tokens from BERT Devlin et al.
+(2019).
+Additionally, we make use of the sentence-transformer library [^12] https://www.sbert.net/ when deploying E5, BGE and GTE.
+The visual retrieval models deployed are namely DSEwiki-ss, DSEdocmatix, and ColPali.
+We use pre-trained checkpoints available on HuggingFace  [^13] https://huggingface.co/; the specific checkpoint information can be found in Table D.1 alongside other configuration details.
+Appendix E Dataset Demonstration
+E.1 Document Pages by Domains
+The documents in MMDocIR can be categorized into 10 types. We provide examples of each type as below.
+  • Admin & Industry: These documents primarily consist of instructional and overview content on industry, reflected by the dominance of text-based questions (78.0%) and a smaller reliance on visual evidence (image questions only 20.3%), which shows a text-heavy structure (70.1%).
+Some detailed examples are shown in Figure 5(b).
+  • Tut & Workshop: Documents in this category focus on slides or tutorials, which exhibit a balanced question modality: 61.7% text, 24.5% image, and 9.5% table questions.
+Strong visual components are present, with 57.4% of its content being images—the highest among all categories.
+Some detailed examples are shown in Figure 5(c).
+  • Academic Paper: These documents are formal publications with structured layouts, citations, and academic pictures.
+The questions span multiple modalities: 28.8% text, 25.7% image, and 50.0% table.
+Text modality dominates content distribution (74.6%), with the presence of tables (11.1%) and images (12.8%) demonstrating rich multimodal alignment and explicit questions with answers.
+Some detailed examples are shown in Figure 6(a).
+  • Brochure: Designed for promotional purposes, the brochure category contains highly visual documents.
+Over 52.6% of questions are image-based—the highest among all domains—while text-based questions account for only 60.5%.
+Modality distribution is similarly diverse: 50.8% image, showcasing their visually complex layout.
+Some detailed examples are shown in Figure 6(b).
+  • Financial Report: These documents involve massive numerical and quantitative data, reflected in a high proportion of table questions (54.5%) and strong table content distribution (29.2%).
+While text remains significant (60.3%), the inclusion of tabular and numerical analysis is essential for understanding these documents.
+Some detailed examples are shown in Figure 6(c).
+  • Guidebook: Instruction manuals for electronics and tools, guidebooks exhibit the most balanced question modality: 51.8% text, 54.4% image, and 26.8% table, indicating multimodal instructional designs.
+Some detailed examples are shown in Figure 7(a).
+  • Government: This category covers policy files and governmental reports.
+It is highly text-centric with 69.9% text questions and 88.2% text content.
+This reflects the formal and regulatory nature of such documents.
+Some detailed examples are shown in Figure 7(b).
+  • Laws: Legal documents exhibit strong textual dominance both in questions (62.1%) and content (83.8%), with very limited visual presence (image content only 1.6%).
+They often maintain specific formats and focus on linguistic interpretation rather than visual layout.
+Some detailed examples are shown in Figure 7(c).
+  • News: Although based on only one document, the “News” domain shows notable multimodal richness.
+It includes a significant image portion (39.8%), high text presence (48.5%), and 11.6% titles.
+This reflects the use of images and headlines typical of news articles.
+Some detailed examples are shown in Figure 7(d).
+E.2 Document Layouts
+In this section, we present 9 pages along with their detected layouts, which are highlighted for better visualizations, as shown in Figure 8, 9, and 10.
+Specifically, layout detection identifies the spatial location of different content types, such as images, tables, and text within a document.
+With the help of layout detection, we can precisely locate the specific position of an answer, whether it is an image, a text paragraph, or a table.
+This enables a more fine-grained layout-level evaluation of multimodal retrieval capabilities.
+E.3 Annotation Examples
+In this section, we present 4 annotation examples that illustrate typical multimodal retrieval and reasoning patterns, which help explain the construction and retrieval process.
+Each annotation includes the following primary components: question, answer, page-level labels, and layout-level labels.
+The page-level labels show the selected pages that contain ground truth evidence.
+Based on these selected pages, layout-level labels further display the specific layout box detection of evidence.
+These examples frequently require reasoning across multiple pages and modalities.
+The evidence encompasses diverse formats such as figures, charts, tables, and texts, highlighting the complexity and richness of the multimodal retrieval tasks.
+(a) Page screenshots in Research Report domain.
+(b) Page screenshots in Administration & Industry domain.
+(c) Page screenshots in Tutorial & Workshop domain.
+Figure 5: The screenshot examples of typical document pages for (a) Research Report, (b) Administration & Industry, and (c) Tutorial & Workshop domain.
+(a) Page screenshots in Academic Paper domain.
+(b) Page screenshots in Brochure domain.
+(c) Page screenshots in Financial Report domain.
+Figure 6: The screenshot examples of typical document pages for (a) Academic Paper, (b) Brochure, and (c) Financial Report domain.
+(a) Page screenshots in Guidebook domain.
+(b) Page screenshots in Government domain.
+(c) Page screenshots in Laws domain.
+(d) Page screenshots in News domain.
+Figure 7: The screenshot examples of typical document pages for (a) Guidebook, (b) Government, (c) Laws domain, and (d) News domain.
+(a) Example 1: original page vs. page highlighted with layout bounding boxes.
+(b) Example 2: original page vs. page highlighted with layout bounding boxes.
+(c) Example 3: original page vs. page highlighted with layout bounding boxes.
+Figure 8: The 3 examples illustrate the function and effectiveness of layout detection on document pages.
+(a) Example 1: original page vs. page highlighted with layout bounding boxes.
+(b) Example 2: original page vs. page highlighted with layout bounding boxes.
+(c) Example 3: original page vs. page highlighted with layout bounding boxes.
+Figure 9: The 3 examples illustrate the function and effectiveness of layout detection on document pages.
+(a) Example 1: original page vs. page highlighted with layout bounding boxes.
+(b) Example 2: original page vs. page highlighted with layout bounding boxes.
+(c) Example 3: original page vs. page highlighted with layout bounding boxes.
+Figure 10: The 3 examples illustrate the function and effectiveness of layout detection on document pages.
+Figure 11: This example shows a typical image retrieval and reasoning task that requires synthesizing information from pie chart.
+Figure 12: This example shows a typical multi-page retrieval task that requires synthesizing information from text passages across multiple pages.
+Figure 13: This example shows a typical image reasoning task that requires synthesizing information from specific image.
+Figure 14: This example shows a typical multi-page image and text reasoning task that requires synthesizing cross-modal information from image and text.
+Sub dataset #Not #Barely #Suit- #Total HuggingFace Resource
+suitable suitable able
+arXiVQA 245 203 52 500 vidore/arxivqa_test_subsampled
+DocVQA 345 130 25 500 vidore/docvqa_test_subsampled
+InfoVQA 139 284 77 500 vidore/infovqa_test_subsampled
+TAT-DQA 373 121 6 500 vidore/tatdqa_test
+Industrial 78 260 62 400
+vidore/syntheticDocQA_energy_test
+vidore/syntheticDocQA_healthcare_industry_test
+vidore/syntheticDocQA_government_reports_test
+vidore/syntheticDocQA_artificial_intelligence_test
+Sum 1180 998 222 2,400 -
+Percentage 49.1% 41.5% 9.25% - -
+Table 14: Document statistics for ViDoRe Benchmark.
+Appendix F Detailed Analysis of ViDoRe Benchmark
+F.1 Query and Annotation Analysis
+As mentioned in Section 6, ViDoRe Faysse et al.
+(2024) is the most relevant benchmark to MMDocIR.
+It integrates several datasets such as DocVQA Mathew et al.
+(2021), InfoVQA Mathew et al.
+(2022), TAT-DQA Zhu et al.
+(2022), arXiVQA Li et al.
+(2024), and providing new documents in scientific, medical, administrative, and environment domains.
+In this Appendix, we elaborate our analysis of the sampled 2,400 questions sampled from ViDoRe.
+The statistics are shown in Table 14.
+ViDoRe test set contains questions in either English or French.
+In our work, we examine only the English questions.
+For academic subsets, we examine all 1,500 questions: 500 questions from DocVQA, 500 questions from InfoVQA, and 500 questions from arXiVQA.
+In TAT-DQA, we sample and examine the first 500 questions.
+For the industrial documents, we select 100 questions from each domain (i.e., energy, healthcare, government, and artificial intelligence).
+We examine sampled questions and summarize these questions into 3 categories:
+  • Unsuitable Queries.
+Queries that are not well-suited for IR systems can often burden these systems by generating numerous irrelevant results.
+For example, a query such as “What’s the x-axis of the figure” is likely to prompt matches from multiple passages within document corpora that mention figures with an x-axis.
+This tends to happen because the query is overly broad and lacks contextual specificity.
+When such queries stem from Document Visual Question Answering (DocVQA) tasks targeting a single image, the challenge is exacerbated, as the reliance on precise context increases while the target remains too vague, undermining the fundamental principles of effective IR.
+  • Barely Suitable Queries.
+Queries that fall into this category provide some guidance towards locating useful passages, yet suffer from a lack of precise detail.
+These queries often fetch moderate number of passages, where both relevance and focus may not be as sharp.
+For example, the query “What was the total assets from AMER in 2018?” is meant for Visual Question Answering (VQA) focused on a specific financial topic.
+Although this seems specific, the issue arises when multiple sections within an annual report discuss AMER’s total assets for the year.
+This causes significant confusion since ViDoRe is set to acknowledge only a single passage as the verified answer.
+This lack of uniqueness in the ground truth makes it hard to evaluate the actual performance of IR system.
+  • Suitable Queries.
+The most effective queries for IR systems are characterized by their specificity and ability to distinguish between different sections of texts.
+These queries often involve precise facts or detailed inquiries that facilitate pinpointing exact passages.
+For instance, the question “What was the magnitude of the earthquake that occurred in Maule on 2/27/2010?” incorporates significant keywords and details that guide the retrieval system directly to the necessary data.
+Such queries align perfectly with the objectives of IR, leveraging specificity and detailed context to efficiently retrieve most relevant information.
+The comprehensive analysis of our queries, as presented in Table 14, reveals a significant challenge in adapting questions from Visual Question Answering (VQA) datasets (such as DocVQA, InfoVQA, TAT-DQA, and arXiVQA) for Information Retrieval (IR) purposes.
+Only 8% of these queries prove suitable for effective IR usage.
+In comparison, queries derived from industrial documents perform slightly better, with 15.5% deemed suitable.
+A common issue identified is that these queries are either excessively simplistic or highly specific to a particular context.
+Our findings suggest that the primary difficulty stems from the inherent differences between DocIR and DocVQA.
+VQA queries are typically crafted to address content on a specific page or within a particular image, inherently limiting their scope and specificity.
+This specificity and simplism are functional within the confines of the intended VQA context but pose substantial limitations when such queries are repurposed for IR tasks.
+Due to this gap, we exclude ViDoRe benchmark from our experiments.
+F.2 Document Corpora Analysis
+ViDoRe bootstrap document corpora directly from existing DocVQA benchmarks (i.e., DocVQA, InfoVQA, TAT-DQA, arXiVQA) that perform single-page VQA.
+In the DocVQA setting, only selected pages are provided for VQA, rather than the entire document pages.
+For arXivQA, the retrieved passages are not document pages, but are cropped images (e.g., figures, tables, and charts).
+In our experiments, we need to rely on the entire documents pages to evaluate retrieval on long documents.
+To bridge the gap of missing complete document corpora, we put in considerable efforts to collect the original documents of existing DocVQA datasets, as mentioned in Section 4.1.
+Appendix G License Agreements
+We ensure that the distribution of each dataset complies with the corresponding licenses, all of which are listed below:
+  • MMLongBench-Doc: is under Apache-2.0 license agreement for academic research purposes.
+  • DocBench: we achieved the agreement of usage as academic research from the dataset’s author.
+  • MP-DocVQA: is under “MIT License” license agreement for academic research purposes.
+  • SlideVQA: is under “NTT License” license agreement for academic research purposes.
+  • TAT-DQA: is under “CC-BY-4.0” license agreement for academic research purposes.
+  • ArXivQA: is under “CC-BY-SA-4.0” license agreement for academic research purposes.
+  • SciQAG: is under “CC-BY-4.0” license agreement for academic research purposes.
+  • DUDE: is under “GPL-3.0” license agreement for academic research purposes.
+  • CUAD: is under “CC-BY-4.0” license agreement for academic research purposes.
+For the new annotations contributed in MMDocIR, including but not limited to the questions, page and layout annotations, we make them available solely for research purposes.
+Users are permitted to use, modify, and share these annotations for academic and non-commercial research activities.
+Any other use, including commercial exploitation, is not permitted without explicit written permission from the authors.
+Appendix H Ethical Considerations
+The introduction and broader adoption of MMDocIR may have potential ethical impacts spanning both positive and negative dimensions.
+Below, we outline possible negative consequences and discuss potential mitigation strategies:
+Privacy Risks: MMDocIR enables models to retrieve relevant information over lengthy, multimodal documents, which may include sensitive personal, financial, or health information.
+There is a risk that such technologies could be leveraged for large-scale surveillance, unauthorized extraction of personal data, or other privacy violations.
+Fairness and Bias: If benchmarked models are trained or evaluated on data that does not reflect diverse demographic, linguistic, and backgrounds, outputs may exhibit biases.
+This may lead to unfair decision-making or stereotypes.
+Mitigation Strategies: To mitigate these risks, we make sure that: (i) Benchmark development uses only publicly available, carefully vetted datasets, with sensitive information anonymized or removed; (ii) Retrieval outputs are monitored for bias and fairness.
+We encourage researchers and practitioners employing MMDocIR to be mindful of these factors and to actively work toward responsible development and deployment, including transparency about limitations and proactive safeguards where needed.
+We welcome community feedback and collaboration on best practices to further reduce risks as this technology evolves.

### v3

@@
 Red boundary boxes represent layout-level labels.
-$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally
-Figure 2: Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark.
+[…]
+Figure 2: Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark. Note that the white spaces, headers, and footers are removed from the area counting.
 1 Introduction
@@
 MMLongBench-Doc VQA question 1,082 TXT/C/TAB/I Multiple 47.5
-Wiki-SS Natural question 3,610 TXT Wikipedia 1.0
+Wiki-$S$ Natural question 3,610 TXT Wikipedia 1.0
 DocMatix-IR VQA question 5.61m TXT/C/TAB/I Multiple 4.2
@@
 (2022); Chen et al.
-(2023) which primarily deals with textual data, multimodal document retrieval imposes substantially greater demands on understanding multimodal elements such as images, tables, charts, and layout designs.
+(2023); Dong et al.
+(2024); Wang et al.
+(2023) and long-context Yang et al.
+(2025) QA which primarily deals with textual data, multimodal document retrieval imposes substantially greater demands on understanding multimodal elements such as images, tables, charts, and layout designs.
 Such elements often carry significant information that plain text fails to convey Cui et al.
@@
 (2024b) in Figure 2 shows that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively.
-This highlights the need for retrieval systems that effectively handle multimodal information.
-However, as shown on (see Table 1), existing benchmarks exhibit several critical limitations that undermine comprehensive evaluation of multimodal retrieval systems.
+This highlights the need for retrieval systems that effectively handle multimodal and cross modal Zhang et al.
+(2025); Dong et al.
+(2025b) information.
+However, as shown on Table 1, existing benchmarks exhibit several critical limitations that undermine comprehensive evaluation of multimodal retrieval systems.
 The key limitations include: 1.
@@
 (1) The page-level retrieval identifies the most relevant pages within a document to answer user query.
-(2) The layout-level retrieval targets the most relevant fine-grained elements such as paragraphs, equations, figures, tables, and charts (see Appendix E.2 for more examples).
+(2) The layout-level retrieval targets the most relevant layouts.
+A layout is an element on the document page where the element could be a paragraph, a heading, an equation, a table, a figure, or a chart (see Appendix E.2 for more examples).
 Such task supports more precise and context-aware retrieval that pinpoint specific elements to address user queries.
@@
 In summary, our contributions are threefold:
-  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework (§$\S$ 2) that supports page-level and fine-grained layout-level multimodal document retrieval.
+  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework ($\S$ 2) that supports page-level and fine-grained layout-level multimodal document retrieval.
   • MMDocIR Benchmark: We introduce Multimodal Document Information Retrieval benchmark.
-The evaluation set (§$\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 questions.
-The training set (§$\S$ 4) consists of 6,878 documents and labels for 73,843 questions.
-  • We conduct extensive experiments and comparisons of both text and visual retrievers (§$\S$ 5), demonstrating clear advantage of incorporating visual content in multimodal retrieval tasks.
+The evaluation set ($\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 questions.
+The training set ($\S$ 4) consists of 6,878 documents and labels for 73,843 questions.
+  • We conduct extensive experiments and comparisons of both text and visual retrievers ($\S$ 5), demonstrating clear advantage of incorporating visual content in multimodal retrieval tasks.
 2 Dual-Task Retrieval Definition
-Let $\mathcal{D}$ be a document corpora consisting of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection.
+Let $\mathcal{D}$ be a document corpora consisting of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: $\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection.
+[^2] A document page usually comprises about 5 to 15 layouts, depending on its complexity..
 The objective is to perform document retrieval at both page-level and layout-level.
@@
 Given that documents in MMDocIR contain 65.1 pages on average, pinpointing relevant pages is highly non-trivial, akin to finding a needle in haystack, which demands careful inspection and document understanding.
+Our annotation process is described as follows:
   • For DocBench: we manually annotate page labels for all 864 questions from scratch, by carefully reviewing each document and locating the pages containing answer evidence.
@@
 This effort results in corrections to 10 answers and 169 page labels [^3] Common errors in page labeling: annotators starting page indexing at 1 rather than 0, missing labels for questions spanning multiple pages, and incorrect or absent page labels..
+Through these efforts, we successfully obtain page-level labels for a total of 1,658 questions in MMDocIR.
 Layout-level Annotation.
@@
 Compared to page annotation, layout-level labeling is significantly more complex and labor-intensive.
-  • Layout Detection. We begin by utilizing MinerU Wang et al. (2024) to automatically parse all documents and detect all layouts (e.g., layout type and bounding boxes).
+Our method of annotations is as follows:
+  • Layout Detection.
+We begin by utilizing MinerU Wang et al.
+(2024) to automatically parse all documents and detect all layouts (e.g., layout type and bounding boxes).
+[^4] These elements are categorized into five main types: text, image, table, title, and equation.
+The bounding boxes are represented by coordinates of top-right and bottom-left corners.
   • Evidence Identification.
 We identify the layouts that contain necessary answer evidence.
-In case where MinerU fails to detect evidentiary element, we manually annotate the the bounding boxes, accounting for 7% of the total layout-level labels.
+In case where MinerU fails to detect evidentiary element, we manually annotate the bounding boxes, accounting for 7% of the total layout-level labels.
+Ultimately, this meticulous process leads to the annotation of 2,638 layout labels in MMDocIR.
 3.3 Quality Control
@@
 Each group is responsible for annotating approximately 1,000 questions, with an overlap of 400 questions serving the need for cross-validation.
+The process for quality control works as follows:
 […]
@@
 Notably, the length of VLM-text is 1.5 and 3.8 times of OCR-text for table (with more structured numbers) and figure (mostly with visual elaboration), respectively.
+[^5] Tables typically contain structured numbers, which can be well recorded by text.
+Figures contain mostly visual elaboration, making it hard to extract raw text.
 For pages, we construct two variants by combining the natural text with either OCR-text and VLM-text for each page, resulting in OCR-page and VLM-page representations.
@@
 Train Subsets assorted docs 6,878 73,843 32.6 6.32 42.6 49.3 34.3 10.8 4.9
-[…]
 - MP-DocVQA health/ind. docs 875 15,266 46.8 6.9 38.8 57.3 18.0 22.7 1.9
@@
 Different domains feature distinct multimodal distribution.
+For instance, reports, tutorials, workshops, and brochures predominantly contain images, whereas financial and industry documents are table-rich.
+In contrast, government and legal documents primarily comprise text.
 The overall modality distribution is as follows: text (60.4%), image (18.8%), table (16.7%), and others (4.1%), with fine-grained distribution shown in Figure 4(a).
@@
 MMDocIR includes 1,658 questions, and 2,107 page and 2,638 layout labels.
-The evidence spans 4 modalities: text (44.7%), image (21.7%), table (37.4%), and layout/meta (11.5%).
+The evidence spans 4 modalities: text (44.7%), image (21.7%), table (37.4%), and layout/meta [^6] The “layout/meta” refers to evidence related to layout information and meta-data statistics.
+(11.5%).
 Notably, MMDocIR presents several challenges: 254 questions require cross-modal understanding, 313 questions require evidence across multiple pages, and 637 questions require reasoning over multiple layouts.
@@
 The overall statistics (e.g., document information, modality distribution, domain, etc) of MMDocIR training set are summarized in Table 4.
-{tabu}
-lll|cccccccccc|cc
 Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
 Report &Indu. Worksh. Paper ure Report book ment Macro Micro
-Recall@$k=1$
-VLM-text
-DPR 32.3 25.5 27.0 31.0 28.4 18.8 23.5 31.2 38.3 16.1 27.2 26.9
+Recall@$k=1$ VLM-text DPR 32.3 25.5 27.0 31.0 28.4 18.8 23.5 31.2 38.3 16.1 27.2 26.9
 ColBERT 48.6 42.8 51.1 46.2 36.0 36.8 49.6 60.9 59.5 26.3 45.8 44.9
@@
 GTE 46.5 26.3 48.7 38.9 35.9 27.0 46.2 50.1 45.8 24.1 38.9 37.9
-[…]
-Image
-DSEwiki-ss 53.0 50.0 54.0 48.7 45.1 43.0 51.5 46.9 54.2 33.6 48.0 47.5
+Image DSEwiki-$s$ 53.0 50.0 54.0 48.7 45.1 43.0 51.5 46.9 54.2 33.6 48.0 47.5
 DSEdocmatix 52.3 40.4 56.1 51.7 45.8 43.5 53.8 53.7 58.3 46.7 50.2 50.1
@@
 Col-Phi3ours 56.7 50.4 56.9 61.3 54.8 50.7 60.8 61.3 63.6 54.0 57.0 57.1
-Recall@$k=3$
-VLM-text
-DPR 52.2 44.2 43.5 54.6 52.0 35.1 44.4 53.9 57.2 25.5 46.3 46.2
+Recall@$k=3$ VLM-text DPR 52.2 44.2 43.5 54.6 52.0 35.1 44.4 53.9 57.2 25.5 46.3 46.2
 ColBERT 70.1 64.4 70.3 72.3 59.1 55.3 71.1 81.3 70.8 34.3 64.9 64.8
@@
 GTE 71.1 44.5 67.2 64.4 54.3 43.0 70.6 71.9 68.2 31.4 58.7 58.3
-[…]
-Image
-DSEwiki-ss 75.4 65.0 73.9 79.8 69.5 63.5 75.4 71.5 81.4 50.4 70.6 71.4
+Image DSEwiki-$s$ 75.4 65.0 73.9 79.8 69.5 63.5 75.4 71.5 81.4 50.4 70.6 71.4
 DSEdocmatix 75.4 67.5 73.3 80.0 66.3 61.6 72.8 76.4 82.6 57.7 71.4 71.8
@@
 Col-Phi3ours 80.2 74.1 77.4 84.8 69.1 67.7 78.7 79.5 81.8 69.3 76.3 76.8
-Recall@$k=5$
-VLM-text
-DPR 66.5 60.1 56.0 68.9 58.8 43.8 57.1 68.6 64.8 33.6 57.8 57.8
+Recall@$k=5$ VLM-text DPR 66.5 60.1 56.0 68.9 58.8 43.8 57.1 68.6 64.8 33.6 57.8 57.8
 ColBERT 78.8 74.0 78.7 82.3 66.1 60.8 77.0 88.5 78.0 38.7 72.3 72.3
@@
 GTE 77.4 62.6 74.7 75.8 62.0 51.8 77.8 80.0 75.0 39.4 67.6 67.2
-[…]
-Image
-DSEwiki-ss 84.0 80.2 78.7 87.0 75.7 73.0 82.0 77.3 88.3 58.4 78.5 79.2
+Image DSEwiki-$s$ 84.0 80.2 78.7 87.0 75.7 73.0 82.0 77.3 88.3 58.4 78.5 79.2
 DSEdocmatix 82.1 77.2 79.6 87.8 73.9 72.4 81.7 83.1 89.4 67.9 79.5 80.1
@@
 Col-Phi3ours 86.3 78.8 81.2 92.4 79.0 73.8 85.3 85.1 87.1 73.0 82.2 83.0
-Table 5: Main results for page-level retrieval, with the best results in boldface and second best results underlined. For clarity, we omit results using VLM-text (Refer to Table A for full results).
+Table 5: Main results for page-level retrieval, with the best results in boldface and second best results underlined. For clarity, we omit results using VLM-text (Refer to Table 11 for full results).
 5 Experiment
-{tabu}
-lll|cccccccccc|cc
 Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
 Report &Indu. Worksh. Paper ure Report book ment Macro Micro
-Recall@$k=1$
-VLM-text
-DPR 11.6 9.5 19.2 19.2 14.9 15.9 15.8 25.6 34.7 27.0 19.3 19.2
+Recall@$k=1$ VLM-text DPR 11.6 9.5 19.2 19.2 14.9 15.9 15.8 25.6 34.7 27.0 19.3 19.2
 ColBERT 22.0 14.9 28.0 28.3 17.9 29.7 21.1 52.6 54.5 44.5 31.3 31.4
@@
 GTE 17.5 10.5 23.0 27.2 14.5 26.3 14.4 39.8 49.2 38.3 26.1 27.1
-[…]
-Image
-DSEwiki-ss 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
+Image DSEwiki-$s$ 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
 DSEdocmatix 19.9 11.4 31.5 30.1 17.8 30.0 20.8 46.5 39.4 31.4 27.9 29.1
@@
 Col-Phi3ours 22.6 22.0 37.5 34.9 28.9 30.3 22.7 50.2 45.1 26.3 31.1 31.6
-Recall@$k=5$
-VLM-text
-DPR 31.0 25.7 36.7 44.9 33.0 34.1 34.9 49.9 56.3 51.1 39.8 40.4
+Recall@$k=5$ VLM-text DPR 31.0 25.7 36.7 44.9 33.0 34.1 34.9 49.9 56.3 51.1 39.8 40.4
 ColBERT 41.8 37.7 53.7 61.8 35.1 52.4 46.1 83.2 70.1 62.5 54.4 56.0
@@
 GTE 36.7 25.6 51.2 56.6 39.7 46.6 46.4 72.2 74.3 63.2 51.3 52.3
-[…]
-Image
-DSEwiki-ss 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
+Image DSEwiki-$s$ 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
 DSEdocmatix 39.6 36.2 53.9 57.5 33.7 52.5 42.8 69.4 63.1 48.9 49.8 51.9
@@
 Col-Phi3ours 46.4 38.2 53.1 61.8 45.0 54.6 45.7 68.8 65.7 43.8 52.3 54.5
-Recall@$k=10$
-VLM-text
-DPR 42.2 33.1 52.1 56.2 39.9 43.5 44.0 62.8 61.7 59.7 49.5 50.5
+Recall@$k=10$ VLM-text DPR 42.2 33.1 52.1 56.2 39.9 43.5 44.0 62.8 61.7 59.7 49.5 50.5
 ColBERT 51.0 48.7 60.6 69.8 43.9 61.6 53.7 88.4 74.8 66.4 61.9 63.7
@@
 GTE 48.6 41.1 61.5 68.8 44.3 56.9 58.0 83.0 77.5 66.9 60.7 62.2
-[…]
-Image
-DSEwiki-ss 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
+Image DSEwiki-$s$ 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
 DSEdocmatix 53.7 43.3 59.6 66.5 44.7 59.1 50.3 75.4 69.2 53.7 57.5 59.9
@@
 Col-Phi3ours 57.7 50.5 66.6 72.3 50.7 59.3 53.6 68.5 74.8 57.5 61.1 63.3
-Table 6: Main results for layout-level retrieval (Refer to Table A for full results with VLM-text and Hybrid inputs).
+Table 6: Main results for layout-level retrieval (Refer to Table 12 for full results with VLM-text and Hybrid inputs).
 5.1 Evaluation Metric
 The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores.
-Recall@$k$ is defined as the proportion of ground truth page/layout evidence successfully retrieved.
+Recall@k@k is defined as the proportion of ground truth page/layout evidence successfully retrieved.
+For page matching, the recall is straightforwardly computed with page indices.
 For layout matching, we calculate recall based on the overlaps between the bounding boxes of retrieved layouts and gold-standard layouts.
+Unlike page retrieval, where boundaries are unambiguous, layout detection tools can produce differing bounding boxes for the same content.
+Our ground-truth layouts include both MinerU outputs and manual annotations for cases where MinerU misses elements.
+During our evaluation, retrievers are provided with MinerU predicted layouts (e.g., bboxes and types), but some ground truth bboxes cannot be exactly matched to them.
+Therefore, simple binary classifications (matched or not matched) are insufficient.
+Overlap-based recall offers a nuanced and realistic evaluation, especially where perfect alignment is not guaranteed.
 5.2 Baseline Models and Setting
@@
   • Layout Retrieval: Text retrievers process multimodal layouts using OCR or VLM text (see Section 3.4).
-Visual retrievers process textual layouts using either Image input (cropped image of textual area) or Hybrid input (original text, as VLM can directly encode text).
+Visual retrievers [^7] Most visual retrievers are not explicitly trained on text query-doc pairs, this setup constitutes out-of-domain data.
+process textual layouts using either Image input (cropped image of textual area) or Hybrid input (original text, as VLM can directly encode text).
 5.3 Main Results for Page-level Retrieval
-Table 4.2 presents the main results for page-level retrieval. Our key findings are as follows:
+Table 5 presents the main results for page-level retrieval. Our key findings are as follows:
   • Superiority of Visual Retrievers: Visual retrievers consistently outperform text retrievers across various domains and retrieval metrics, highlighting the advantage of using screenshots to retain multimodal cues often lost in text conversion.
   • Effectiveness of MMDocIR: The visual retrievers trained on the MMDocIR training set demonstrate superior performance, demonstrating the value of high-quality training data.
-  • Necessity of Token-level Embeddings: Compared to dense-level retrievers (e.g., BGE, DSE, DPR-Phi3ours), token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) achieve more advantageous results in Recall@1 and have marginal performance improvement in Recall@3/5.
+  • Effectiveness of VLM-Text: Although VLM-text approaches underperform visual retrievers, they achieve much better performance than the OCR-text methods.
+This indicates benefits of using GPT-4o to preserve visual cues in text.
+  • Effect of Token-level Embeddings: Compared to dense-level retrievers (e.g., BGE, DSE, DPR-Phi3ours), token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) achieve more advantageous results in Recall@1 and have marginal performance improvement in Recall@3/5.
 However, token-level embedding can incur storage costs of 10 times more than a single embedding (DSE requires 0.24GB for indexing MMDocIR while ColPali requires 10.0GB).
@@
 This greatly benefits textual retrievers in multimodal understanding.
-  • Necessity of Token-level Embeddings: For layout retrieval tasks, token-level retrievers marginally outperform dense-level retrievers, demonstrating its importance of in such task.
+  • Comparison of Hybrid vs.
+Pure Image Sequences: Visual retrievers relying on hybrid image-text sequences generally perform less effectively than those utilizing pure image sequences.
+This suggests that current VLMs may have stronger capabilities in modeling images than text within the multimodal framework.
+  • Effect of Token-level Embeddings: For layout retrieval tasks, token-level retrievers marginally outperform dense-level retrievers, demonstrating its importance of such task.
   • Top 10 Coverage: For layout retrieval tasks, retrieving top 10 layouts does not guarantee comprehensive coverage of the ground truth layout labels, emphasizing the complexity of the tasks.
-{tabu}
-ll|ccc|ccc
 Method Page recall Layout recall
 OCR VLM $\Delta$ OCR VLM $\Delta$
-$k=1$
-DPR 22.3 27.2 +4.9 12.6 19.3 +6.7
+$k=1$ DPR 22.3 27.2 +4.9 12.6 19.3 +6.7
 ColBERT 40.3 45.8 +5.5 19.8 31.3 +11.5
@@
 GTE 35.4 38.9 +3.5 18.2 26.1 +7.9
-$k=3\text{ or }5$
-DPR 39.4 46.3 +6.9 23.7 39.8 +16.1
+$k=3\text{ or }5$ DPR 39.4 46.3 +6.9 23.7 39.8 +16.1
 ColBERT 58.8 64.9 +6.1 33.2 54.4 +21.2
@@
 GTE 54.9 58.7 +3.8 33.5 51.3 +17.8
-$k=5\text{ or }10$
-DPR 49.0 57.8 +8.8 29.9 49.5 +19.6
+$k=5\text{ or }10$ DPR 49.0 57.8 +8.8 29.9 49.5 +19.6
 ColBERT 66.0 72.3 +6.3 37.6 61.9 +24.3
@@
 In contrast, only a small fraction of layouts contain more than 380 tokens (3.9% for OCR-text, 4.8% for VLM-text, 0.5% for natural-text).
-Hence, as reflected in Table 4.2 and 5, text retriever demonstrates stronger performance on layout-level retrieval than on page-level retrieval.
-{tabu}
-ll|ccc
+Hence, as reflected in Table 5 and 6, text retriever demonstrates stronger performance on layout-level retrieval than on page-level retrieval.
 Method Layout recall
 Hybrid Image $\Delta$
-$k=1$
-DSEwiki-ss 24.6 28.2 +3.6
+$k=1$ DSEwiki-$s$ 24.6 28.2 +3.6
 DSEdocmatix 27.5 27.9 +0.4
@@
 Col-Phi3ours 29.8 31.1 +1.3
-$k=5$
-DSEwiki-ss 46.7 50.2 +3.5
+$k=5$ DSEwiki-$s$ 46.7 50.2 +3.5
 DSEdocmatix 48.2 49.8 +1.6
@@
 Col-Phi3ours 50.0 52.3 +2.3
-$k=10$
-DSEwiki-ss 55.8 58.5 +2.7
+$k=10$ DSEwiki-$s$ 55.8 58.5 +2.7
 DSEdocmatix 57.4 57.5 +0.1
@@
 Representing text as image tokens requires significantly more resources than native text encoding.
-To address this inefficiency and promote balanced retrieval capabilities, we advocate for future visual retrievers to be jointly trained or fine-tuned on both text and visual retrieval tasks.
+To address this inefficiency and promote balanced retrieval capabilities Dumitru et al.
+(2025); Liang et al.
+(2025), we advocate for future visual retrievers to be jointly trained on both text and visual retrieval tasks using SFT or RL Duong et al.
+(2025).
 Such hybrid training would enable models to efficiently process text when appropriate, without compromising performance on visual inputs.
-{tabu}
-l|ll|ccc Model Store Index Search
-(GB) (MM:SS) (MM:SS)
-Page Processing
-Text
-DPR 0.06 6:53 00:02
+5.7 Cascade Retrieval
+As shown in Table 6, directly performing layout retrieval can be challenging.
+Hence, we propose alternative methods, by perform page-retrieval first, subsequently followed by layout-retrieval within the retrieved page.
+We term such retrieval to be cascade retrieval.
+Note that the page retrieval is not perfect, such error can propagate to layout retrieval and affect the final results.
+page(1st) layout(2nd) Top1 Top5 Top10
+BGE: direct layout 29.0 53.2 62.4
+BGE BGE 24.3 49.0 58.6
+E5: direct layout 26.4 51.8 61.2
+E5 E5 22.2 47.8 58.2
+ColBERT: direct layout 31.4 56.0 63.7
+ColBERT ColBERT 28.5 53.0 61.3
+ColPali: direct layout 32.5 54.3 63.2
+ColPali ColPali 32.7 54.5 63.2
+ColPali ColBERT 31.8 57.0 64.2
+DSE: direct layout 29.1 51.9 59.9
+DSE DSE 29.6 54.0 62.0
+DSE ColBERT 29.0 56.4 64.4
+DPR-Phi3: direct layout 30.2 53.9 62.8
+DPR-Phi3 DPR-Phi3 31.1 54.2 61.7
+DPR-Phi3 ColBERT 30.6 56.6 64.5
+Col-Phi3: direct layout 31.6 54.5 63.3
+Col-Phi3 Col-Phi3 33.3 58.6 63.7
+Col-Phi3 ColBERT 35.3 58.8 65.4
+Table 9: Comparison of 1-stage vs 2-stage approaches across different models
+In this setting, we retrieve top-$k$ pages first, then rerank all layouts belonging to retrieved $k$ pages, and get top-$k$ layouts.
+The cascade retrieval results are shown in Table 9.
+We can observe that method with high page retrieval recall can significantly improve layout retrieval in the reranking paradigm.
+Model Store Index Search
+(GB) ($M$:$S$) ($M$:$S$)
+Page Processing Text DPR 0.06 6:53 00:02
 ColBERT 3.45 14:12 00:04
@@
 E5 0.08 8:26 00:03
-[…]
-[.2pt on 1pt off 1.5pt]2-6
-Image
-DPR-Phi3 0.24 101:20 00:04
+Image DPR-Phi3 0.24 101:20 00:04
 ColPali 10.00 47:14 00:05
 Col-Phi3 24.56 106:23 00:07
-Layout Processing
-Text
-DPR 0.51 41:33 00:15
+Layout Processing Text DPR 0.51 41:33 00:15
 ColBERT 26.72 94:56 02:25
@@
 E5 0.66 60:21 00:18
-[…]
-[.2pt on 1pt off 1.5pt]2-6
-Image
-DPR-Phi3 1.99 735:51 00:44
+Image DPR-Phi3 1.99 735:51 00:44
 ColPali 83.50 262:29 09:06
 Col-Phi3 204.32 784:07 10:56
-[…]
-Hybrid
-DPR-Phi3 1.84 130:38 01:09
+Hybrid DPR-Phi3 1.84 130:38 01:09
 ColPali 12.06 73:50 04:24
@@
 5.8 Efficiency Analysis
-We evaluate the inference efficiency by measuring three key metrics : storage consumption, indexing time and, retrieval latency, as shown in Table  5.6.
-Experiments are conducted on a NVIDIA A100 40GB GPU with batch size of 4 for image and 256 for text.
+We evaluate the inference efficiency by measuring three key metrics : storage consumption, indexing time and, retrieval latency, as shown in Table  10.
+Experiments are conducted with batch size of 4 for image and 256 for text.
 DPR-styled retrievers which generate single vector embeddings, demonstrates higher efficiency and lower computation across all metrics, compared to ColBERT-styled retrievers that produce token-level embeddings.
@@
 (2024) rather than human experts, and are not de-contextualized for retrieval task.
+MMDocRAG Dong et al.
+(2025a) is constructed upon MMDocIR to support multimodal generation.
 ViDoRe Faysse et al.
@@
 In conclusion, multimodal document retrieval presents a complex challenge that requires both understanding and integrating diverse data modalities beyond plain text.
-To more effectively evaluate these capabilities, we introduce the MMDocIR benchmark, which features the innovative dual-task retrieval capabilities targeting page-level and layout-level document granularity.
+To more effectively evaluate these capabilities, we introduce the MMDocIR benchmark, which features the innovative dual‐task retrieval capabilities targeting page-level and layout-level document granularity.
 The MMDocIR includes a rich dataset featuring expertly annotated labels for 1,685 questions and bootstrapped labels for 73,843 questions, serving as a valuable resource for both training and evaluation of multimodal document retrieval.
 Our comprehensive empirical studies show that visual-driven retrievers significantly outperform text-driven ones, underscoring the importance of visual information in improving retrieval performance.
-Future work can expand upon these findings by optimizing retrieval algorithms to enhance both accuracy and efficiency of multimodal document retrieval systems.
+Future work can expand upon these findings by optimizing retrieval algorithms to enhance both accuracy and efficiency of multimodal document retrieval systems, as well as the multilingual capability Liang et al.
+(2020).
 Limitations
@@
 Appendix A Supplementary Experimental Results
-In this section, we provide the full results of page-level and layout-level retrievals, which supplement our main results discussion in Section 5.3 and Section 5.4 respectively.
-Specifically, Table A extends the main page-level results shown in Table 4.2 with the results of text retrievers using OCR-text.
-Table A extends the main layout-level results shown in Table 5 with the results of (i) text retrievers using OCR-text and (ii) visual retrievers using hybrid inputs.
-{tabu}
-lll|cccccccccc|cc
+In this section, we provide the full results of page-level and layout-level retrievals, which supplement our main results discussion in Section 5.3 and Section 7 respectively.
+Specifically, Table 11 extends the main page-level results shown in Table 5 with the results of text retrievers using OCR-text.
+Table 12 extends the main layout-level results shown in Table 6 with the results of (i) text retrievers using OCR-text and (ii) visual retrievers using hybrid inputs.
 Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
 Report &Indu. Worksh. Paper ure Report book ment Macro Micro
-Recall@$k=1$
-OCR-text
-DPR 21.2 22.1 27.7 23.3 24.4 16.7 21.1 20.7 31.0 15.1 22.3 21.7
+Recall@$k=1$ OCR-text DPR 21.2 22.1 27.7 23.3 24.4 16.7 21.1 20.7 31.0 15.1 22.3 21.7
 ColBERT 43.8 39.8 42.4 39.3 39.2 38.7 46.3 50.6 46.1 17.1 40.3 40.0
@@
 GTE 44.6 32.6 45.0 33.2 37.2 31.8 40.0 39.9 35.2 14.5 35.4 34.6
-[…]
-VLM-text
-DPR 32.3 25.5 27.0 31.0 28.4 18.8 23.5 31.2 38.3 16.1 27.2 26.9
+VLM-text DPR 32.3 25.5 27.0 31.0 28.4 18.8 23.5 31.2 38.3 16.1 27.2 26.9
 ColBERT 48.6 42.8 51.1 46.2 36.0 36.8 49.6 60.9 59.5 26.3 45.8 44.9
@@
 GTE 46.5 26.3 48.7 38.9 35.9 27.0 46.2 50.1 45.8 24.1 38.9 37.9
-[…]
-Image
-DSEwiki-ss 53.0 50.0 54.0 48.7 45.1 43.0 51.5 46.9 54.2 33.6 48.0 47.5
+Image DSEwiki-$s$ 53.0 50.0 54.0 48.7 45.1 43.0 51.5 46.9 54.2 33.6 48.0 47.5
 DSEdocmatix 52.3 40.4 56.1 51.7 45.8 43.5 53.8 53.7 58.3 46.7 50.2 50.1
@@
 Col-Phi3ours 56.7 50.4 56.9 61.3 54.8 50.7 60.8 61.3 63.6 54.0 57.0 57.1
-Recall@$k=3$
-OCR-text
-DPR 46.1 40.6 38.9 46.7 43.9 32.4 38.4 37.0 50.0 20.4 39.4 39.8
+Recall@$k=3$ OCR-text DPR 46.1 40.6 38.9 46.7 43.9 32.4 38.4 37.0 50.0 20.4 39.4 39.8
 ColBERT 72.6 59.7 57.8 66.7 60.0 53.7 63.8 68.5 61.4 23.7 58.8 59.5
@@
 GTE 69.2 47.0 58.7 59.5 61.8 46.6 65.5 59.1 61.4 19.7 54.9 54.7
-[…]
-VLM-text
-DPR 52.2 44.2 43.5 54.6 52.0 35.1 44.4 53.9 57.2 25.5 46.3 46.2
+VLM-text DPR 52.2 44.2 43.5 54.6 52.0 35.1 44.4 53.9 57.2 25.5 46.3 46.2
 ColBERT 70.1 64.4 70.3 72.3 59.1 55.3 71.1 81.3 70.8 34.3 64.9 64.8
@@
 GTE 71.1 44.5 67.2 64.4 54.3 43.0 70.6 71.9 68.2 31.4 58.7 58.3
-[…]
-Image
-DSEwiki-ss 75.4 65.0 73.9 79.8 69.5 63.5 75.4 71.5 81.4 50.4 70.6 71.4
+Image DSEwiki-$s$ 75.4 65.0 73.9 79.8 69.5 63.5 75.4 71.5 81.4 50.4 70.6 71.4
 DSEdocmatix 75.4 67.5 73.3 80.0 66.3 61.6 72.8 76.4 82.6 57.7 71.4 71.8
@@
 Col-Phi3ours 80.2 74.1 77.4 84.8 69.1 67.7 78.7 79.5 81.8 69.3 76.3 76.8
-Recall@$k=5$
-OCR-text
-DPR 59.5 55.8 43.4 59.1 56.2 41.2 50.7 45.5 56.0 23.0 49.0 49.4
+Recall@$k=5$ OCR-text DPR 59.5 55.8 43.4 59.1 56.2 41.2 50.7 45.5 56.0 23.0 49.0 49.4
 ColBERT 78.4 71.1 63.3 75.2 68.8 60.5 72.0 72.7 67.5 30.3 66.0 66.5
@@
 GTE 78.3 61.9 67.3 72.3 68.7 55.2 72.6 64.0 67.5 24.3 63.2 63.5
-[…]
-VLM-text
-DPR 66.5 60.1 56.0 68.9 58.8 43.8 57.1 68.6 64.8 33.6 57.8 57.8
+VLM-text DPR 66.5 60.1 56.0 68.9 58.8 43.8 57.1 68.6 64.8 33.6 57.8 57.8
 ColBERT 78.8 74.0 78.7 82.3 66.1 60.8 77.0 88.5 78.0 38.7 72.3 72.3
@@
 GTE 77.4 62.6 74.7 75.8 62.0 51.8 77.8 80.0 75.0 39.4 67.6 67.2
-[…]
-Image
-DSEwiki-ss 84.0 80.2 78.7 87.0 75.7 73.0 82.0 77.3 88.3 58.4 78.5 79.2
+Image DSEwiki-$s$ 84.0 80.2 78.7 87.0 75.7 73.0 82.0 77.3 88.3 58.4 78.5 79.2
 DSEdocmatix 82.1 77.2 79.6 87.8 73.9 72.4 81.7 83.1 89.4 67.9 79.5 80.1
@@
 “Image” refers to processing document page as screenshot image.
-{tabu}
-llll|cccccccccc|cc
 Method Domain Resear. Admin Tutori.& Acade. Broch- Finance Guide- Govern- Laws News Average
 Report &Indu. Worksh. Paper ure Report book ment Macro Micro
-Recall@$k=1$
-Textual Retrieval
-OCR-text
-DPR 3.4 7.2 1.2 11.3 3.0 9.8 8.2 24.7 30.9 26.3 12.6 12.4
+Recall@$k=1$ Textual Retrieval OCR-text DPR 3.4 7.2 1.2 11.3 3.0 9.8 8.2 24.7 30.9 26.3 12.6 12.4
 ColBERT 5.0 8.8 4.7 16.4 2.0 13.2 4.6 50.8 47.7 45.3 19.8 19.1
@@
 GTE 5.7 7.0 3.9 17.2 2.8 15.8 9.4 40.7 41.6 38.3 18.2 18.4
-[…]
-VLM-text
-DPR 11.6 9.5 19.2 19.2 14.9 15.9 15.8 25.6 34.7 27.0 19.3 19.2
+VLM-text DPR 11.6 9.5 19.2 19.2 14.9 15.9 15.8 25.6 34.7 27.0 19.3 19.2
 ColBERT 22.0 14.9 28.0 28.3 17.9 29.7 21.1 52.6 54.5 44.5 31.3 31.4
@@
 GTE 17.5 10.5 23.0 27.2 14.5 26.3 14.4 39.8 49.2 38.3 26.1 27.1
-[…]
-Visual Retrieval
-Pure-Image
-DSEwiki-ss 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
+Visual Retrieval Pure-Image DSEwiki-$s$ 20.6 15.1 31.0 31.1 20.1 29.2 22.0 39.3 37.5 35.8 28.2 29.2
 DSEdocmatix 19.9 11.4 31.5 30.1 17.8 30.0 20.8 46.5 39.4 31.4 27.9 29.1
@@
 Col-Phi3ours 22.6 22.0 37.5 34.9 28.9 30.3 22.7 50.2 45.1 26.3 31.1 31.6
-[…]
-Hybrid
-DSEwiki-ss 14.0 10.4 29.8 18.0 13.7 20.4 13.5 46.0 45.1 34.7 24.6 23.4
+Hybrid DSEwiki-$s$ 14.0 10.4 29.8 18.0 13.7 20.4 13.5 46.0 45.1 34.7 24.6 23.4
 DSEdocmatix 18.2 11.6 32.7 24.0 17.7 27.2 16.7 48.1 45.5 33.0 27.5 27.4
@@
 Col-Phi3ours 26.4 12.6 33.7 27.3 30.1 27.9 24.6 46.2 47.4 21.9 29.8 29.6
-Recall@$k=5$
-Textual Retrieval
-OCR-text
-DPR 7.3 12.5 5.6 24.0 8.9 16.9 13.6 47.2 50.2 51.1 23.7 23.5
+Recall@$k=5$ Textual Retrieval OCR-text DPR 7.3 12.5 5.6 24.0 8.9 16.9 13.6 47.2 50.2 51.1 23.7 23.5
 ColBERT 10.9 23.8 10.2 32.2 6.8 25.5 17.0 78.7 63.4 63.2 33.2 32.6
@@
 GTE 10.0 18.2 12.8 32.9 15.2 29.4 21.0 67.7 65.3 62.4 33.5 33.5
-[…]
-VLM-text
-DPR 31.0 25.7 36.7 44.9 33.0 34.1 34.9 49.9 56.3 51.1 39.8 40.4
+VLM-text DPR 31.0 25.7 36.7 44.9 33.0 34.1 34.9 49.9 56.3 51.1 39.8 40.4
 ColBERT 41.8 37.7 53.7 61.8 35.1 52.4 46.1 83.2 70.1 62.5 54.4 56.0
@@
 GTE 36.7 25.6 51.2 56.6 39.7 46.6 46.4 72.2 74.3 63.2 51.3 52.3
-[…]
-Visual Retrieval
-Pure-Image
-DSEwiki-ss 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
+Visual Retrieval Pure-Image DSEwiki-$s$ 42.4 32.9 56.3 58.5 39.8 50.6 41.6 68.6 60.9 50.0 50.2 52.1
 DSEdocmatix 39.6 36.2 53.9 57.5 33.7 52.5 42.8 69.4 63.1 48.9 49.8 51.9
@@
 Col-Phi3ours 46.4 38.2 53.1 61.8 45.0 54.6 45.7 68.8 65.7 43.8 52.3 54.5
-[…]
-Hybrid
-DSEwiki-ss 31.8 29.5 51.1 43.0 34.5 39.3 38.3 71.3 71.3 57.1 46.7 45.6
+Hybrid DSEwiki-$s$ 31.8 29.5 51.1 43.0 34.5 39.3 38.3 71.3 71.3 57.1 46.7 45.6
 DSEdocmatix 37.3 26.7 48.2 49.7 34.5 48.6 41.0 72.2 69.4 54.2 48.2 49.3
@@
 Col-Phi3ours 50.9 25.5 49.1 58.3 41.9 48.1 49.2 62.3 60.6 48.5 50.0 51.8
-Recall@$k=10$
-Textual Retrieval
-OCR-text
-DPR 10.5 21.1 8.8 32.0 14.9 19.6 17.0 59.7 56.4 58.9 29.9 29.3
+Recall@$k=10$ Textual Retrieval OCR-text DPR 10.5 21.1 8.8 32.0 14.9 19.6 17.0 59.7 56.4 58.9 29.9 29.3
 ColBERT 14.1 31.1 13.1 38.4 9.1 31.0 22.4 83.9 67.9 65.4 37.6 37.4
@@
 GTE 19.1 29.0 21.4 39.0 19.2 32.9 29.2 78.8 74.5 66.2 40.9 40.1
-[…]
-VLM-text
-DPR 42.2 33.1 52.1 56.2 39.9 43.5 44.0 62.8 61.7 59.7 49.5 50.5
+VLM-text DPR 42.2 33.1 52.1 56.2 39.9 43.5 44.0 62.8 61.7 59.7 49.5 50.5
 ColBERT 51.0 48.7 60.6 69.8 43.9 61.6 53.7 88.4 74.8 66.4 61.9 63.7
@@
 GTE 48.6 41.1 61.5 68.8 44.3 56.9 58.0 83.0 77.5 66.9 60.7 62.2
-[…]
-Visual Retrieval
-Pure-Image
-DSEwiki-ss 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
+Visual Retrieval Pure-Image DSEwiki-$s$ 55.9 41.3 61.5 68.1 47.8 60.7 54.2 72.9 68.3 54.4 58.5 61.1
 DSEdocmatix 53.7 43.3 59.6 66.5 44.7 59.1 50.3 75.4 69.2 53.7 57.5 59.9
@@
 Col-Phi3ours 57.7 50.5 66.6 72.3 50.7 59.3 53.6 68.5 74.8 57.5 61.1 63.3
-[…]
-Hybrid
-DSEwiki-ss 44.1 34.3 57.6 56.3 42.6 50.7 48.6 81.1 79.1 63.7 55.8 56.0
+Hybrid DSEwiki-$s$ 44.1 34.3 57.6 56.3 42.6 50.7 48.6 81.1 79.1 63.7 55.8 56.0
 DSEdocmatix 49.9 37.3 57.3 61.3 45.9 57.9 50.1 77.9 74.9 61.5 57.4 58.9
@@
 (2024b) and DocBench Zou et al.
-(2024), treat DocVQA as a long-context task, accommodating entire documents that average between 50–70 pages in length.
+(2024), treat DocVQA as a long-context task, accommodating entire documents that average between 50 to 70 pages in length.
 As document lengths increase, retrieval becomes essential.
@@
 B.2 Document Corpora Collection Criteria
-To facilitate the development of MMDocIR, we leverage resources from existing DocVQA benchmarks described in Appendix B.1.
+Early document related benchmarks Dong et al.
+(2021) are mostly textual only, which are not considered in MMDocIR.
+To facilitate the development of MMDocIR, we leverage visually-rich documents from recent DocVQA benchmarks described in Appendix B.1.
 Despite not being curated for IR, they offer valuable document corpora and questions that can be adapted for IR tasks.
@@
   • DUDE Landeghem et al.
-(2023) provides 5,019 documents from aggregator websites7771: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/.
+(2023) provides 5,019 documents from aggregator websites1212121: archive.org, 2: http://commons.wikimedia.org/, 3: http://documentcloud.org/.
 It covers a broad range of domains, including medical, legal, technical, and financial, among others, to evaluate models’ ability to handle diverse topics and the specific knowledge each requires.
@@
 B.6 Hard Negative Sampling
-In addition to annotating ground truth (positive) page labels, we enhance our training data with negative labels.
+In addition to annotating ground truth (positive) page labels, we enhance our training data with negative labels Li et al.
+(2025).
 In the context of retrieval, hard negatives are particularly informative non-relevant documents that closely resemble true positives according to the model’s current scoring function.
@@
 The irrelevant pages with the highest top-$k$ scores (i.e., those most likely to be confused with positives) are selected as hard negatives for training.
+In the future, we consider to incorporate content planning Bao et al.
+(2022b) and synthetic methods Bao et al.
+(2023, 2022a) for hard negative generation.
 Artifacts Purpose Referred Section Resource URL
@@
 DocBench Eval-set curation https://github.com/Anni-Zou/DocBench
-MP-DocVQA Train-set curation Section 4 Appendix B.4 https://rrc.cvc.uab.es/?ch=17&com=tasks
+MP-DocVQA Train-set curation
+Section 4
+Appendix B.4
+https://rrc.cvc.uab.es/?ch=17&com=tasks
 SlideVQA Train-set curation https://github.com/nttmdlab-nlp/SlideVQA
@@
 OpenCV Image matching https://opencv.org/
-Text Retriever Layout location BGE and E5 (see Table D.1)
-Visual Retriever Hard negatives Appendix B.6 Colpali (see Table D.1)
+Text Retriever Layout location BGE and E5 (see Table 14)
+Visual Retriever Hard negatives Appendix B.6 Colpali (see Table 14)
 Table 13: Artifacts used to facilitate construction of MMDocIR evaluation & train set.
@@
 $\text{Sim}(q,d)_{dpr}=\frac{\langle\mathrm{E_{q}^{dpr}}|\mathrm{E_{d}^{dpr}}\rangle}{\left\|\mathrm{E_{q}^{dpr}}\right\|\cdot\left\|\mathrm{E_{d}^{dpr}}\right\|}$ (3)
-[…]
+where $Sim(q,d)_{dpr}$ is computed as the cosine similarity between their embeddings. and $\langle\cdot|\cdot\rangle$ is the dot product.
 $\text{Sim}(q,d)_{col}=\sum_{i\in[1,N_{q}]}\max_{j\in[1,N_{d}]}\langle\mathrm{E_{q}^{col}}^{(i)}|\mathrm{E_{d}^{col}}^{(j)}\rangle$ (4)
-[…]
+where $Sim(q,d)_{col}$ is the sum over all query vectors $\mathrm{E_{q}^{col}}^{(i)}$, of its maximum dot product $\langle\cdot|\cdot\rangle$ with each of the $N_{d}$ document embedding vectors $\mathrm{E_{d}^{col}}^{(j)}$.
 C.3 Contrastive Loss
@@
 We calculate the loss as:
-($\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$ (5)
+$\mathcal{L}^{dpr}_{(q,d^{+},d^{-})}=-\log\frac{\exp(\text{Sim}^{dpr}_{(q,d^{+})}/\tau)}{\sum_{d_{i}\in d^{+}\cup d^{-}}\exp(\text{Sim}^{dpr}_{(q,d_{i})}/\tau)}$ (5)
 where DPR-Phi3 is trained on the InfoNCE loss, and the temperature parameter $\tau=0.02$ in our experiments.
-($\mathcal{L}^{col}_{(q,d^{+},d^{-})}=$ $\log\Big{(}1+\exp\big{(}\max_{d_{i}\in d^{-}}(\text{Sim}^{col}_{(q,d_{i})})$
-$-\text{Sim}^{col}_{(q,d^{+})}\big{)}\Big{)}$ (6)
+[…]
+$)\displaystyle-\text{Sim}^{col}_{(q,d^{+})}\big$$\Big)$ (6)
 where Col-Phi3 is trained via the softplus loss based on the positive scores w.r.t. to the maximal negative scores.
@@
 (2020)) and a list of token-level embeddings (following ColBERT Khattab and Zaharia (2020)), respectively.
-To train the model, we employ memory-efficient techniques such as LoRA Hu et al.
+To train the model, we employ memory-efficient techniques such as PERF Zhang et al.
+(2024b, a), LoRA Hu et al.
 (2022), FlashAttention Dao (2024), and DeepSpeed Rasley et al.
@@
 In both tasks, each training query is paired with one positive document and one hard negative document.
-The document screenshots are resized to $344\times 1,$ 344 × 1 , 344 pixels and cropped into $4\times 4$ sub-images.
-The training process is performed on four A100 80GB GPUs.
+The document screenshots are resized to $1,344\times 1,344$ pixels and cropped into $4\times 4$ sub-images.
 Appendix D Retrievers: Introduction and Implementation Details
@@
 (2023); Xu et al.
-(2020, 2021), Chunking Chen et al.
-(2024a); Raina and Gales (2024); Dong et al.
-(2024), and Image Captioning You et al.
+(2020, 2021), Information extraction Dong et al.
+(2022, 2023a), Chunking Chen et al.
+(2024a); Raina and Gales (2024), and Image Captioning You et al.
 (2016); Aneja et al.
 (2018).
-These steps are time-consuming and can introduce errors that impact the overall retrieval performance.
-Current text retrieval are primarily categorized as sparse or dense retrieval.
+These steps are time-consuming and can introduce errors that impact the overall retrieval performance Wu et al.
+(2025); Nie et al.
+(2023); Li et al.
+(2023a).
+Current text retrieval are primarily categorized as sparse or dense retrieval on chunks Dong et al.
+(2023b).
 For two widely-used sparse retrievers: TF-IDF Salton et al.
@@
 Moreover, GTE Li et al.
-(2023) integrates graph-based techniques to enhance dense embedding.
+(2023b) integrates graph-based techniques to enhance dense embedding.
 However, most text retrieval systems overlook valuable visual information present in documents.
-{tabu}
-l|l|c|c|l Model Dimension Base Model HuggingFace Checkpoint
-Text
-DPR 768 BERT-base facebook/dpr-ctx_encoder-multiset-base facebook/dpr-question_encoder-multiset-base
-[…]
+Model Dimension Base Model HuggingFace Checkpoint
+Text DPR 768 BERT-base
+facebook/dpr-ctx_encoder-multiset-base
+facebook/dpr-question_encoder-multiset-base
 ColBERT $N_{\mathrm{tok}}\times$768 BERT-base colbert-ir/colbertv2.0
-[…]
 Contriever 768 BERT-base facebook/contriever-msmarco
-[…]
 E5 1,024 BERT-large intfloat/e5-large-v2
-[…]
 BGE 1,024 RetroMAE BAAI/bge-large-en-v1.5
-[…]
 GTE 1,024 BERT-large thenlper/gte-large
-Visual
-DSEwiki-ss 3,072 Phi-3-Vision Tevatron/dse-phi3-v1.0
-[…]
+Visual DSEwiki-$s$ 3,072 Phi-3-Vision Tevatron/dse-phi3-v1.0
 DSEdocmatix 3,072 Phi-3-Vision Tevatron/dse-phi3-docmatix-v2
-[…]
-ColPali $N_{\mathrm{tok}}\times$1,024 PaliGemma vidore/colpali
+ColPali $N_{\mathrm{tok}}\times$128 PaliGemma vidore/colpali
 Table 14: Implementation details for Text and Vision Retrieval Models
@@
 The visual retrieval models deployed are namely DSEwiki-ss, DSEdocmatix, and ColPali.
-We use pre-trained checkpoints available on HuggingFace  [^13] https://huggingface.co/; the specific checkpoint information can be found in Table D.1 alongside other configuration details.
+We use pre-trained checkpoints available on HuggingFace  [^18] https://huggingface.co/; the specific checkpoint information can be found in Table 14 alongside other configuration details.
 Appendix E Dataset Demonstration
@@
 This specificity and simplism are functional within the confines of the intended VQA context but pose substantial limitations when such queries are repurposed for IR tasks.
-Due to this gap, we exclude ViDoRe benchmark from our experiments.
 F.2 Document Corpora Analysis